### v1

# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

Abstract
Offline preference-based reinforcement learning (PbRL) provides an effective way to overcome the challenges of designing reward and the high costs of online interaction. However, since labeling preference needs real-time human feedback, acquiring sufficient preference labels is challenging. To solve this, this paper proposes a offLine prEference-bAsed RL with high Sample Efficiency (LEASE) algorithm, where a learned transition model is leveraged to generate unlabeled preference data. Considering the pretrained reward model may generate incorrect labels for unlabeled data, we design an uncertainty-aware mechanism to ensure the performance of reward model, where only high confidence and low variance data are selected. Moreover, we provide the generalization bound of reward model to analyze the factors influencing reward accuracy, and demonstrate that the policy learned by LEASE has theoretical improvement guarantee. The developed theory is based on state-action pair, which can be easily combined with other offline algorithms. The experimental results show that LEASE can achieve comparable performance to baseline under fewer preference data without online interaction.
Index Terms:
Preference-based reinforcement learning; Sample efficiency; Reward learning; Theoretical guarantee
I Introduction
Reinforcement learning (RL) has been widely applied in robot control [1, 10] and healthcare [12, 11] fields. However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [40]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9]. Many researches have been explored to solve the above two challenges.
Offline RL, learning policy from previous collected dataset without interaction with environment, has become an effective way for challenge 1 [4, 8]. However, due to limited data coverage, offline RL suffers from the distribution shift between the learned policy and the behavior policy [28], which leads to poor agent performance. To address this problem, offline RL either directly constrain the policy closed to the behavior policy [2] or penalize value function for out-of-distribution (OOD) data [4, 8]. The above methods improve the performance of agent through incorporating conservatism into policy learning.
Preference-based RL (PbRL), also known as RL from human feedback (RLHF), which learns the reward function from human preference data without the demand for tedious hand-engineered reward design [14, 43], has been developed recently for challenge 2. However, compared with data $(s,a,s^{\prime})$, the collecting cost of preference data is higher since it demands human real-time feedback to label preference, which often brings high sample complexity (low sample efficiency) [13, 3]. Some data augmentation techniques, such as generating unlabeled data through interaction with simulator [3] or temporal cropping [35], have been proposed to solve problem of the limitation of preference data. However, the above methods need online interaction with environment (simulator) and still demand large human feedback. Thus, in PbRL, the sample efficiency is still of utmost importance.
In offline PbRL setting, existing methods [34, 17, 33] typically involve two steps: reward learning and policy learning. They learns the policy based on offline static dataset without reward and preference dataset, and introduce transformer [17] or diffusion model [33] to improve the performance of reward model and agent performance. However, the theoretical guarantees of reward generalization and policy improvement are not provided. Therefore, this paper aims to design a high sample efficiency offline PbRL algorithm that can achieve comparable performance with baseline from fewer preference data without interaction with environment, and develop the related theory for reward generalization bound and policy improvement guarantee. The key questions can be summarized as follow:
  1. 1.
    In algorithm, the collected preference data fails to cover all state-action region and the collecting cost is too high. How to use the limited preference data to learn accurate reward model and guarantee agent performance?
  2. 2.
    In theory, what factors affect the reward generalization and what is the detailed relationship between offline RL algorithm itself, the reward gap and the improvement of policy performance?
For question 1, motivated by model-based offline RL, which learns the environment model to broaden data coverage [37, 36], we also train the transition model to achieve data augmentation (improve sample efficiency). Specifically, the agent generates two different trajectories through interaction with the learned transition model, and then the pretrained reward model generates pseudo label for these two trajectories. However, the preference model may generate incorrect pseudo label, which may bring unstable training and low performance.
For question 2, there are very fewer algorithms for offline PbRL theory. The most relevant work is [18], where the theory of policy improvement guarantee is established. However, the generalization bound for reward model is not considered, and this theory is based on the whole trajectory rather than state-action pair in offline algorithms, which is not conducive to the theoretical analysis of specific offline PbRL algorithms since many offline RL are based on state-action pair.
Figure 1: The learning framework of LEASE. The offline dataset without reward is used to train transition model and limited labeled preference dataset is used to pretrain reward model. Then the generated unlabeled preference data is screened through uncertainty-aware selecting mechanism. The reward model is updated based on labeled and generated dataset. Finally, the agent learns the policy based on offline dataset and learned reward model.
To solve the above problems, this paper proposes a novel offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE), where a selecting mechanism is designed to guarantee the quantity of generated dataset. This selecting mechanism is based on the principles of high confidence and low uncertainty, and can effectively reduce the error of pseudo-labels. Moreover, we develop a new generalization bound for reward model, and provide the theory of policy improvement guarantee for offline PbRL based on state-action pair. Fig. 1 shows the learning framework of LEASE. The theoretical and experimental results show that LEASE has policy improvement guarantee and can achieve comparable performance to baseline under fewer preference data on D4RL benchmark. The contributions are given below:
  1. 1.
    A novel learning framework, named LEASE, is proposed for offline PbRL, where an innovative preference augmentation technique is utilized to improve sample efficiency.
  2. 2.
    An uncertainty-aware mechanism is designed for screening the generated preference data so as to guarantee the stability of reward training and improve the accuracy of reward prediction.
  3. 3.
    The generalization bound of reward model and the theory of policy improvement based on state-action pair are developed to analyze factors that influence reward and policy performance.
The framework of this paper is as follows: Section II introduces the related works about offline PbRL. Section III provides the basic notations for algorithm. Section IV illustrates the two main problems that this paper aims to solve. Section V introduces the details of designed selecting mechanism and derives the generalization bound for reward model. Section VI illustrates the implementation details of the proposed method and provides theory for offline PbRL. Section VII conducts experiments to validate the effectiveness of LEASE and analyze the effects of each part for agent performance. Section VIII further discusses the offline PbRL theory and algorithm. Section IX summarizes the entire work.
II Related Work
This section presents a brief overview for the algoriths and related theories of offline PbRL, and the differences and advantages between our work and the previous related works.
The algorithm for offline PbRL. Offline PbRL eliminates the demand for interaction with environment and handcraft designed reward. OPAL [34] is the first algorithm that combining offline RL and PbRL. PT [17] utilized transformer-based architecture to design preference model capable of generating non-Markovian rewards. OPPO [30] directly optimized the policy in a high-level embedding space without learning a separate reward function. However, the above methods demands a large amount of preference dataset. FTB [33] trains a diffusion model to achieve data augmentation, but this consumers larger training cost. Instead, our method LEASE trains simple transition model to generate unlabeled data and designed selecting mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer dataset and less time.
The theory for offline PbRL. There are few algorithms that provide theoretical guarantees for offline PbRL, including the generalization bound of reward model and the guarantee of policy improvement. Zhu et al.[27] studied offline PbRL, but the analysis are restricted to linear model. Zhan et al.[18] extended to general function approximation, but the generalization analysis of reward model is not provided and the theory is based on trajectory. LEASE gives the theoretical analysis for reward model, and provides the theoretical guarantee for policy based on state-action pair. The generalization bound of reward model has generality and can be applicable to methods that train reward model using pseudo-labels, and the theory of policy improvement guarantee of offline PbRL can be easily combined with other offline RL.
Data selection techniques. Data screening mechanism essentially belongs to the field of semi-supervised learning. The goal of semi-supervised learning is using unlabeled data to improve model performance when labeled data is limited. Data selection techniques is used to filter the data with clean labels from a noisy dataset. Han et al.[7] selected unlabeled data with small losses from one network. Li et al.[5] used a two-component GMM to separate the dataset into a clean set and a noisy set. Rizve et al.[31] modeled the prediction uncertainty of unlabeled data to screen data. Xiao et al.[6] selected data with high confidence as clean data. Motivated by [31] and [6], we trained ensemble reward model to select data with high confidence and low variance to guarantee the quality of unlabeled data.
In PbRL methods, Surf [3] is the most similar to our method LEASE that uses data augmentation technique. The differences between them mainly includes the below three aspects: 1) Surf belongs to online RL, but LEASE belongs to offline RL. Surf generates data through interaction with environment (simulator) while LEASE generates data through the learned transition model. Note that, in most human-in-loop-control scenarios, the cost of online interaction is large and the realistic simulator is difficult to create; 2) Surf only uses confidence for label filtering, whereas LEASE employs both confidence and uncertainty principles, which effectively reduce pseudo-label error. The results in Table III have validated this; 3) LEASE provides the general theoretical framework for offline PbRL, but Surf don’t provide theoretical analysis.
III Preliminaries
In this section, we provides the basic notations for reinforcement learning algorithm, including offline RL, preference-based RL and model-based RL.
Offline reinforcement learning. The framework of RL is based on the Markov Decision Process (MDP) that is described by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},R,T,\rho,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})$ is the transition dynamics, $\rho$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor [40]. The term $\Delta({\Omega})$ denotes the set of probability distribution over space $\Omega$. The goal of RL is to optimize the policy $\pi$ that maximizes the expected discounted return $J(\pi,R):=\mathbb{E}_{(s,a)\sim d_{T}^{\pi}(s,a)}[R(s,a)]/(1-\gamma)$, where $d_{T}^{\pi}(s,a):=d_{T}^{\pi}(s)\pi(a|s)$ is the state-action marginal distribution under the learned policy $\pi$. The discounted state marginal distribution $d_{T}^{\pi}(s)$ is denoted as $(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathcal{P}(s_{t}=s|\pi)$, where $\mathcal{P}(s_{t}=s|\pi)$ is the probability of reaching state $s$ at time $t$ by rolling out $\pi$. The policy $\pi$ can be derived from $Q$-learning [41], which learns the $Q$-function that satisfies Bellman operator: $\mathcal{T}Q(s,a):=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T_{\mathcal{M}}(s^{\prime}\mid s,a)}[\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})]$.
The goal of offline RL is to learn a policy from offline dataset $\mathcal{D}_{\text{offline}}=\{\left(s_{i},a_{i},r_{i},s_{i}^{\prime}\right)\}_{i=1}^{C}$ collected by the behavior policy $\mu$. The policy learning includes two parts: policy evaluation (minimizing the Bellman error) and policy improvement (maximizing the $Q$-function), that is
$\widehat{Q}\leftarrow\arg\min_{Q}\mathbb{E}_{s,a,s^{\prime}\sim\mathcal{D}}\big{[}Q(s,a)-\widehat{\mathcal{T}}\widehat{Q}(s,a)\big{]}^{2},$ (1)
$\widehat{\pi}\leftarrow\arg\max_{\pi}\mathbb{E}_{s\sim\mathcal{D},a\sim\pi}\big{[}\widehat{Q}(s,a)\big{]},$
where $\mathcal{D}$ can be a fixed offline dataset or replay buffer generated by the current policy $\widehat{\pi}$ interacting with the environment. The operator $\widehat{\mathcal{T}}$ is the Bellman operator based on sample, denoted as $\widehat{\mathcal{T}}Q(s,a):=R(s,a)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q\left(s^{\prime},a^{\prime}\right)$ [4].
Preference-based reinforcement learning. Different from standard RL setting, the reward function is not available in PbRL. Instead, PbRL learns the reward function $\widehat{R}$ from preferences between pairs of trajectory segments to align the human intention [15], where a trajectory segment $\sigma$ of length $L$ is defined as a sequence of states and actions $\{s_{k},a_{k},...,s_{k+L-1},a_{k+L-1}\}\in(\mathcal{S}\times\mathcal{A})^{L}$. Given a pair of segments $(\sigma_{0},\sigma_{1})$, human choose which segment is preferred, i.e., $y\in\{0,1\}$ [14]. The preference label $y=1$ indicates $\sigma_{1}\succ\sigma_{0}$ and $y=0$ indicates $\sigma_{0}\succ\sigma_{1}$, where $\sigma_{i}\succ\sigma_{j}$ denotes that the segment $i$ is preferable to the segment $j$.
The preference data is stored in the dataset $\mathcal{D}_{l}$ as a tuple $(\sigma_{0},\sigma_{1},y)$. To obtain the reward function $\widehat{R}$ parameterized by $\psi$, the Bradley-Terry model [16] is utilized to define a preference predictor following previous works [17, 32]:
$P_{\psi}\left(\sigma_{0}\succ\sigma_{1}\right)=\frac{\exp{\sum_{t}\widehat{R}_{\psi}(s_{t}^{0},a_{t}^{0})}}{\exp{\sum_{t}\widehat{R}_{\psi}(s_{t}^{0},a_{t}^{0})}+\exp{\sum_{t}\widehat{R}_{\psi}(s_{t}^{1},a_{t}^{1})}}.$ (2)
Then, based on the collected preference dataset $\mathcal{D}_{l}$, the reward function $\widehat{R}_{\psi}$ can be updated through minimizing the cross-entropy loss between the predicted and the true preference labels [32]:
$\mathcal{L}_{R}(\psi)=-\mathbb{E}_{(\sigma_{0},\sigma_{1},y)\sim\mathcal{D}_{l}}\big{[}$ $y\log P_{\psi}\left(\sigma_{1}\succ\sigma_{0}\right)$ (3)
$+$ $(1-y)\log P_{\psi}\left(\sigma_{0}\succ\sigma_{1}\right)\big{]}.$
After obtaining the reward function $\widehat{R}$, offline PbRL optimizes the policy by maximizing the expected discounted return $J(\pi,R)$ like standard offline RL.
Model-based reinforcement learning. Model-based RL learns the dynamics model to improve sample efficiency [38, 36, 39]. They used the offline dataset $\mathcal{D}_{\text{offline }}$ to estimate transition model $\widehat{T}(s^{\prime}|s,a)$. The transition model $\widehat{T}_{\varphi}$ parameterized by $\varphi$ is typically trained via maximum likelihood estimation (MLE):
$\mathcal{L}_{T}(\varphi)=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\text{offline }}}\big{[}\log\widehat{T}_{\varphi}(s^{\prime}~{}|~{}s,a)\big{]}.$ (4)
In practical implantation, the transition model is approximated by Gaussian distribution and the MLE method is employed to train $N_{T}$ ensemble transition models $\{T^{i}_{\varphi}=\mathcal{N}(\mu^{i}_{\varphi},\sigma^{i}_{\varphi})\}_{i=1}^{N_{T}}$ [37]. The samples are generated through $H$-step rollouts. Here, we use the trained transition to generate more trajectory to achieve data augmentation.
IV Problem Formulation
The cost of collecting preference data is high since it needs real-time human feedback. Therefore, this paper aims to improve sample efficiency to learn accurate reward model from limited preference data and guarantee agent performance. For PbRL, the ideal form is that the learned reward function $\widehat{R}$ from collected preference data can be consistent with true reward $R^{*}$. Here, we define the function class as =$\mathcal{R}=\{R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$. Then we assume the reward class $\mathcal{R}$ is realizable.
Assumption 1 (Realizability).
Let $d(s,a)\in\Delta(\mathcal{S}\times\mathcal{A})$ be the arbitrary data distribution. Then, for any distribution $d(s,a)$, $\inf_{R\in\mathcal{R}}\mathbb{E}_{(s,a)\sim d(s,a)}[R^{*}(s,a)-R(s,a)]^{2}<\varepsilon_{R}$ holds.
Based on the above assumption, the optimal reward function can be obtained if the preference data can cover all data space. However, preference data tend to be limited, even less than hundred in real-world scenarios, such as rehabilitation robot field. Therefore, different from previous PbRL [3, 35, 33], this paper learns reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$. We train the transition model through Eq. (4) to generate more unlabeled preference data $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$. The pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
$\widehat{y}\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)=\mathbbm{1}\big{[}P\left(\sigma_{1}^{u}\succ\sigma_{0}^{u};{\psi}\right)>0.5\big{]},$ (5)
where $\mathbbm{1}(\cdot)$ is indicator function.The reward function is updated through collected labeled dataset $\mathcal{D}_{l}$ and generated unlabeled dataset $\mathcal{D}_{u}$. Then, the reward model can be optimized through minimizing
$\mathcal{L}^{\prime}_{R}(\psi)=$ $\mathbb{E}_{(\sigma_{0}^{l},\sigma_{1}^{l},y)\sim\mathcal{D}_{l}}\big{[}L\left((\sigma_{0}^{l},\sigma_{1}^{l}),y\right)\big{]}$ (6)
$+$ $\mathbb{E}_{(\sigma_{0}^{u},\sigma_{1}^{u})\sim\mathcal{D}_{u}}\big{[}L\left((\sigma_{0}^{u},\sigma_{1}^{u}),\widehat{y}\right)\big{]},$
where $L((\sigma_{0},\sigma_{1}),y)=-(1-y)\log P(\sigma_{0}\succ\sigma_{1};{\psi})-y\log P(\sigma_{1}\succ\sigma_{0};{\psi})$. However, the pretrained reward model may generate incorrect pseudo-labels for unlabeled dataset, leading to noisy training and poor generalization [31]. Therefore, one key question is how to design a data selecting mechanism to improve prediction accuracy and guarantee training stability.
The another key aspect is the theory for offline PbRL. There are very few algorithms specifically designed for offline PbRL with strong guarantee, including generalization bound for reward model and safe improvement guarantee for policy learning. Therefore, another key question is to develop a systematic theory for offline PbRL, including generalization bound and improvement guarantee.
V Reward Learning
The reward learning involves two stages: pretraining based on collected labeled data $\mathcal{D}_{l}$ and updating based on $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$ during policy learning, which is essentially semi-supervised learning [26]. This section focuses on designing a data selecting function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of generated preference data, and explaining the factors that influence the generalization ability of reward model.
V-A Uncertainty-aware Pseudo-label Selection
Motivated by the previous pseudo-labeling work [31], we select data from unlabeled dataset $\mathcal{D}_{u}$ according two principles: confidence and uncertainty. The data with high confidence and low uncertainty can be chosen for reward training. High confidence refers that pre-trained reward model discriminates the preference of two trajectories into $\widehat{y}$ with high probability $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$. Low uncertainty refers that the $N_{R}$ reward models (model ensembles) predicts with small variance $\tau(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})$. The probability confidence $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$ and uncertainty variance $\tau(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})$ can be denoted as
$p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})=(1-\widehat{y})\bar{P}_{\psi}\left(\sigma_{0}^{u}\succ\sigma_{1}^{u}\right)+\widehat{y}\bar{P}_{\psi}\left(\sigma_{1}^{u}\succ\sigma_{0}^{u}\right)$ (7)
$\tau(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})=\text{{{Std} }}\big{\{}\bm{P}_{\psi}\left(\sigma_{0}^{u}\succ\sigma_{1}^{u}\right)\big{\}},$
where $N_{R}$ is the number of reward model, Std $\{\bm{P}(\cdot)\}$ denotes the variance of output probability of $N_{R}$ reward models, and $\bar{P}(\cdot)$ is the mean probability of $N_{R}$ reward models. Therefore, according to high confidence and low uncertainty, the $f(\sigma_{0}^{u},\sigma_{1}^{u})$ can be denoted as:
$f(\sigma_{0}^{u},\sigma_{1}^{u})=\mathbbm{1}\big{[}p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})>\kappa_{p}\big{]}\cdot\mathbbm{1}\big{[}\tau(\sigma_{0}^{u},\sigma_{1}^{u})<\kappa_{\tau}\big{]},$ (8)
where $\kappa_{p}$ and $\kappa_{\tau}$ are thresholds of confidence and uncertainty respectively. $f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes the generated data $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected, and $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ denotes the data is not selected. Then combing Eq. (6), the reward model can be optimized through minimizing
^$\widehat{\mathcal{L}}_{R}(\psi)=\frac{1}{\widetilde{N}_{u}}\sum\nolimits_{j=1}^{{N}_{u}}f\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)^{(j)}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$ (9)
$\frac{1}{N_{l}}\sum\nolimits_{i=1}^{N_{l}}L\left(\sigma_{0}^{l},\sigma_{1}^{l})^{(i)},y^{i}\right),$
where $\widetilde{N}_{u}$ is the number of generated dataset after screening, and ^$\widehat{\mathcal{L}}_{R}(\psi)$ is the empirical risk. It contains two parts: labeled loss ^$\widehat{\mathcal{L}}_{l}(\psi)$ and unlabeled loss ^$\widehat{\mathcal{L}}_{u}(\psi)$. Note that the label in unlabeled loss is pseudo-label, which may be different from the true label. Therefore, there is the gap between the unlabeled loss with pseudo-label ^$\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$. Before bounding this gap, we firstly give the below assumption without loss of generality.
Assumption 2.
For the pretrained reward model $\widehat{R}_{\psi}$ through the limited labeled dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$, if the pseudo label $\widehat{y}$ is defined in Eq. (5), then the pseudo-labeling error for the unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(j)}\}_{j=1}^{N_{u}}$ is smaller than $\eta$, i.e., $\sum_{j=1}^{N_{u}}\mathbbm{1}[\widehat{y}^{j}\neq y^{j}]/N_{u}\leq\eta$.
Proposition 1.
Suppose that the loss $L((\sigma_{0},\sigma_{1}),y))$ is bounded by $\Omega$. Then, for any $R\in\mathcal{R}$, under Assumption 2 the following equation holds:
|^$\big{|}\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\big{|}\leq\eta\Omega.$ (10)
The proof of Proposition 1 can be found in Appendix A-A. This gap is mainly influenced by the term $\eta$. Through selecting high confidence and low uncertainty samples, the $\eta$ can be significantly reduced. Combining Eqs. (5) and (8), we find that if $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$ is small, the $L((\sigma_{0}^{u},\sigma_{1}^{u}),\widehat{y})$ would become larger. Too large error can lead to unstable training and low performance. Therefore, through selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, a more accurate subset of pseudo-labels can be used in reward training, and through setting a higher confidence threshold for pseudo labeling, a lower prediction error and the stability of training can be guaranteed.
V-B Generalization Bound for Reward Model
Before developing generalization bound, we define the expected error $\mathcal{L}_{R}(\psi)$ with respect to the reward model $R(s,a;\psi)$: $\mathcal{L}_{R}(\psi)=\mathbb{E}_{(\sigma_{0},\sigma_{1},y)\sim\mathcal{D}}[L((\sigma_{0},\sigma_{1}),y)]$, where $\mathcal{D}$ is data distribution. The reward model is trained through minimizing the empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)$ in Eq. (9). Through developing generalization error bound, we can analyse the factors that influence generalization ability.
Similar to the previous generalization bound works [29], Rademacher complexity, which measures the richness of a certain hypothesis space [19], is introduced firstly. The definition is given below:
Definition 1 (Empirical Rademacher complexity).
Let $\mathcal{F}$ be a family of functions mapping from $\mathcal{Z}$ to $\mathbb{R}$ and $\widehat{\mathcal{S}}=\left\{z_{1},\ldots,z_{N_{s}}\right\}$ be a fixed sample of size $N_{s}$ drawn from the distribution $\mathcal{S}$ over $\mathcal{Z}$. The empirical Rademacher complexity of $\mathcal{G}$ for sample $\widehat{\mathcal{S}}$ is defined as
^$\widehat{\Re}_{\widehat{\mathcal{S}}}(\mathcal{F})=\mathbb{E}_{\bm{\sigma}}\big{[}\sup_{f\in\mathcal{F}}\frac{1}{N_{s}}\sum\nolimits_{i=1}^{N_{s}}\sigma_{i}f\left(z_{i}\right)\big{]},$ (11)
where $\bm{\sigma}=(\sigma_{1},...,\sigma_{N})$ are independent uniform random variables taking values in $\{-1,+1\}$.
Definition 2 (Induced Reward Function Families).
Given for a space $\mathcal{R}$ of reward function $R$, the induced reward function families $\Pi(\mathcal{F})$ is defined as
$\Pi(\mathcal{R})=\big{\{}(\sigma_{0},\sigma_{1})\rightarrow-\log P(\sigma_{0}\succ\sigma_{1})~{}|~{}R\in\mathcal{R}\big{\}},$ (12)
where $P(\cdot)$ is defined in Eq. (2) and $\Pi(\mathcal{R})$ is the union of projections of $\mathcal{R}$ onto each dimension.
In general, lower Rademacher complexity corresponds to better generalization performance. Then, based on the above definition and Proposition 1, we develop the new generalization bound for reward model trained through Eq. (9). The specific theorem is given below.
Theorem 1.
Let reward model be trained on the labeled dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$ and unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$. Then, for any $\delta>0$, with probability at least $1-\delta$, under Assumption 1 and 2 the following holds for any reward function $R\in\mathcal{R}$,
$\mathcal{L}_{R}(\psi)\leq\widehat{\mathcal{L}}_{R}(\psi)+\eta\Omega+4\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}\big{(}\Pi(\mathcal{R})\big{)}+3\sqrt{\frac{\log(2/\delta)}{2(N_{l}+N_{u})}},$
where $\widehat{\mathcal{D}}$ is the input combination of labeled and unlabeled dataset, i.e. $\{(\sigma_{0},\sigma_{1})^{(i)}\}_{i=1}^{N_{l}+N_{u}}$.
The proof of Theorem 1 can be found in Appendix A-B. Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)$, pseudo-labeling error $\eta\Omega$, Rademacher complexity and constant terms. The constant term can become small when the number of unlabeled data increases, but it may cause unstable training when labels are inaccurate. According to Proposition 1, empirical error and pseudo-labeling error can be reduced through selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, thus the better generalization ability, that is tighter upper bound, can be achieved. Note that Theorem 1 has generality and can be applicable to methods that train reward model using pseudo-labels. The difference among various methods may lie in how they train reward model to improve the accuracy of pseudo-labels, that is reducing $\eta$. If one method fails to reduce the pseudo-labeling error $\eta$, the upper generalization bound would be looser than that of no data augmentation.
VI Policy Learning
This paper directly uses the proposed offline RL algorithms, such as CQL [4] and IQL [25], to perform policy learning. To solve the problem of high cost for labeling preference, we propose a offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE). This section aims to describe the details of LEASE and establish the theory for safe policy improvement guarantee.
VI-A The Implementation Details of LEASE
1
20:  offline dataset $\mathcal{D}_{\text{offline}}$, limited labeled preference dataset $\mathcal{D}_{l}$, critic $Q_{\omega}$, policy $\pi_{\theta}$, transition model ${T}_{\varphi}$, and reward model $R_{\psi}$.
30:  reward $\widehat{R}$ and policy $\widehat{\pi}$ learned by LEASE.
41:  Initialization: Randomly initializing all networks and generated unlabeled dataset $\mathcal{D}_{u}=\varnothing$.
2:  Train an ensemble transition model $\{{T}_{\varphi}^{i}\}_{i=1}^{N_{T}}$ on $\mathcal{D}_{\text{offline}}$ according to Eq. (4).
3:  Pre-train an ensemble reward model $\{{R}_{\psi}^{i}\}_{i=1}^{N_{R}}$ on $\mathcal{D}_{l}$ according to Eq. (3).
4:  for $t=1,2,\cdots,N_{\text{iter}}$ do
5:     if $t$ rollout frequency $=0$ then
6:        Generate synthetic $H$-step rollouts by ${T}_{\varphi}$. Add transition trajectory into $\mathcal{D}_{u}$.
7:     end if
8:     if reward update condition is TRUE then
9:        Give pseudo label for unlabeled dataset $\mathcal{D}_{u}$ based on pre-trained reward model.
10:        Select high confidence and low uncertainty unlabeled data according to Eq. (8).
11:        Train the reward model based on $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ according to Eq. (9).
12:     end if
13:     Train policy using offline RL algorithms, such as CQL and IQL.
14:  end for
Algorithm 1 LEASE: offLine prEference-bAsed RL with high Sample Efficiency
Most offline RL algorithms are based on Eq. (1). They proposed various algorithms to solve the problem of distribution shift. In offline PbRL setting, the reward is unknown for agent. The accuracy of reward model directly influence the performance of policy. Moreover, the unlabeled data are generated through the current policy interaction with learned transition model. Therefore, the reward model and policy influence each other. In practical implementation, the policy, $Q$-function, reward and transition model are parameterized by $\pi_{\theta}$, $Q_{\omega}$, $R_{\psi}$ and $T_{\varphi}$ respectively.
Algorithm 1 gives pseudocode for LEASE. The goal of LEASE is to learn better reward model $R_{\psi}$ from fewer labeled preference dataset $\mathcal{D}_{l}$ and learn better policy $\pi_{\theta}$ from offline dataset $\mathcal{D}_{\text{offline}}$ without interaction with environment. The key of LEASE is how to train reward model in policy learning process. Notably, the reward model only update once instead of updating constantly in this process. The reward update condition is set to when the number of unlabeled data in buffer reaches the maximum buffer capacity. The update time is influenced by rollout frequency and rollout batch size. The reward update condition, rollout length $H$ and selecting mechanism $f(\sigma_{1},\sigma_{2})$ are the important factors for reward model.
VI-B Safe Policy Improvement Guarantee
Offline RL aims to guarantee $\xi$-safe policy improvement over the behavior policy $\mu$ (the policy used to collect offline dataset), i.e. $J(\widehat{\pi},R^{*})\geq J(\mu,R^{*})-\xi$ [4]. This part develops theoretical guarantee of policy improvement for LEASE, that is giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$, and further analyzes sample complexity when reward model is unknown. Firstly, we propose a new single-policy concentrability coefficient for PbRL based on state-action pair instead of trajectory like [18].
Definition 3 (Concentrability coefficient for PbRL).
Concentrability coefficient $\mathscr{C}_{\mathcal{R}}(\pi)$ is used to measure how well reward model errors transfer between offline data distribution $d_{T}^{\mu}$ and the visitation distribution $d_{T}^{\pi}$ under transition $T$ and policy $\pi$, defined as
$\mathscr{C}_{\mathcal{R}}(\pi)=\sup_{\widehat{R}\in\mathcal{R}}\Bigg{|}\frac{\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}}{\mathbb{E}_{(s,a)\sim d_{T}^{\mu}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}}\Bigg{|}.$ (13)
where $R^{*}(s,a)$ is the true reward model and coefficient $\mathscr{C}_{\mathcal{R}}(\pi)$ is upper bounded by $\|d_{T}^{\pi}/d_{T}^{\mu}\|_{\infty}$.
Next, similar to [20, 18], we use $\varepsilon$-bracketing number to measure the complexity of reward function class $\mathcal{R}$, which can be defined as
Definition 4 ($\varepsilon$-bracketing number).
The $\varepsilon$-bracketing number $\mathcal{N}_{\mathcal{R}}(\varepsilon)$ is the minimum number of $\varepsilon$-brackets $(l,u)$ required to cover a function class $\mathcal{R}$, where each bracket $(l,u)$ satisfies $l(\sigma_{0},\sigma_{1})\leq P_{R}(\sigma_{0},\sigma_{1})\leq u(\sigma_{0},\sigma_{1})$ and $\|l-u\|_{1}\leq\varepsilon$ for all $R\in\mathcal{R}$ and all trajectory-pairs $(\sigma_{0},\sigma_{1})$, and $P_{R}(\sigma_{0},\sigma_{1})$ is the probability that segment $\sigma_{0}$ is preferable to segment $\sigma_{1}$ defined in Eq. (2).
Theorem 2.
Under Assumption 1, for any $\delta\in(0,1]$, the policy $\widehat{\pi}$ learned by LEASE, with high probability $1-\delta$, satisfies that $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$ is upper bound by
$\xi+\frac{1+\mathscr{C}_{\mathcal{R}}(\widehat{\pi})}{1-\gamma}\Bigg{(}\sqrt{\frac{4C}{NL^{2}}\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}$ (14)
$+{\color[rgb]{0,0,0}\sqrt{\frac{4R_{max}^{2}\log\big{(}1/\delta\big{)}}{NL}}}\Bigg{)},$
where $C>0$ is the absolute constant defined in Eq. (2), $N$ is the size of preference dataset, and $L$ is the length of trajectory. The term $\xi$ is performance gap depending on offline algorithm itself.
The proof of Theorem 2 can be found in Appendix A-C. The $\xi$ is constant when offline algorithm is determined. Therefore, the tighter bound can be achieved through reducing performance gap caused by reward model (see Proposition 3). It can be reduced to small value $\varpi$ with sample complexity of
$N=\widetilde{\mathcal{O}}\Bigg{(}\frac{4\big{(}1+\mathscr{C}_{\mathcal{R}}(\widehat{\pi})\big{)}^{2}}{\varpi^{2}(1-\gamma)^{2}}\Bigg{(}$ $\sqrt{\frac{C\log(\mathcal{N}_{\mathcal{R}}(1/N)/\delta)}{L^{2}}}$ (15)
$+\sqrt{\frac{R_{max}^{2}\log\big{(}1/\delta\big{)}}{L}}\Bigg{)}^{2}\Bigg{)},$
where concentrability coefficient $\mathscr{C}_{\mathcal{R}}(\widehat{\pi})$ can become smaller when the learned policy $\widehat{\pi}$ is close to behavior policy $\mu$. Bracketing number $\log(\mathcal{N}_{\mathcal{R}}(1/N))$ measures the complexity of function class $\mathcal{R}$ and takes $\widetilde{\mathcal{O}}(d)$ in linear reward model [18]. Notably, LEASE can learn accurate reward model through data augmentation under fewer $N_{l}$ preference data ($N_{l}<N$). Thus, the performance gap can become tighter under fewer data, that is LEASE can have higher sample efficiency compared with [18]. More discussion can be found in Section VIII-A.
VII Experiments
This section gives the details of experiments and focuses on answering the following questions:
  • $\mathbf{Q_{1}}$: How well the accuracy of reward model through data augmentation under fewer preference data?
  • $\mathbf{Q_{2}}$: How well does LEASE perform compared with offline PbRL baseline in standard benchmark tasks?
  • $\mathbf{Q_{3}}$: How the proposed selecting mechanism and parameter settings affect the performance of reward and policy?
We answer these questions on two distinct domains (mujoco and adroit) within the D4RL benchmark [22], covering several locomotion and manipulation tasks. The code for LEASE is available at anonymous.link.
VII-A Experiment Setup
Figure 2: The description of Mujoco tasks (locomotion tasks) and Adroit tasks (manipulation tasks).
Experimental Datasets. The offline dataset and preference dataset of experiments originate from [22] and [21], respectively. The offline datasets are based on the D4RL dataset [22]. We select medium and medium-expert two types dataset for Mujoco tasks, and human and expert for Adroit tasks. The difference between different dataset in certain task lies in the collected policy. Fig. 2 depicts the six tasks for D4RL benchmark. For three mujoco tasks, the goal is to control the robot’s various joints to achieve faster and more stable locomotion. For adroit tasks, they involve controlling a 24-DoF simulated Shadow Hand robot on a sparse reward, high-dimensional robotic manipulation task.
The preference dataset are from the previous work [21], where they provided baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF). One is crowd-sourced labels obtained by crowd-sourcing, and the other is synthetic labels based on ground truth reward. To test LEASE performance for two types preference data, the labels of preference is from human-realistic feedback for mujoco tasks, and ground truth reward for adroit tasks. This paper aims to test the performance under fewer preference data without online interaction. We take the first $100$ data of [21] as $2000$ preference dataset, and denote method without data augmentation as FEWER.
TABLE I: Base parameter configuration of method LEASE.
Model Parameter Value
Transition Model Model learning rate $1\times e^{-3}$
Number of hidden layers 4
Number of hidden units per layer 200
Batch size 256
Number of model networks $N_{T}$ 7
Number of elites 5
Reward Model Model learning rate $3\times e^{-4}$
Number of hidden layers 3
Number of hidden units per layer 256
Batch size (Pretrain) 256(64)
Number of model networks $N_{R}$ 3
Number of labeled dataset $N_{l}$ 100
Policy Model Actor learning rate (IQL) $1(3)\times e^{-4}$
Critic learning rate $3\times e^{-4}$
Number of hidden layers(IQL) 3(2)
Number of hidden units per layer 256
Batch size 256
Experimental Details. Offline PbRL setting involves two steps: reward learning and policy learning. Here, we directly use offline RL algorithms CQL [4] and IQL [25] to perform policy learning. The performance of these algorithms is evaluated based on the cumulative return. For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [22]. The normalized score $\tilde{J}$ is computed by:
$\tilde{J}=\frac{J-J_{r}}{J_{e}-J_{r}}\times 100,$ (16)
where $J_{r}$, $J_{e}$ and $J$ are the expected return of a random policy, an expert policy, and the trained policy by offline RL algorithms, respectively. Table I gives the hyper-parameter configuration for reward and transition training, and policy learning. The details are given the below:
Figure 3: The relationship between preference model accuracy and the number of preference dataset $N_{l}$.
TABLE II: The hyperparameters for LEASE under CQL and IQL algorithms.
Task Name CQL IQL
$H$ $\kappa_{p}$ $\kappa_{\tau}$ $H$ $\kappa_{p}$ $\kappa_{\tau}$
walker2d-m $10$ $0.85$ $0.05$ $10$ $0.75$ $0.08$
walker2d-m-e $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
hopper-m $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
hopper-m-e $100$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
halfcheetah-m $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
halfcheetah-m-e $200$ $0.85$ $0.05$ $10$ $0.75$ $0.08$
pen-human $10$ $0.95$ $0.05$ $10$ $0.85$ $0.08$
pen-expert $10$ $0.90$ $0.05$ $10$ $0.95$ $0.08$
door-human $10$ $0.95$ $0.05$ $10$ $0.90$ $0.08$
door-expert $20$ $0.95$ $0.05$ $10$ $0.95$ $0.08$
hammer-human $10$ $0.95$ $0.05$ $10$ $0.90$ $0.08$
hammer-expert $20$ $0.99$ $0.05$ $10$ $0.95$ $0.08$
1) Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models. Each model consists of a $4$-layer feed-forward neural network with $200$ hidden units. The model training employs maximum likelihood estimation with a learning rate of $1\times e^{-3}$ and Adam optimizer. Following [21], we train an ensemble of $3$ reward models. Each reward model includes a $3$-layer feed-forward neural network with $256$ hidden units. The reward model training uses cross-entropy loss with a learning rate of $3\times e^{-4}$ and Adam optimizer.
2) Policy optimization: The offline algorithms include CQL and IQL. In CQL algorithm, the actor network $\pi_{\theta}$ and critic network $Q_{\omega}$ adopt a $3$-layer feed-forward neural network with learning rate of $1\times e^{-4}$ and $3\times e^{-4}$, respectively. The cql weight and temperature are set as $5.0$ and $1.0$ for all tasks, respectively. In IQL algorithm, the actor network and critic network adopt a $2$-layer neural network with learning rate of $3\times e^{-4}$ and $3\times e^{-4}$, respectively. The expectile and temperature are set as $0.7$ and $3.0$ for mujoco tasks and $0.8$ and $3.0$ for adroit tasks.
3) Parameter tuning: Theorem 1 indicates that the performance of reward model is influenced by the selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$ and the quality of the generated data. Therefore, the key hyperparameters should be tuned for LEASE including: rollout horizon $H\in\{10,20,100,200\}$, probability confidence $\kappa_{p}(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})\in\{0.75,0.85,0.90,0.95,0.99\}$ and uncertainty variance $\kappa_{\tau}(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})\in\{0.05,0.08\}$. Table II shows the above hyperparameters for LEASE under CQL and IQL algorithms.
Figure 4: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) hopper-medium-expert and (b) halfcheetah-medium-expert datasets, where the value predicted by the trained reward model and ground-truth reward value are both normalized to $[0,1]$. From left to right are methods LEASE, FEWER and FRESH, respectively.
VII-B Results for Reward Model
The effect of the number of preference data for preference model. Learning an accurate reward is difficult in the offline PbRL setting when preference labels are scarce. Fig. 3 illustrates the relationship between the prediction accuracy of preference model without data augmentation and the number of preference data $N_{l}$, where preference model is based on reward model (Eq. (2)). Here, we use the last $200$ of $2000$ preference data that are not seen in training stage as evaluation dataset. It indicates that as the number of preference data $N_{l}$ increases, the prediction accuracy of the preference model shows an upward trend. This also validates Theorem 1 that increasing the number of samples $N_{l}$ can reduce the generalization error.
TABLE III: Comparison results of pseudo-labeling accuracy and agent performance between using confidence and uncertainty and only using confidence. The accuracy of pseudo-label generated by reward model is evaluated on all preference dataset.
Selecting Mechanism pen-expert door-expert hammer-expert
label accuracy agent performance label accuracy agent performance label accuracy agent performance
Confidence and uncertainty $87.3\pm 0.5$ $\mathbf{132.5\pm 2.3}$ $89.3\pm 0.6$ $\mathbf{103.2\pm 0.7}$ $85.5\pm 0.2$ $\mathbf{126.3\pm 1.2}$
Only confidence $85.9\pm 0.8$ $126.8\pm 0.1$ $87.8\pm 0.2$ $101.1\pm 2.2$ $84.4\pm 0.4$ $118.2\pm 9.6$
The effect of data augmentation for the performance of reward model. We compare the accuracy of reward model before and after data augmentation under fewer preference dataset. Fig. 4 shows the comparison results between prediction and ground truth of reward, where the predicted and true rewards are both normalized to $[0,1]$. We randomly sample $500$ data from unlabeled datasets that are not seen in training stage for evaluation. This figure indicates the linear relationship between reward predicted by LEASE and ground truth is better than that of other two methods, where method not using selecting mechanism is denoted as FRESH. The prediction of reward model learned by method FRESH is very narrow and the accuracy is greatly reduced compared with FEWER. This is because generated data of FRESH are not selected by $f(\sigma_{0},\sigma_{1})$, which causes substantial errors for the labels of generated data and leads to the collapse of training. This validates that the performance of reward model can be improved through data augmentation and selecting mechanism, where the designed selecting mechanism has a greater impact on reward model performance.
TABLE IV: The performance of the learned transition model, where error is calculated by the sum of the mean square values of predicted and true value in each dimension.
Task Name walker2d-m-e hopper-m-e halfcheetah-m-e
Error $6.57\pm 0.10$ $0.38\pm 0.01$ $16.19\pm 0.48$
Task Name pen-expert door-expert hammer-expert
Error $1.05\pm 0.01$ $0.04\pm 0.01$ $0.58\pm 0.03$
The effect of uncertainty for reward model and agent performance. Table III compares the accuracy of pseudo-labels and performance between the proposed method LEASE and the confidence-only method. The latter method is the Surf [3] that only uses confidence for screening data. This table validates that LEASE outperforms the confidence-only approach in pseudo-labels accuracy and agent performance. The performance improved by $4.5\%$, $2.8\%$ and $6.85\%$ for pen-expert, door-expert and hammer-expert, respectively. Additionally, there is a difference between the accuracy of the preference model $P(\sigma_{0}\succ\sigma_{1})$ (pseudo-label accuracy) and the accuracy of the reward model $R(s,a)$ (See Eq. (2)). A decline in the accuracy of the reward model may not necessarily lead to a significant drop in the accuracy of the preference model, as the preference model is related to the summation of reward model over trajectories.
VII-C Results for Transition model
The accuracy of transition model. Table IV shows the accuracy result for the learned transition model, where the accuracy of transition model is measured by the sum of the mean square values of the predicted value and the true value in each dimension. Note that, for walker2d and halfcheetah tasks, the error of the learned transition model is greatly larger than hopper tasks in mujoco environment. This is mainly because the state space of walker2d and hopper is larger than hopper, and the physical model of them are more complex.
TABLE V: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds. $\pm$ captures the standard deviation over seeds. Bold indicates the performance within 2% of the best performing algorithm for each offline algorithm.
Task Name CQL [4] IQL [25]
URLHF FEWER LEASE URLHF FEWER LEASE
walker2d-m $76.0\pm 0.9$ $77.4\pm 0.6$ $\mathbf{78.4\pm 0.9}$ $\mathbf{78.4\pm 0.5}$ $71.8\pm 0.8$ $74.6\pm 1.8$
walker2d-m-e $92.8\pm 22.4$ $77.7\pm 0.3$ $\mathbf{98.6\pm 18.1}$ $\mathbf{109.4\pm 0.1}$ $105.2\pm 3.1$ $\mathbf{108.1\pm 0.5}$
hopper-m $54.7\pm 3.4$ $55.8\pm 2.8$ $\mathbf{56.5\pm 0.6}$ $50.8\pm 4.2$ $\mathbf{56.7\pm 2.6}$ $\mathbf{56.0\pm 0.5}$
hopper-m-e $\mathbf{57.4\pm 4.9}$ $53.6\pm 0.9$ $56.4\pm 0.8$ $\mathbf{94.3\pm 7.4}$ $56.0\pm 1.4$ $55.9\pm 1.9$
halfcheetah-m $\mathbf{43.4\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.3\pm 0.2}$ $42.2\pm 0.1$ $\mathbf{43.0\pm 0.3}$
halfcheetah-m-e $\mathbf{62.7\pm 7.1}$ $48.3\pm 0.7$ $53.2\pm 3.1$ $\mathbf{91.0\pm 2.3}$ $59.1\pm 4.9$ $62.4\pm 1.4$
Mujoco Average $\mathbf{64.5\pm 6.5}$ $59.4\pm 0.9$ $\mathbf{64.4\pm 4.0}$ $\mathbf{77.9\pm 2.5}$ $65.2\pm 2.2$ $66.7\pm 1.0$
pen-human $\mathbf{9.8\pm 14.1}$ $0.5\pm 3.0$ $3.8\pm 4.6$ $50.2\pm 15.8$ $67.3\pm 10.0$ $\mathbf{75.6\pm 3.3}$
pen-expert $\mathbf{138.3\pm 5.2}$ $128.1\pm 0.7$ $132.5\pm 2.3$ $\mathbf{132.9\pm 4.6}$ $104.1\pm 12.9$ $113.8\pm 6.3$
door-human $\mathbf{4.7\pm 5.9}$ $0.2\pm 1.0$ $\mathbf{4.7\pm 8.8}$ $3.5\pm 3.2$ $4.0\pm 2.5$ $\mathbf{5.9\pm 0.5}$
door-expert $\mathbf{103.9\pm 0.8}$ $103.0\pm 0.9$ $\mathbf{103.2\pm 0.7}$ $\mathbf{105.4\pm 0.4}$ $104.5\pm 0.6$ $\mathbf{105.2\pm 0.2}$
hammer-human $\mathbf{0.9\pm 0.3}$ $0.3\pm 0.0$ $0.3\pm 0.0$ $1.4\pm 1.0$ $1.2\pm 0.2$ $\mathbf{1.7\pm 0.4}$
hammer-expert $120.2\pm 6.8$ $124.1\pm 2.1$ $\mathbf{126.3\pm 1.2}$ $\mathbf{127.4\pm 0.2}$ $125.2\pm 2.3$ $\mathbf{126.3\pm 0.1}$
Adroit Average $\mathbf{63.0\pm 5.5}$ $59.4\pm 1.3$ $61.8\pm 3.0$ $\mathbf{70.1\pm 4.2}$ $67.1\pm 4.8$ $\mathbf{71.4\pm 1.8}$
TABLE VI: The comparison results between the performance using selecting mechanism and that not using. The latter method is denoted as FRESH. $\uparrow$ denotes the improvement of performance.
Task Name walker2d-m hopper-m halfcheetah-m walker2d-m-e hopper-m-e halfcheetah-m-e
CQL FRESH $76.0\pm 1.3$ $54.4\pm 1.6$ $43.4\pm 0.1$ $77.4\pm 2.2$ $55.0\pm 0.4$ $49.9\pm 0.9$
LEASE $\mathbf{78.4\uparrow(3.3\%)}$ $\mathbf{56.5(\uparrow 4.0\%)}$ $\mathbf{43.5\uparrow(0.6\%)}$ $\mathbf{98.6\uparrow(27.4\%)}$ $\mathbf{56.4\uparrow(2.6\%)}$ $\mathbf{53.2\uparrow(6.6\%)}$
IQL FRESH $72.7\pm 4.1$ $53.5\pm 0.8$ $42.5\pm 0.1$ $105.3\pm 1.0$ $53.9\pm 0.1$ $54.5\pm 2.2$
LEASE $\mathbf{74.6\uparrow(2.6\%)}$ $\mathbf{56.0\uparrow(4.5\%)}$ $\mathbf{43.0\uparrow(1.3\%)}$ $\mathbf{108.1\uparrow(2.7\%)}$ $\mathbf{55.9\uparrow(3.7\%)}$ $\mathbf{62.4\uparrow(14.6\%)}$
TABLE VII: The prediction result of preference model for different rollout length $H$. Pretrained represents the accuracy of preference model trained with fewer preference data, and updated represents that updated with generated data.
Task name walker2d-m-e hopper-m-e halfcheetah-m-e
Pretrained $0.85$ $0.77$ $0.72$
Updated $H=10$ $\mathbf{0.85}$ $0.74$ $\mathbf{0.78}$
$H=100$ $0.84$ $\mathbf{0.79}$ $0.74$
$H=200$ $0.83$ $0.76$ $0.77$
The effect of rollout length $H$ of transition model for preference model. Table VII shows the prediction accuracy of the trained preference model for different rollout length $H$. Here, we take CQL as example and select the last $200$ of $2000$ preference data as evaluation dataset that are not seen in in reward training stage. The accuracy is calculated through the gap between predicted preference and true preference based on two trajectories. It shows that long horizon easily brings low performance of reward model since the cumulative error of the trained transition model, but long horizon is beneficial to enhance the performance of reward model under some conditions. The accuracy of preference model can be improved through data augmentation under most tasks.
Note that the rollout length $H$ is equal to the length of generated unlabeled data $L$. The $H$ is influenced by the accuracy of trained transition model. As $H$ increases, the accuracy of prediction decreases and large error will bring the unstable agent training. However, the long horizon of trajectory can better describe the preference of human feedback and can reduce sample complexity (See Eq. (15)). Additionally, the results in table VII can not completely reflect the generalization performance of reward model since the evaluation data are limited and there is difference between preference model and reward model (See Eq. (2)).
TABLE VIII: Results of the effect of the learned transition model for agent performance.
No. hopper-medium pen-expert
Error Performance Error Performance
$1$ $0.35\pm 0.01$ $56.5\pm 0.60$ $1.05\pm 0.01$ $132.5\pm 2.3$
$2$ $0.49\pm 0.04$ $54.8\pm 0.85$ $1.42\pm 0.05$ $126.4\pm 8.6$
$3$ $1.19\pm 0.05$ $52.9\pm 0.92$ $2.36\pm 0.24$ $87.7\pm 41.8$
The effect of the accuracy of transition model for agent performance. Table VIII shows the detailed results for the effect of transition model accuracy for agent performance, where the smaller the model error means the higher the accuracy. The change in transition model error is achieved by changing the number of training steps. The base offline RL algorithm is CQL. It shows that the lower the accuracy of the transition model, the poorer performance of the agent.
VII-D Results for Policy Performance
The comparison results on benchmark tasks. Here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms. Table V gives the results of the average normalized score with standard deviation, where m and m-e denote medium and medium-expert respectively. The offline PbRL algorithm under sufficient feedback (URLHF) are obtained from paper [21]. This table shows that LEASE can greatly improve performance under fewer preference data compared with FEWER, and achieve comparable performance to URLHF that demand large human feedback. This validates the content of Theorem 2 that LEASE can reduce the performance gap caused by reward model under fewer preference data and has high sample efficiency.
TABLE IX: The comparison results for the D4RL tasks under different number of preference data $N_{l}$. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score.
CQL [4] IQL [25]
Task Name $N_{l}=20$ $N_{l}=100$ Task Name $N_{l}=20$ $N_{l}=100$
walker2d-medium $77.7\pm 1.3$ $\mathbf{78.4\pm 0.9}$ pen-human $71.9\pm 7.4$ $\mathbf{75.6\pm 3.3}$
walker2d-medium-expert $98.0\pm 18.6$ $\mathbf{98.6\pm 18.1}$ pen-expert $102.5\pm 12.8$ $\mathbf{113.8\pm 6.3}$
hopper-medium $\mathbf{56.8\pm 1.5}$ ${56.5\pm 0.6}$ door-human $4.4\pm 1.2$ $\mathbf{5.9\pm 0.5}$
hopper-medium-expert $54.5\pm 1.2$ $\mathbf{56.4\pm 0.8}$ door-expert $\mathbf{105.2\pm 0.1}$ $\mathbf{105.2\pm 0.2}$
halfcheetah-medium $43.4\pm 0.4$ $\mathbf{43.5\pm 0.1}$ hammer-human $1.2\pm 0.4$ $\mathbf{1.7\pm 0.4}$
halfcheetah-medium-expert $51.0\pm 0.8$ $\mathbf{53.2\pm 3.1}$ hammer-expert $\mathbf{126.4\pm 0.1}$ $126.3\pm 0.1$
Mujoco Average $63.6\pm 4.0$ $\mathbf{64.4\pm 4.0}$ Adroit Average $68.6\pm 3.7$ $\mathbf{71.4\pm 1.8}$
The effect of selecting mechanism for agent performance. Table VI compares the agent performance between the proposed method LEASE and the method not using selecting mechanism (We denote this method as FRESH). This table indicates that the application of selecting mechanism $f(\sigma_{0},\sigma_{1})$ can efficiently improve agent performance. This results also further validates the effectiveness and advantages of the designed selecting mechanism.
The effect of the number of preference data $N_{l}$ for agent performance. Fig. 4 indicates that the number of preference data $N_{l}$ influences the accuracy of reward model directly. URLHF [21] trains reward model with $2000$ preference dataset. Our method LEASE aims to achieve comparable performance with URLHF under fewer preference dataset. Table IX shows the agent performance when the number of preference dataset is $20$ and $100$, where we use CQL algorithm for mujoco tasks and IQL algorithm for adroit tasks. This table indicates LEASE still can improve agent performance under very fewer preference dataset. Notably, the performance gap between $100$ preference data and $20$ preference data is not much. This indicates that LEASE have potential to perform well under more fewer preference dataset.
VIII Discussion
VIII-A Discussion for Offline PbRL Theory
In offline RL field, high sample efficiency refers that the agent can achieve comparable performance under fewer data compared with the performance under large data. In this paper, the data refers to the preference dataset. The labeled preference dataset, each trajectory of length $L$, is collected through real-time human feedback under policy $\mu$, which demands tremendous human effort, thus the collected cost of preference data is higher than fixed offline data. The unlabeled dateset is generated through trained transition without real-time interaction under learned policy $\pi^{t}$ at time $t$.
Through data augmentation, the sample efficiency can be significantly reduced. The performance gap caused by reward can be reduced to $\varpi$ under fewer labeled dataset. Zhan et al.[18] also developed systematical theory for offline PbRL and introduced the concentrability coefficient $\mathscr{C}_{r}(\pi)$ for PbRL, defined as
$\max\Bigg{\{}0,\sup_{r}\frac{\mathbb{E}_{\sigma^{0}\sim\pi,\sigma^{1}\sim\pi_{\text{ref}}}\big{[}r^{*}(\sigma^{0},\sigma^{1})-{r}(\sigma^{0},\sigma^{1})\big{]}}{\sqrt{\mathbb{E}_{\sigma^{0}\sim\mu_{0},\sigma^{1}\sim\mu_{1}}\big{|}r^{*}(\sigma^{0},\sigma^{1})-{r}(\sigma^{0},\sigma^{1})\big{|}^{2}}}\Bigg{\}},$
where $r(\sigma^{0},\sigma^{1})=\sum_{i}[R(s_{i}^{0},a_{i}^{0})-R(s_{i}^{1},a_{i}^{1})]$, $\pi_{\text{ref}}$ is an arbitrary trajectory distribution (usually set as $\mu_{1}$), and $\mu_{0},\mu_{1}$ are behavior trajectory distribution. However, it is based on trajectory and is difficult to combine with other offline RL theories. The proposed concentrability coefficient in our paper defined in Eq. (13) is based on state-action pairs.
Moreover, the performance gap of offline PbRL between behavior policy and learned policy is influenced by offline algorithm itself and the performance of the learned reward (preference) model. However, [18] fails to consider the gap caused by offline algorithm itself and the sampling error. The theory developed in our paper can be easily combined with other offline algorithm. The method in [18] can learn $\varpi$-optimal policy with a sample complexity of
$N=\widetilde{\mathcal{O}}\Bigg{(}\frac{c^{2}k^{2}\mathscr{C}_{r}^{2}(\widehat{\pi})}{\varpi^{2}}\log\left(\frac{\mathcal{N}_{r}(1/N)}{\delta}\right)\Bigg{)},$ (17)
where $c>0$ is a universal constant, $k=(\inf_{x\in[-r_{max},r_{max}]}\\ \Phi^{\prime}(x))^{-1}$, and $\Phi(x)$ is a monotonically increasing link function. Compared with Eq. (17), the sample complexity of LEASE defined in Eq. (15) contains more useful information. For example, the sample complexity is related to the length of preference data and the range of reward function, and can be reduced when the learned policy is closed to behavior policy.
VIII-B Discussion for Broader Impacts
Actual application. In actual scenarios, especially for human-in-loop-control, such as exoskeleton robot assistance or rehabilitation, designing a high sample efficient RL algorithm is greatly significant. Firstly, many rewards are difficult to describe in mathematical terms in some situation, such as the comfort of interaction. Human feedback or preference is the better way to reflect the above indexes. In addition, in human in-the-loop control, obtaining preference data through interaction can easily cause fatigue, and inappropriate interaction may cause damage to human. It is a necessity to learn reward model from limited preference dataset. Therefore, PbRL has a great potential to improve control performance in some scenarios where the reward function is difficult to describe and human is in the control loop.
Theoretical study. The theory of LEASE shows that the performance gap between the behavior policy $\mu$ and $\widehat{\pi}$ learned by LEASE includes two part: the gap $\xi$ caused by offline algorithm itself and the gap $\xi_{1}$ caused by reward model gap (the details see Theorem 2), where the term $\xi_{1}$ only depends on preference dataset. Moreover, the theory of LEASE is based on state-action pairs, which is consistent with most offline algorithms theory. Therefore, the theory developed in this paper can be easily combined with other offline algorithms theory and easily used to build theory of policy improvement guarantee, which can provide the theoretical basis for offline PbRL and facilitate the development of offline PbRL theory.
IX Conclusion
This paper proposes a novel offline PbRL algorithm (LEASE) with high sample efficiency. LEASE can achieve comparable performance under fewer preference dataset. By selecting high confidence and low variance data, the stability and accuracy of reward model are guaranteed. Moreover, this paper provides the theoretical analysis for LEASE, including generalization of reward model and policy improvement guarantee. This theory can be easily connected with other offline algorithms. The theoretical and experimental results demonstrate that the data selecting mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve performance of reward model and the performance learned by LEASE can be guaranteed under fewer preference dataset. However, there is still the gap between the true and the learned reward model. Future works can focus on how to further reduce this gap under the limited preference data or how to achieve conservative estimation for state-action pairs where the learned reward model predicts inaccurately.
Appendix A Related Proofs
A-A Proof of Proposition 1
The Proof for Proposition 1: This proof follows the previous work [29]. Obviously, the largest pseudo-labeling error $\eta$ in Assumption 2 includes exactly two types of pseudo-labeling error:
$\eta_{1}=\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]},$ (A.1)
$\eta_{2}=\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]},$
where $\eta_{1}$ represents the error ratio of classifying the first category as the second category, and $\eta_{2}$ is the error ratio of classifying the second category as the first category. Then, we prove the gap between the unlabeled loss with pseudo-label ^$\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ from two sides: ^$\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$ and ^$\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$. Notably, we ignore the selecting mechanism in the below proof since it doesn’t influence the proof result.
Step 1: Proving upper bound : ^$\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$.
^$\widehat{\mathcal{L}}_{u}(\psi)$
$=$ $\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$
$=$ $-\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5\big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$+\mathbbm{1}\big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5\big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
$\leq$ $-\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\big{[}y^{j}=0\big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$+\mathbbm{1}\big{[}y^{j}=1\big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
$+\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$+\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
$\leq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta_{1}\max_{j}\Big{\{}-\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}\Big{\}}$
$+\eta_{2}\max_{j}\Big{\{}-\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}\Big{\}}$
$\leq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega.$
Step 2: Proving low bound : ^$\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$.
^$\widehat{\mathcal{L}}_{u}(\psi)$
$=$ $\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$
$\geq$ $-\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\big{[}y^{j}=0\big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$+\mathbbm{1}\big{[}y^{j}=1\big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
$-\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$-\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
$\geq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega.$
Combing Step 1 and Step 2, we can obtain the following result:
|^$\Big{|}\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\Big{|}\leq\eta\Omega.$ (A.2)
This completes the proof of Theorem 1.
A-B Proof of Theorem 1
Lemma 1.
Let $\mathcal{F}$ be a family of functions mapping from $\mathcal{X}$ to $\mathbb{R}$ and $\widehat{\mathcal{D}}$ be empirical datasets sampled from an i.i.d. sample ${\mathcal{D}}$ of size $N$. Then, for any $\delta>0$, with probability at least $1-\delta$, the following holds for all $f\in\mathcal{F}$,
$\Big{|}\mathbb{E}_{\mathcal{D}}\left[f\right]-\mathbb{E}_{\widehat{\mathcal{D}}}\left[f\right]\Big{|}\leq 2\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}(\mathcal{F})+3\sqrt{\frac{\log(2/\delta)}{2N}},$ (A.3)
where $\mathbb{E}_{\widehat{\mathcal{D}}}\left[f\right]=\sum_{i=1}^{m}\left[g(x_{i})\right]/m$ is the empirical form of $\mathbb{E}_{\mathcal{D}}\left[f\right]$. This proof can be found in Theorem 3.3 of work [24].
Proposition 2.
Let $\mathcal{F}$ be a family of loss functions defined in Eq. (6) and $\Pi(\mathcal{R})$ be a family of functions defined in Definition 2. Then, for any sample $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the following relation holds between the empirical Rademacher complexities of $\Pi(\mathcal{R})$ and $\mathcal{F}$:
^$\widehat{\Re}_{\widehat{\mathcal{D}}}(\mathcal{F})\leq 2\widehat{\Re}_{\widehat{\mathcal{D}}}\big{(}\Pi(\mathcal{R})\big{)}.$ (A.4)
Proof. According to Definition 1, the empirical Rademacher complexity of $\mathcal{F}$ can be written as:
^$\widehat{\Re}_{\widehat{\mathcal{D}}}(\mathcal{F})$ (A.5)
$=$ $\mathbb{E}_{\bm{\sigma}}\bigg{[}\sup_{R\in\mathcal{R}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\big{[}-(1-y)\log P\left(\sigma_{0}\succ\sigma_{1}\right)$
$-y\log P\left(\sigma_{1}\succ\sigma_{0}\right)\big{]}\bigg{]}$
$\leq$ $\mathbb{E}_{\bm{\sigma}}\bigg{[}\sup_{R\in\mathcal{R}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\big{[}-\log P\left(\sigma_{0}\succ\sigma_{1}\right)\big{]}\bigg{]}$
$+\mathbb{E}_{\bm{\sigma}}\bigg{[}\sup_{R\in\mathcal{R}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\big{[}-\log P\left(\sigma_{1}\succ\sigma_{0}\right)\big{]}\bigg{]}$
$=$ $2\widehat{\Re}_{\widehat{\mathcal{D}}}\big{(}\Pi(\mathcal{R})\big{)}.$
This completes the proof of Proposition 2.
The Proof for Theorem 1: This proof is based on Proposition 1 and 2. Combing Eqs. (A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error ^$\widehat{\mathcal{L}}(\psi)$ holds:
$\mathcal{L}(\psi)\leq\widehat{\mathcal{L}}(\psi)+4\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}\big{(}\Pi(\mathcal{R})\big{)}+3\sqrt{\frac{\log(2/\delta)}{2N}}.$ (A.6)
However, the dataset for training reward model includes two parts: the labeled dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$ and unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$, and empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)=\widehat{\mathcal{L}}_{l}(\psi)+\widehat{\mathcal{L}}_{u}(\psi)$. Let ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ be the empirical error under unlabeled dataset with true label, then
$\mathcal{L}_{R}(\psi)\leq\widehat{\mathcal{L}}_{l}(\psi)+\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+4\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}\big{(}\Pi(\mathcal{R})\big{)}+3\sqrt{\frac{\log(2/\delta)}{2(N_{l}+N_{u})}},$
where $\widehat{\mathcal{D}}$ is the input combination of labeled and unlabeled dataset, denoted as $\{(\sigma_{0},\sigma_{1})^{(i)}\}_{i=1}^{N_{l}+N_{u}}$. According to Proposition 1, that is ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\leq\widehat{\mathcal{L}}_{u}(\psi)+\eta\Omega.$, we can derive
$\mathcal{L}_{R}(\psi)\leq\widehat{\mathcal{L}}_{R}(\psi)+\eta\Omega+4\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}\big{(}\Pi(\mathcal{R})\big{)}+3\sqrt{\frac{\log(2/\delta)}{2(N_{l}+N_{u})}}.$
This completes the proof of Theorem 1.
A-C Proof of Theorem 2
Before proof, we give the related Assumption and Lemmas for Theorem 2. Firstly, we assume the the reward class $\mathcal{R}$ is realizable (Assumption 1) and bounded (Assumption 3).
Assumption 3 (Boundedness).
For any $R\in\mathcal{R}$ and any state-action pairs $(s,a)$, the equation $|R(s,a)|\leq R_{max}$ holds.
Proposition 3.
Consider a set of trajectories $\{\sigma_{0}^{i},\sigma_{1}^{i}\}_{i=1}^{N}$, each of length $L$, collected from an offline dataset following the distribution $d_{T}^{\mu}(s,a)$. Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$, the following holds for all $\widehat{R}\in\mathcal{R}$
$\bigg{|}\mathbb{E}_{(s,a)\sim d_{T}^{\mu}(s,a)}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}\bigg{|}$ (A.7)
$\leq$ $\sqrt{\frac{4C}{NL^{2}}\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}+{\color[rgb]{0,0,0}\sqrt{\frac{4R_{max}^{2}\log\big{(}1/\delta\big{)}}{NL}}},$
where $R^{*}$ is the true reward model and $\widehat{R}$ is the learned reward model. $\mathcal{N}_{\mathcal{R}}(1/N)$ is the bracketing number defined in Definition 3.
The proof of Proposition 3 can be found in Appendix A-D. Note that, the above reward model is only trained from offline preference dataset. Through generating more preference data, the upper bound can be tighter. However, the generated data may be inaccuracy, which may improve the reward gap. The selecting mechanism can effectively solve this problem.
The Proof for Theorem 2: This Theorem aims to give the lower bound of term $J(\widehat{\pi},R^{*})-J(\mu,R^{*})$. We prove this from ideal case and actual case.
1) Ideal case: There is no need to consider distribution shift and empirical error problem. The policy $\widehat{\pi}$ is directly learned by $\widehat{\pi}=\max_{\pi}J(\pi,\widehat{R})$. Thus, $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})$ holds. Then, we can derive
$J(\mu,R^{*})-J(\widehat{\pi},R^{*})$ (A.8)
$=$ $J(\mu,R^{*})-J(\mu,\widehat{R})+J(\mu,\widehat{R})-J(\widehat{\pi},R^{*})$
$\leq$ $J(\mu,R^{*})-J(\mu,\widehat{R})+J(\widehat{\pi},\widehat{R})-J(\widehat{\pi},R^{*})$
$\leq$ $\big{|}J(\mu,R^{*})-J(\mu,\widehat{R})\big{|}+\big{|}J(\widehat{\pi},R^{*})-J(\widehat{\pi},\widehat{R})\big{|}.$
Since $J(\pi,R):=\mathbb{E}_{(s,a)\sim d_{T}^{\pi}(s,a)}[R(s,a)]/(1-\gamma)$, $|J(\pi,R^{*})-J(\pi,\widehat{R})|$ can be written as
$\big{|}J(\pi,R^{*})-J(\pi,\widehat{R})\big{|}=\frac{1}{1-\gamma}\Big{|}\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}\Big{|}.$
Then, according to Definition 3, we have
$\Big{|}J(\pi,R^{*})-J(\pi,\widehat{R})\Big{|}\leq\frac{\mathscr{C}_{\mathcal{R}}(\pi)}{1-\gamma}\Big{|}\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}\Big{|}.$
By Proposition 3, the following inequality holds
$J(\mu,R^{*})-J(\widehat{\pi},R^{*})\leq\frac{1+\mathscr{C}_{\mathcal{R}}(\widehat{\pi})}{1-\gamma}$ (A.9)
$\bigg{(}\sqrt{\frac{4C}{NL^{2}}\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}+{\color[rgb]{0,0,0}\sqrt{\frac{4R_{max}^{2}\log\big{(}1/\delta\big{)}}{NL}}}\bigg{)}.$
2) Actual case: Offline RL suffers from distribution shift problem and it is a necessity to consider empirical error. Therefore, many offline RL algorithms incorporate conservatism into policy to overcome distribution shift, that is learning policy through $\widehat{\pi}=\max_{\pi}J(\pi,\widehat{R})-P(\pi)$. Therefore, $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})-\xi$ instead of $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})$ in actual implantation. The term $\xi$ depends on the algorithm itself. Here, we take CQL as a example. The policy gap is proven in Theorem 3.6 of [4]. The detailed is given as follow.
$\xi=$ $\underbrace{\frac{\gamma C_{T,R}R_{max}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d_{T}^{\mu}(s)}\sqrt{\frac{|\mathcal{A}|(1+D(s))}{|\mathcal{D}(s)|}}}_{:=\xi_{1}}$ (A.10)
$-\underbrace{\frac{\alpha}{1-\gamma}\mathbb{E}_{s\sim d_{T}^{\mu}(s)}\bigg{[}D(s)\bigg{]}}_{:=\xi_{2}},$
where ||$\cdot$|| ⋅ | denotes cardinality of a specific set, $D(s)=\sum_{a}\widehat{\pi}(a|s)(\frac{\widehat{\pi}(a|s)}{\mu(a|s)}-1)$ and $C_{T,R}$ is the empirical coefficient. It consists of two terms: the first term $\xi_{1}$ captures the decrease in policy performance due to sampling error. The second term $\xi_{2}$ captures the increase in policy performance in empirical setting [4].
Therefore, when considering distribution shift and empirical error, we can derive
$J(\mu,R^{*})-J(\widehat{\pi},R^{*})$
$\leq$ $J(\mu,R^{*})-J(\mu,\widehat{R})+J(\widehat{\pi},\widehat{R})+\xi-J(\widehat{\pi},R^{*})$
$\leq$ $\xi+\big{|}J(\mu,R^{*})-J(\mu,\widehat{R})\big{|}+\big{|}J(\widehat{\pi},R^{*})-J(\widehat{\pi},\widehat{R})\big{|}.$
Furthermore, according to Eq. (A.9), we have
$J(\mu,R^{*})-J(\widehat{\pi},R^{*})\leq\xi+\frac{1+\mathscr{C}_{\mathcal{R}}(\widehat{\pi})}{1-\gamma}$
$\bigg{(}\sqrt{\frac{4C}{NL^{2}}\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}+\sqrt{\frac{4R_{max}^{2}\log\big{(}1/\delta\big{)}}{NL}}\bigg{)}.$
This completes the proof for Theorem 2.
A-D Proof of Proposition 3
Lemma 2.
There exists an absolute constant $C$ such that for any $\delta\in(0,1]$, with probability at least $1-\delta$, the following holds for all $\widehat{R}\in\mathcal{R}$
$\sum_{i=1}^{N}\Big{(}{P_{\widehat{R}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})}-{P_{R^{*}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})}\Big{)}^{2}\leq C\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right),$
where $y\in\{0,1\}$, $R^{*}$ is the true reward model and $\widehat{R}$ is the learned reward model. $P_{R}(0~{}|~{}\sigma_{0},\sigma_{1})$ is the probability that $\sigma_{0}$ is preferable $\sigma_{1}$ under reward model $R$. $\mathcal{N}_{\mathcal{R}}(1/N)$ is the bracketing number defined in Definition 3. This proof can be found in Lemma 2 of previous work [18] and Proposition 14 of [23].
The Proof for Proposition 3: We prove this Proposition from two steps. Firstly, we bound the difference between $P_{R^{*}}(\cdot~{}|~{}\sigma_{0},\sigma_{1})$ and $P_{\widehat{R}}(\cdot~{}|~{}\sigma_{0},\sigma_{1})$. Then, we bound the difference between $R^{*}(s,a)$ and $\widehat{R}(s,a)$.
Step 1: Bound the probability difference $|P_{R^{*}}(\cdot~{}|~{}\sigma_{0},\sigma_{1})-P_{\widehat{R}}(\cdot~{}|~{}\sigma_{0},\sigma_{1})|$. By Cauchy-Schwarz inequality $\big{(}\sum_{i}a_{i}b_{i}\big{)}^{2}\leq\big{(}\sum_{i}a_{i}^{2}\big{)}\big{(}\sum_{i}b_{i}^{2}\big{)}$, we set $a_{i}=P_{\widehat{R}}-P_{R^{*}}$ and $b_{i}$ that is chosen from $\{-1,1\}$ and satisfies $a_{i}b_{i}>0$. Then, we have
$\sum_{i=1}^{N}\Big{(}{P_{\widehat{R}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})}-{P_{R^{*}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})}\Big{)}^{2}$ (A.11)
$\geq$ $\frac{1}{N}\bigg{(}\sum_{i=1}^{N}\Big{|}P_{\widehat{R}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})-P_{R^{*}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})\Big{|}\bigg{)}^{2}.$
Then, by Lemma 2, the probability difference can be written as
$\sum_{i=1}^{N}\Big{|}P_{\widehat{R}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})-P_{R^{*}}(y^{i}|\sigma_{0}^{i},\sigma_{1}^{i})\Big{|}$ (A.12)
$\leq$ $\sqrt{CN\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}.$
Step 2: Bound the reward difference $|R^{*}(s,a)-\widehat{R}(s,a)|$. Let $f(\sigma)=\sum_{i}R^{*}(s^{i},a^{i})\in[-LR_{max},LR_{max}]$, $g(\sigma)=\sum_{i}\widehat{R}(s^{i},a^{i})\in[-LR_{max},LR_{max}]$, and $F(x_{1},x_{2})=e^{x_{1}}/(e^{x_{1}}+e^{x_{2}})$. Then, according to Eq. (2), we can derive
$\Big{|}P_{\widehat{R}}(y|\sigma_{0},\sigma_{1})-P_{R^{*}}(y|\sigma_{0},\sigma_{1})\Big{|}$ (A.13)
$=$ $\Big{|}F\big{(}f(\sigma_{0}),f(\sigma_{1})\big{)}-F\big{(}g(\sigma_{0}),g(\sigma_{1})\big{)}\Big{|}.$
Notably, the above equation only considers the condition $y=0$ since the result of $y=1$ is similar to $y=0$. Then, according to error propagation, we have
$\Big{|}F\big{(}f(\sigma_{0}),f(\sigma_{1})\big{)}-F\big{(}g(\sigma_{0}),g(\sigma_{1})\big{)}\Big{|}$ (A.14)
$\approx$ $\Bigg{|}\frac{\partial F}{\partial x_{1}}\Bigg{|}\Big{|}f(\sigma_{0})-g(\sigma_{0})\Big{|}+\Bigg{|}\frac{\partial F}{\partial x_{2}}\Bigg{|}\Big{|}f(\sigma_{1})-g(\sigma_{1})\Big{|}$
$\geq$ $\frac{e^{x_{1}+x_{2}}}{\big{(}e^{x_{1}}+e^{x_{2}}\big{)}^{2}}\Big{|}\Big{(}f(\sigma_{0})-g(\sigma_{0})\Big{)}+\Big{(}f(\sigma_{1})-g(\sigma_{1})\Big{)}\Big{|}$
$\geq$ $\frac{1}{4}\bigg{|}\sum_{l=1}^{L}\sum_{j=0}^{1}\Big{(}R^{*}(s_{j}^{l},a_{j}^{l})-\widehat{R}(s_{j}^{l},a_{j}^{l})\Big{)}\bigg{|}.$
Then, combing Eqs. (A.12) and (A.14), we can further derive
$\bigg{|}\sum_{i=1}^{N}\sum_{l=1}^{L}\sum_{j=0}^{1}\Big{(}R^{*}(s_{j}^{i,l},a_{j}^{i,l})-\widehat{R}(s_{j}^{i,l},a_{j}^{i,l})\Big{)}\bigg{|}$ (A.15)
$\leq$ $4\sqrt{CN\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}.$
According to Chernoff-Hoeffding bound [42], for independent random variables $X_{1},X_{2},...,X_{n}$, with high probability $1-\delta$, the below equation holds
${\color[rgb]{0,0,0}\mathbb{E}[X]\leq\frac{1}{n}\sum_{i=1}^{n}X_{i}+(b-a)\sqrt{\frac{\log(1/\delta)}{2n}},}$ (A.16)
where $[a,b]$ is the range of values that each $X_{i}$ can take. Then, for Eq. (A.15), we set $X_{i}$ as $R^{*}(s_{j},a_{j})-R(s_{j},a_{j})$ and $X_{i}\in[-2R_{max},2R_{max}])$. Then, the below equation holds
$\bigg{|}\mathbb{E}_{(s,a)\sim d_{T}^{\mu}(s,a)}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}\bigg{|}$ (A.17)
$\leq$ $\sqrt{\frac{4C}{NL^{2}}\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}+{\color[rgb]{0,0,0}\sqrt{\frac{4R_{max}^{2}\log\big{(}1/\delta\big{)}}{NL}}},$
where $d_{T}^{\mu}(s,a)$ is the distribution of offline dataset. The above equation holds since the trajectories $\{\sigma_{0}^{i},\sigma_{1}^{i}\}_{i=1}^{N}$ are collected from offline dataset. This completes proof of Proposition 3.

### v2

@@
 Abstract
-Offline preference-based reinforcement learning (PbRL) provides an effective way to overcome the challenges of designing reward and the high costs of online interaction.
+Offline preference-based reinforcement learning (PbRL) offers an effective approach to addressing the challenges of designing rewards and mitigating the high costs associated with online interaction.
 However, since labeling preference needs real-time human feedback, acquiring sufficient preference labels is challenging.
-To solve this, this paper proposes a offLine prEference-bAsed RL with high Sample Efficiency (LEASE) algorithm, where a learned transition model is leveraged to generate unlabeled preference data.
-Considering the pretrained reward model may generate incorrect labels for unlabeled data, we design an uncertainty-aware mechanism to ensure the performance of reward model, where only high confidence and low variance data are selected.
-Moreover, we provide the generalization bound of reward model to analyze the factors influencing reward accuracy, and demonstrate that the policy learned by LEASE has theoretical improvement guarantee.
-The developed theory is based on state-action pair, which can be easily combined with other offline algorithms.
-The experimental results show that LEASE can achieve comparable performance to baseline under fewer preference data without online interaction.
-Index Terms:
-Preference-based reinforcement learning; Sample efficiency; Reward learning; Theoretical guarantee
+To solve this, this paper proposes an offLine prEference-bAsed RL with high Sample Efficiency (LEASE) algorithm, where a learned transition model is leveraged to generate unlabeled preference data.
+Considering the pretrained reward model may generate incorrect labels for unlabeled data, we design an uncertainty-aware mechanism to ensure the performance of the reward model, where only high-confidence and low-variance data are selected.
+Moreover, the generalization bound of the reward model is provided to analyze the factors influencing reward accuracy, and the policy learned by LEASE has a theoretical improvement guarantee.
+The above developed theory is based on a state-action pair, which can be easily combined with other offline algorithms.
+The experimental results show that LEASE can achieve comparable performance to the baseline under fewer preference data without online interaction.
 I Introduction
@@
 However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [39]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9].
-Many researches have been explored to solve the above two challenges.
-Offline RL, learning policy from previous collected dataset without interaction with environment, has become an effective way for challenge 1 [4, 8].
-However, due to limited data coverage, offline RL suffers from the distribution shift between the learned policy and the behavior policy [28], which leads to poor agent performance.
-To address this problem, offline RL either directly constrain the policy closed to the behavior policy [2] or penalize value function for out-of-distribution (OOD) data [4, 8].
-The above methods improve the performance of agent through incorporating conservatism into policy learning.
+Much research has been conducted to solve the above two challenges.
+Offline RL, learning policy from previously collected dataset without interaction with the environment, has become an effective way for challenge 1 [4, 8].
+However, due to limited data coverage, offline RL suffers from the distribution shift between the learned policy and the behavior policy [28, 48], which leads to poor agent performance.
+To address this problem, offline RL either directly constrains the policy close to the behavior policy [2, 47] or penalizes the value function for out-of-distribution (OOD) data [4, 8, 46].
+The above methods improve the performance of the agent by incorporating conservatism into policy learning.
 Preference-based RL (PbRL), also known as RL from human feedback (RLHF), which learns the reward function from human preference data without the demand for tedious hand-engineered reward design [14, 42], has been developed recently for challenge 2.
 However, compared with data $(s,a,s^{\prime})$, the collecting cost of preference data is higher since it demands human real-time feedback to label preference, which often brings high sample complexity (low sample efficiency) [13, 3].
-Some data augmentation techniques, such as generating unlabeled data through interaction with simulator [3] or temporal cropping [35], have been proposed to solve problem of the limitation of preference data.
-However, the above methods need online interaction with environment (simulator) and still demand large human feedback.
+Some data augmentation techniques, such as generating unlabeled data through interaction with a simulator [3] or temporal cropping [35], have been proposed to solve the problem of the limitation of preference data.
+However, the above methods need online interaction with the environment (simulator) and still demand large human feedback.
 Thus, in PbRL, the sample efficiency is still of utmost importance.
-In offline PbRL setting, existing methods [34, 17, 33] typically involve two steps: reward learning and policy learning.
-They learns the policy based on offline static dataset without reward and preference dataset, and introduce transformer [17] or diffusion model [33] to improve the performance of reward model and agent performance.
+In offline PbRL settings, existing methods [34, 17, 33] typically involve two steps: reward learning and policy learning.
+They learns the policy based on an offline static dataset without a reward and preference dataset, and introduce transformer [17] or diffusion model [33] to improve the performance of the reward model and agent performance.
 However, the theoretical guarantees of reward generalization and policy improvement are not provided.
-Therefore, this paper aims to design a high sample efficiency offline PbRL algorithm that can achieve comparable performance with baseline from fewer preference data without interaction with environment, and develop the related theory for reward generalization bound and policy improvement guarantee.
-The key questions can be summarized as follow:
+Therefore, this paper aims to design a high sample efficiency offline PbRL algorithm that can achieve comparable performance with the baseline from fewer preference data without interaction with the environment, and develop the related theory for reward generalization bound and policy improvement guarantee.
+The key questions can be summarized as follows:
   1. 1.
-    In algorithm, the collected preference data fails to cover all state-action region and the collecting cost is too high.
-How to use the limited preference data to learn accurate reward model and guarantee agent performance?
+    In the algorithm, the collected preference data fails to cover all state-action regions, and the collection cost is too high.
+How to use the limited preference data to learn an accurate reward model and guarantee agent performance?
   2. 2.
-    In theory, what factors affect the reward generalization and what is the detailed relationship between offline RL algorithm itself, the reward gap and the improvement of policy performance?
+    In theory, what factors affect the reward generalization, and what is the detailed relationship between the offline RL algorithm itself, the reward gap, and the improvement of policy performance?
 For question 1, motivated by model-based offline RL, which learns the environment model to broaden data coverage [37, 36], we also train the transition model to achieve data augmentation (improve sample efficiency).
-Specifically, the agent generates two different trajectories through interaction with the learned transition model, and then the pretrained reward model generates pseudo label for these two trajectories.
-However, the preference model may generate incorrect pseudo label, which may bring unstable training and low performance.
-For question 2, there are very fewer algorithms for offline PbRL theory.
+Specifically, the agent generates two different trajectories through interaction with the learned transition model, and then the pretrained reward model generates pseudo-labels for these two trajectories.
+However, the preference model may generate incorrect pseudo-labels, which may lead to unstable training and low performance.
+For question 2, there are very few algorithms for offline PbRL theory.
 The most relevant work is [18], where the theory of policy improvement guarantee is established.
-However, the generalization bound for reward model is not considered, and this theory is based on the whole trajectory rather than state-action pair in offline algorithms, which is not conducive to the theoretical analysis of specific offline PbRL algorithms since many offline RL are based on state-action pair.
+However, the generalization bound for the reward model is not considered, and this theory is based on the whole trajectory rather than the state-action pair in offline algorithms.
+This is not conducive to the theoretical analysis of specific offline PbRL algorithms, since many offline RL algorithms are based on the state-action pair.
 Figure 1: The learning framework of LEASE.
-The offline dataset without reward is used to train transition model and limited labeled preference dataset is used to pretrain reward model.
-Then the generated unlabeled preference data is screened through uncertainty-aware selecting mechanism.
-The reward model is updated based on labeled and generated dataset.
-Finally, the agent learns the policy based on offline dataset and learned reward model.
-To solve the above problems, this paper proposes a novel offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE), where a selecting mechanism is designed to guarantee the quantity of generated dataset.
-This selecting mechanism is based on the principles of high confidence and low uncertainty, and can effectively reduce the error of pseudo-labels.
-Moreover, we develop a new generalization bound for reward model, and provide the theory of policy improvement guarantee for offline PbRL based on state-action pair.
+The offline dataset without reward is used to train the transition model, and a limited labeled preference dataset is used to pretrain the reward model.
+Then, the generated unlabeled preference data is screened through the uncertainty-aware selection mechanism.
+The reward model is updated based on the labeled and generated dataset.
+Finally, the agent learns the policy based on an offline dataset and a learned reward model.
+To solve the above problems, this paper proposes a novel offLine prEference-bAsed RL with a high Sample Efficiency algorithm (LEASE), where a selecting mechanism is designed to guarantee the quantity of the generated dataset.
+This selection mechanism is based on the principles of high confidence and low uncertainty, and can effectively reduce the error of pseudo-labels.
+Moreover, we develop a new generalization bound for the reward model and provide a theory of policy improvement guarantee for offline PbRL based on the state-action pair.
 Fig.
 1 shows the learning framework of LEASE.
-The theoretical and experimental results show that LEASE has policy improvement guarantee and can achieve comparable performance to baseline under fewer preference data on D4RL benchmark.
+The theoretical and experimental results show that LEASE has a policy improvement guarantee and can achieve comparable performance to the baseline under fewer preference data on the D4RL benchmark.
 The contributions are given below:
@@
   2. 2.
-    An uncertainty-aware mechanism is designed for screening the generated preference data so as to guarantee the stability of reward training and improve the accuracy of reward prediction.
+    An uncertainty-aware mechanism is designed for screening the generated preference data to guarantee the stability of reward training and improve the accuracy of reward prediction.
   3. 3.
-    The generalization bound of reward model and the theory of policy improvement based on state-action pair are developed to analyze factors that influence reward and policy performance.
+    The generalization bound of the reward model and the theory of policy improvement based on the state-action pair are developed to analyze factors that influence reward and policy performance.
 The framework of this paper is as follows: Section II introduces the related works about offline PbRL.
-Section III provides the basic notations for algorithm.
-Section IV illustrates the two main problems that this paper aims to solve.
-Section V introduces the details of designed selecting mechanism and derives the generalization bound for reward model.
-Section VI illustrates the implementation details of the proposed method and provides theory for offline PbRL.
-Section VII conducts experiments to validate the effectiveness of LEASE and analyze the effects of each part for agent performance.
-Section VIII further discusses the offline PbRL theory and algorithm.
-Section IX summarizes the entire work.
-II Related Work
-This section presents a brief overview for the algoriths and related theories of offline PbRL, and the differences and advantages between our work and the previous related works.
+Section III provides the basic notations for the algorithm and illustrates the two main problems that this paper aims to solve.
+Section IV introduces the details of the designed selection mechanism and derives the generalization bound for the reward model.
+Section V illustrates the implementation details of the proposed method and provides theory for offline PbRL.
+Section VI conducts experiments to validate the effectiveness of LEASE and analyze the effects of each part on agent performance.
+Section VII further discusses the offline PbRL theory and algorithm.
+Section VIII summarizes the entire work.
+II Related Works
+This section presents a brief overview of the algorithms and related theories of offline PbRL, and discusses the differences and advantages between our work and the previous related works.
 The algorithm for offline PbRL.
-Offline PbRL eliminates the demand for interaction with environment and handcraft designed reward.
-OPAL [34] is the first algorithm that combining offline RL and PbRL.
-PT [17] utilized transformer-based architecture to design preference model capable of generating non-Markovian rewards.
+Offline PbRL eliminates the demand for interaction with the environment and handcrafted-designed reward.
+OPAL [34] is the first algorithm that combines offline RL and PbRL.
+PT [17] utilized a transformer-based architecture to design a preference model capable of generating non-Markovian rewards.
 OPPO [30] directly optimized the policy in a high-level embedding space without learning a separate reward function.
-However, the above methods demands a large amount of preference dataset.
-FTB [33] trains a diffusion model to achieve data augmentation, but this consumers larger training cost.
-Instead, our method LEASE trains simple transition model to generate unlabeled data and designed selecting mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer dataset and less time.
+However, the above methods demand a large amount of preference data.
+FTB [33] trains a diffusion model to achieve data augmentation, but this consumes a larger training cost.
+Instead, our method LEASE trains a simple transition model to generate unlabeled data and designs a selection mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer datasets and less time.
 The theory for offline PbRL.
-There are few algorithms that provide theoretical guarantees for offline PbRL, including the generalization bound of reward model and the guarantee of policy improvement.
-Zhu et al.[27] studied offline PbRL, but the analysis are restricted to linear model.
-Zhan et al.[18] extended to general function approximation, but the generalization analysis of reward model is not provided and the theory is based on trajectory.
-LEASE gives the theoretical analysis for reward model, and provides the theoretical guarantee for policy based on state-action pair.
-The generalization bound of reward model has generality and can be applicable to methods that train reward model using pseudo-labels, and the theory of policy improvement guarantee of offline PbRL can be easily combined with other offline RL.
+There are a few algorithms that provide theoretical guarantees for offline PbRL, including the generalization bound of the reward model and the guarantee of policy improvement.
+Zhu et al.[27] studied offline PbRL, but the analysis is restricted to the linear model.
+Zhan et al.[18] extended to general function approximation, but the generalization analysis of the reward model is not provided, and the theory is based on the trajectory.
+LEASE gives the theoretical analysis for the reward model, and provides the theoretical guarantee for the policy based on the state-action pair.
+The generalization bound of the reward model has generality.
+It can be applied to methods that train a reward model using pseudo-labels, and the theory of policy improvement guarantees in offline PbRL can be easily combined with other offline RL approaches.
 Data selection techniques.
-Data screening mechanism essentially belongs to the field of semi-supervised learning.
-The goal of semi-supervised learning is using unlabeled data to improve model performance when labeled data is limited.
-Data selection techniques is used to filter the data with clean labels from a noisy dataset.
+The data screening mechanism essentially belongs to the field of semi-supervised learning.
+The goal of semi-supervised learning is to use unlabeled data to improve model performance when labeled data is limited.
+Data selection techniques are used to filter the data with clean labels from a noisy dataset.
 Han et al.[7] selected unlabeled data with small losses from one network.
@@
 Xiao et al.[6] selected data with high confidence as clean data.
-Motivated by [31] and [6], we trained ensemble reward model to select data with high confidence and low variance to guarantee the quality of unlabeled data.
-In PbRL methods, Surf [3] is the most similar to our method LEASE that uses data augmentation technique.
-The differences between them mainly includes the below three aspects: 1) Surf belongs to online RL, but LEASE belongs to offline RL.
-Surf generates data through interaction with environment (simulator) while LEASE generates data through the learned transition model.
-Note that, in most human-in-loop-control scenarios, the cost of online interaction is large and the realistic simulator is difficult to create; 2) Surf only uses confidence for label filtering, whereas LEASE employs both confidence and uncertainty principles, which effectively reduce pseudo-label error.
-The results in Table III have validated this; 3) LEASE provides the general theoretical framework for offline PbRL, but Surf don’t provide theoretical analysis.
-III Preliminaries
-In this section, we provides the basic notations for reinforcement learning algorithm, including offline RL, preference-based RL and model-based RL.
+Motivated by [31] and [6], we trained an ensemble reward model to select data with high confidence and low variance to guarantee the quality of unlabeled data.
+III Preliminaries and Problem Formulation
+In this section, we provide the basic notations for reinforcement learning algorithms, including offline RL, preference-based RL, and model-based RL.
+Then, we introduce the two main problems that this paper aims to solve from both algorithmic and theoretical perspectives.
+III-A Preliminaries for RL
 Offline reinforcement learning.
 The framework of RL is based on the Markov Decision Process (MDP) that is described by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},R,T,\rho,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})$ is the transition dynamics, $\rho$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor [39].
-The term $\Delta({\Omega})$ denotes the set of probability distribution over space $\Omega$.
+The term $\Delta({\Omega})$ denotes the set of probability distributions over space $\Omega$.
 The goal of RL is to optimize the policy $\pi$ that maximizes the expected discounted return $J(\pi,R):=\mathbb{E}_{(s,a)\sim d_{T}^{\pi}(s,a)}[R(s,a)]/(1-\gamma)$, where $d_{T}^{\pi}(s,a):=d_{T}^{\pi}(s)\pi(a|s)$ is the state-action marginal distribution under the learned policy $\pi$.
@@
 Preference-based reinforcement learning.
-Different from standard RL setting, the reward function is not available in PbRL.
+Different from the standard RL setting, the reward function is not available in PbRL.
 Instead, PbRL learns the reward function $\widehat{R}$ from preferences between pairs of trajectory segments to align the human intention [15], where a trajectory segment $\sigma$ of length $L$ is defined as a sequence of states and actions $\{s_{k},a_{k},...,s_{k+L-1},a_{k+L-1}\}\in(\mathcal{S}\times\mathcal{A})^{L}$.
@@
 After obtaining the reward function $\widehat{R}$, offline PbRL optimizes the policy by maximizing the expected discounted return $J(\pi,R)$ like standard offline RL.
-Model-based reinforcement learning.
-Model-based RL learns the dynamics model to improve sample efficiency [38, 36, 39].
-They used the offline dataset $\mathcal{D}_{\text{offline }}$ to estimate transition model $\widehat{T}(s^{\prime}|s,a)$.
-The transition model $\widehat{T}_{\varphi}$ parameterized by $\varphi$ is typically trained via maximum likelihood estimation (MLE):
+[…]
 $\mathcal{L}_{T}(\varphi)=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\text{offline }}}\big[\log\widehat{T}_{\varphi}(s^{\prime}~|~s,a)\big].$ (4)
@@
 The samples are generated through $H$-step rollouts.
-Here, we use the trained transition to generate more trajectory to achieve data augmentation.
-IV Problem Formulation
+Here, we use the trained transition to generate more trajectories to achieve data augmentation.
+III-B Problem Formulation
 The cost of collecting preference data is high since it needs real-time human feedback.
-Therefore, this paper aims to improve sample efficiency to learn accurate reward model from limited preference data and guarantee agent performance.
-For PbRL, the ideal form is that the learned reward function $\widehat{R}$ from collected preference data can be consistent with true reward $R^{*}$.
-Here, we define the function class as =$\mathcal{R}=\{R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$.
-Then we assume the reward class $\mathcal{R}$ is realizable.
+Therefore, this paper aims to improve sample efficiency to learn an accurate reward model from limited preference data and guarantee agent performance.
+For PbRL, the ideal form is that the learned reward function $\widehat{R}$ from collected preference data can be consistent with the true reward $R^{*}$.
+Here, we define the function class as $\mathcal{R}=\{R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$.
+Then, similar to [44, 43], we make the below assumption for the reward model:
 Assumption 1 (Realizability).
@@
 Then, for any distribution $d(s,a)$, $\inf_{R\in\mathcal{R}}\mathbb{E}_{(s,a)\sim d(s,a)}[R^{*}(s,a)-R(s,a)]^{2}<\varepsilon_{R}$ holds.
-Based on the above assumption, the optimal reward function can be obtained if the preference data can cover all data space.
-However, preference data tend to be limited, even less than hundred in real-world scenarios, such as rehabilitation robot field.
-Therefore, different from previous PbRL [3, 35, 33], this paper learns reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$.
-We train the transition model through Eq.
+This assumption states that there must exist a reward function $R\in\mathcal{R}$ which can well-approximate the optimal reward function when the preference data can cover all data space.
+However, preference data tend to be limited, even fewer than a hundred in real-world scenarios.
+Therefore, different from previous PbRL [3, 35, 33], this paper aims to study how to learn the better reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$.
+To broaden the coverage of preference data, motivated by model-based offline RL methods, we train the transition model through Eq.
 (4) to generate more unlabeled preference data $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$.
-The pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
+Then, the pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
 $\widehat{y}\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)=\mathbbm{1}\big[P\left(\sigma_{1}^{u}\succ\sigma_{0}^{u};{\psi}\right)>0.5\big],$ (5)
-where $\mathbbm{1}(\cdot)$ is indicator function.The reward function is updated through collected labeled dataset $\mathcal{D}_{l}$ and generated unlabeled dataset $\mathcal{D}_{u}$.
-Then, the reward model can be optimized through minimizing
+where $\mathbbm{1}(\cdot)$ is indicator function.
+The reward function can be updated through the collected labeled dataset $\mathcal{D}_{l}$ and the generated unlabeled dataset $\mathcal{D}_{u}$.
+Then, the reward model can be optimized by minimizing
 $\mathcal{L}^{\prime}_{R}(\psi)=$ $\mathbb{E}_{(\sigma_{0}^{l},\sigma_{1}^{l},y)\sim\mathcal{D}_{l}}\big[L\left((\sigma_{0}^{l},\sigma_{1}^{l}),y\right)\big]$ (6)
-$+$ $\mathbb{E}_{(\sigma_{0}^{u},\sigma_{1}^{u})\sim\mathcal{D}_{u}}\big{[}L\left((\sigma_{0}^{u},\sigma_{1}^{u}),\widehat{y}\right)\big{]},$
+[…]
 where $L((\sigma_{0},\sigma_{1}),y)=-(1-y)\log P(\sigma_{0}\succ\sigma_{1};{\psi})-y\log P(\sigma_{1}\succ\sigma_{0};{\psi})$.
-However, the pretrained reward model may generate incorrect pseudo-labels for unlabeled dataset, leading to noisy training and poor generalization [31].
-Therefore, one key question is how to design a data selecting mechanism to improve prediction accuracy and guarantee training stability.
-The another key aspect is the theory for offline PbRL.
-There are very few algorithms specifically designed for offline PbRL with strong guarantee, including generalization bound for reward model and safe improvement guarantee for policy learning.
-Therefore, another key question is to develop a systematic theory for offline PbRL, including generalization bound and improvement guarantee.
-V Reward Learning
+However, the pretrained reward model may generate incorrect pseudo-labels for the unlabeled dataset, leading to noisy training and poor generalization [31].
+Therefore, one key question is how to design a data selection mechanism to improve prediction accuracy and guarantee training stability.
+Another key aspect is the theory for offline PbRL.
+There are very few algorithms specifically designed for offline PbRL with strong guarantees, including a generalization bound for the reward model and a safe improvement guarantee for policy learning.
+Although Zhan et al.[18] established the lower bounds of policy improvement under partial coverage for offline PbRL, the general reward functions are defined over the whole trajectory rather than just state-action pairs, and the generalization bound of the reward model is not considered.
+Therefore, another key question is to develop a systematic theory for offline PbRL based on state-action pairs, including a generalization bound for the reward model and an improvement guarantee for policy learning.
+IV Reward Learning
 The reward learning involves two stages: pretraining based on collected labeled data $\mathcal{D}_{l}$ and updating based on $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$ during policy learning, which is essentially semi-supervised learning [26].
-This section focuses on designing a data selecting function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of generated preference data, and explaining the factors that influence the generalization ability of reward model.
-V-A Uncertainty-aware Pseudo-label Selection
-Motivated by the previous pseudo-labeling work [31], we select data from unlabeled dataset $\mathcal{D}_{u}$ according two principles: confidence and uncertainty.
-The data with high confidence and low uncertainty can be chosen for reward training.
-High confidence refers that pre-trained reward model discriminates the preference of two trajectories into $\widehat{y}$ with high probability $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$.
-Low uncertainty refers that the $N_{R}$ reward models (model ensembles) predicts with small variance $\tau(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})$.
-The probability confidence $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$ and uncertainty variance $\tau(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})$ can be denoted as
+This section focuses on designing a data selection function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of generated preference data, and explaining the factors that influence the generalization ability of the reward model.
+IV-A Uncertainty-aware Pseudo-label Selection
+[…]
 $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})=(1-\widehat{y})\bar{P}_{\psi}\left(\sigma_{0}^{u}\succ\sigma_{1}^{u}\right)+\widehat{y}\bar{P}_{\psi}\left(\sigma_{1}^{u}\succ\sigma_{0}^{u}\right)$ (7)
@@
 $f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes the generated data $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected, and $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ denotes the data is not selected.
-Then combing Eq.
-(6), the reward model can be optimized through minimizing
-^$\widehat{\mathcal{L}}_{R}(\psi)=\frac{1}{\widetilde{N}_{u}}\sum\nolimits_{j=1}^{{N}_{u}}f\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)^{(j)}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$ (9)
+Intuitively, more rigorous data selection criteria, characterized by higher confidence and lower variance, lead to a reduction in the number of qualified samples.
+While such a strategy improves the quality of the generated data, it could lead to a narrower data distribution, which might adversely affect the precision of the trained reward model.
+Therefore, appropriate selection of parameters $\kappa_{p}$ and $\kappa_{\tau}$ is required to enhance the performance of the reward model.
+Then, combining Eq.
+(6) and the designed selection mechanism, the reward model can be optimized by minimizing
+$\widehat{\mathcal{L}}_{R}(\psi)=\frac{1}{\widetilde{N}_{u}}\sum\nolimits_{j=1}^{{N}_{u}}f\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)^{(j)}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$ (9)
 $\frac{1}{N_{l}}\sum\nolimits_{i=1}^{N_{l}}L\left(\sigma_{0}^{l},\sigma_{1}^{l})^{(i)},y^{i}\right),$
-where $\widetilde{N}_{u}$ is the number of generated dataset after screening, and ^$\widehat{\mathcal{L}}_{R}(\psi)$ is the empirical risk.
-It contains two parts: labeled loss ^$\widehat{\mathcal{L}}_{l}(\psi)$ and unlabeled loss ^$\widehat{\mathcal{L}}_{u}(\psi)$.
-Note that the label in unlabeled loss is pseudo-label, which may be different from the true label.
-Therefore, there is the gap between the unlabeled loss with pseudo-label ^$\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$.
-Before bounding this gap, we firstly give the below assumption without loss of generality.
+where $\widetilde{N}_{u}$ is the number of generated dataset after screening, and $\widehat{\mathcal{L}}_{R}(\psi)$ is the empirical risk.
+It contains two parts: labeled loss $\widehat{\mathcal{L}}_{l}(\psi)$ and unlabeled loss $\widehat{\mathcal{L}}_{u}(\psi)$.
+Note that the label in the unlabeled loss is a pseudo-label, which may be different from the true label.
+Therefore, there is a gap between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$.
+Before bounding this gap, we first give the following assumption without loss of generality.
 Assumption 2.
@@
 (5), then the pseudo-labeling error for the unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(j)}\}_{j=1}^{N_{u}}$ is smaller than $\eta$, i.e., $\sum_{j=1}^{N_{u}}\mathbbm{1}[\widehat{y}^{j}\neq y^{j}]/N_{u}\leq\eta$.
+Note that the bound of pseudo-label error rate $\eta$ of the generated data filtered by the proposed selection mechanism is significantly lower than that of the unfiltered data, as the mechanism effectively eliminates low-confidence and high-uncertainty samples.
+This observation is experimentally validated in Table II and Fig.
+3.
+Then, based on the above assumption, we bound the difference between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ and give the following proposition.
 Proposition 1.
 Suppose that the loss $L((\sigma_{0},\sigma_{1}),y))$ is bounded by $\Omega$. Then, for any $R\in\mathcal{R}$, under Assumption 2 the following equation holds:
-|^$\big{|}\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\big{|}\leq\eta\Omega.$ (10)
-The proof of Proposition 1 can be found in Appendix A-A.
-This gap is mainly influenced by the term $\eta$.
-Through selecting high confidence and low uncertainty samples, the $\eta$ can be significantly reduced.
-Combining Eqs.
-(5) and (8), we find that if $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$ is small, the $L((\sigma_{0}^{u},\sigma_{1}^{u}),\widehat{y})$ would become larger.
-Too large error can lead to unstable training and low performance.
-Therefore, through selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, a more accurate subset of pseudo-labels can be used in reward training, and through setting a higher confidence threshold for pseudo labeling, a lower prediction error and the stability of training can be guaranteed.
-V-B Generalization Bound for Reward Model
+$\big|\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\big|\leq\eta\Omega.$ (10)
+[…]
+IV-B Generalization Bound for Reward Model
+This part studies the generalization performance of the reward model.
 Before developing generalization bound, we define the expected error $\mathcal{L}_{R}(\psi)$ with respect to the reward model $R(s,a;\psi)$: $\mathcal{L}_{R}(\psi)=\mathbb{E}_{(\sigma_{0},\sigma_{1},y)\sim\mathcal{D}}[L((\sigma_{0},\sigma_{1}),y)]$, where $\mathcal{D}$ is data distribution.
-The reward model is trained through minimizing the empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)$ in Eq.
+The reward model is trained through minimizing the empirical error $\widehat{\mathcal{L}}_{R}(\psi)$ in Eq.
 (9).
-Through developing generalization error bound, we can analyse the factors that influence generalization ability.
-Similar to the previous generalization bound works [29], Rademacher complexity, which measures the richness of a certain hypothesis space [19], is introduced firstly. The definition is given below:
+Through developing generalization error bound, we can analyze the factors that influence generalization ability.
+Similar to the previous generalization bound works [29], Rademacher complexity, which measures the richness of a certain hypothesis space [19], is introduced first. The definition is given below:
 Definition 1 (Empirical Rademacher complexity).
@@
 The empirical Rademacher complexity of $\mathcal{G}$ for sample $\widehat{\mathcal{S}}$ is defined as
-^$\widehat{\Re}_{\widehat{\mathcal{S}}}(\mathcal{F})=\mathbb{E}_{\bm{\sigma}}\big{[}\sup_{f\in\mathcal{F}}\frac{1}{N_{s}}\sum\nolimits_{i=1}^{N_{s}}\sigma_{i}f\left(z_{i}\right)\big{]},$ (11)
+$\widehat{\Re}_{\widehat{\mathcal{S}}}(\mathcal{F})=\mathbb{E}_{\bm{\sigma}}\big[\sup_{f\in\mathcal{F}}\frac{1}{N_{s}}\sum\nolimits_{i=1}^{N_{s}}\sigma_{i}f\left(z_{i}\right)\big],$ (11)
 where $\bm{\sigma}=(\sigma_{1},...,\sigma_{N})$ are independent uniform random variables taking values in $\{-1,+1\}$.
@@
 In general, lower Rademacher complexity corresponds to better generalization performance.
-Then, based on the above definition and Proposition 1, we develop the new generalization bound for reward model trained through Eq.
+Then, based on the above definition and Proposition 1, we develop the new generalization bound for the reward model trained through Eq.
 (9).
@@
 where $\widehat{\mathcal{D}}$ is the input combination of labeled and unlabeled dataset, i.e. $\{(\sigma_{0},\sigma_{1})^{(i)}\}_{i=1}^{N_{l}+N_{u}}$.
-The proof of Theorem 1 can be found in Appendix A-B.
-Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)$, pseudo-labeling error $\eta\Omega$, Rademacher complexity and constant terms.
+The proof of Theorem 1 can be found in Appendix B.
+Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by empirical error $\widehat{\mathcal{L}}_{R}(\psi)$, pseudo-labeling error $\eta\Omega$, Rademacher complexity and constant terms.
 The constant term can become small when the number of unlabeled data increases, but it may cause unstable training when labels are inaccurate.
-According to Proposition 1, empirical error and pseudo-labeling error can be reduced through selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, thus the better generalization ability, that is tighter upper bound, can be achieved.
-Note that Theorem 1 has generality and can be applicable to methods that train reward model using pseudo-labels.
-The difference among various methods may lie in how they train reward model to improve the accuracy of pseudo-labels, that is reducing $\eta$.
-If one method fails to reduce the pseudo-labeling error $\eta$, the upper generalization bound would be looser than that of no data augmentation.
-VI Policy Learning
-This paper directly uses the proposed offline RL algorithms, such as CQL [4] and IQL [25], to perform policy learning.
+According to Proposition 1, empirical error and pseudo-labeling error can be reduced through the selection mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, thus, a better generalization ability, that is tighter upper bound, can be achieved.
+Moreover, Theorem 1 has generality and can apply to methods that train a reward model using pseudo-labels.
+V Policy Learning
 To solve the problem of high cost for labeling preference, we propose a offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE).
-This section aims to describe the details of LEASE and establish the theory for safe policy improvement guarantee.
-VI-A The Implementation Details of LEASE
+This section aims to describe the details of LEASE and establish the theory for a safe policy improvement guarantee.
+In this paper, we directly use the proposed offline RL algorithms, such as CQL [4] and IQL [25], to perform policy learning, and do not improve the offline RL algorithm itself.
+V-A The Implementation Details of LEASE
+Most offline RL algorithms are based on Eq.
+(1).
+They proposed various algorithms to solve the problem of distribution shift.
+In an offline PbRL setting, the reward is unknown to the agent, and the accuracy of the reward model directly influences the performance of the policy.
+Moreover, the unlabeled data are generated through the current policy interaction with the learned transition model.
+The learned policy affects the distribution of the generated unlabeled preference dataset.
+In practical implementation, the policy, $Q$-function, reward and transition model are parameterized by related neural networks $\pi_{\theta}$, $Q_{\omega}$, $R_{\psi}$ and $T_{\varphi}$, respectively.
+Algorithm 1 gives pseudo-code for LEASE.
+The goal of LEASE is to learn a better reward model $R_{\psi}$ from a smaller labeled preference dataset $\mathcal{D}_{l}$ and learn a better policy $\pi_{\theta}$ from an offline dataset $\mathcal{D}_{\text{offline}}$ without interaction with the environment.
+The key to LEASE is how to train the reward model in the policy learning process.
+The proposed algorithmic framework comprises three key components: reward model training, transition model training, and policy training.
+Initially, the transition model is trained using limited offline datasets, while the reward model undergoes pretraining with a small-scale preference dataset.
+The reward update condition is triggered when the number of unlabeled data in the buffer reaches its maximum capacity.
+Subsequently, the policy training phase commences.
+During this phase, the transition model is periodically employed to generate unlabeled synthetic data, and the reward model undergoes iterative updates to enhance its performance.
+V-B Safe Policy Improvement Guarantee
+Offline RL aims to guarantee $\xi$-safe policy improvement over the behavior policy $\mu$ (the policy used to collect offline dataset), i.e.
+$J(\widehat{\pi},R^{*})\geq J(\mu,R^{*})-\xi$ [4].
+For offline PbRL, we further investigate whether similar policy improvement guarantees can be established.
+Therefore, this part aims to develop a theoretical guarantee of policy improvement for LEASE, that is, giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$.
+Then, based on the bounds of policy improvement, we further analyze the sample complexity for LEASE when the reward model is unknown.
 1
@@
 4:for $t=1,2,\cdots,N_{\text{iter}}$ do
-5:     if $t$ rollout frequency $=0$ then
-6:        Generate synthetic $H$-step rollouts by ${T}_{\varphi}$. Add transition trajectory into $\mathcal{D}_{u}$.
+5:  if $t$ % rollout frequency =0=0 then
+6:   Generate data by transition model ${T}_{\varphi}$, and add transition trajectory into $\mathcal{D}_{u}$.
 7:  end if
-8:     if reward update condition is TRUE then
-9:        Give pseudo label for unlabeled dataset $\mathcal{D}_{u}$ based on pre-trained reward model.
-10:        Select high confidence and low uncertainty unlabeled data according to Eq. (8).
-11:        Train the reward model based on $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ according to Eq. (9).
-12:     end if
-13:     Train policy using offline RL algorithms, such as CQL and IQL.
-14:  end for
+8:  if $t$ % reward update frequency =0=0 then
+9:   Generate pseudo label for unlabeled dataset $\mathcal{D}_{u}$, and screen unlabeled data by Eq. (8).
+10:   Train the reward model based on $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ according to Eq. (9).
+11:  end if
+12:  Train policy using offline RL algorithms.
+13:end for
 Algorithm 1 LEASE: offLine prEference-bAsed RL with high Sample Efficiency
-Most offline RL algorithms are based on Eq.
-(1).
-They proposed various algorithms to solve the problem of distribution shift.
-In offline PbRL setting, the reward is unknown for agent.
-The accuracy of reward model directly influence the performance of policy.
-Moreover, the unlabeled data are generated through the current policy interaction with learned transition model.
-Therefore, the reward model and policy influence each other.
-In practical implementation, the policy, $Q$-function, reward and transition model are parameterized by $\pi_{\theta}$, $Q_{\omega}$, $R_{\psi}$ and $T_{\varphi}$ respectively.
-Algorithm 1 gives pseudocode for LEASE.
-The goal of LEASE is to learn better reward model $R_{\psi}$ from fewer labeled preference dataset $\mathcal{D}_{l}$ and learn better policy $\pi_{\theta}$ from offline dataset $\mathcal{D}_{\text{offline}}$ without interaction with environment.
-The key of LEASE is how to train reward model in policy learning process.
-Notably, the reward model only update once instead of updating constantly in this process.
-The reward update condition is set to when the number of unlabeled data in buffer reaches the maximum buffer capacity.
-The update time is influenced by rollout frequency and rollout batch size.
-The reward update condition, rollout length $H$ and selecting mechanism $f(\sigma_{1},\sigma_{2})$ are the important factors for reward model.
-VI-B Safe Policy Improvement Guarantee
-Offline RL aims to guarantee $\xi$-safe policy improvement over the behavior policy $\mu$ (the policy used to collect offline dataset), i.e.
-$J(\widehat{\pi},R^{*})\geq J(\mu,R^{*})-\xi$ [4].
-This part develops theoretical guarantee of policy improvement for LEASE, that is giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$, and further analyzes sample complexity when reward model is unknown.
-Firstly, we propose a new single-policy concentrability coefficient for PbRL based on state-action pair instead of trajectory like [18].
+Before deriving the bounds of policy improvement, we first define a new single-policy concentrability coefficient for PbRL based on the state-action pair.
+This coefficient can quantify the discrepancy between the offline distribution and the visitation distribution of policy $\pi$.
 Definition 3 (Concentrability coefficient for PbRL).
 Concentrability coefficient $\mathscr{C}_{\mathcal{R}}(\pi)$ is used to measure how well reward model errors transfer between offline data distribution $d_{T}^{\mu}$ and the visitation distribution $d_{T}^{\pi}$ under transition $T$ and policy $\pi$, defined as
-$\mathscr{C}_{\mathcal{R}}(\pi)=\sup_{\widehat{R}\in\mathcal{R}}\Bigg{|}\frac{\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}}{\mathbb{E}_{(s,a)\sim d_{T}^{\mu}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}}\Bigg{|}.$ (13)
-where $R^{*}(s,a)$ is the true reward model and coefficient $\mathscr{C}_{\mathcal{R}}(\pi)$ is upper bounded by $\|d_{T}^{\pi}/d_{T}^{\mu}\|_{\infty}$.
-Next, similar to [20, 18], we use $\varepsilon$-bracketing number to measure the complexity of reward function class $\mathcal{R}$, which can be defined as
+$\mathscr{C}_{\mathcal{R}}(\pi)=\Bigg|\frac{\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]}{\mathbb{E}_{(s,a)\sim d_{T}^{\mu}}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]}\Bigg|,$ (13)
+where $R^{*}(s,a)$ is the true reward model and $\hat{R}(s,a)$ is the learned reward model.
+[…]
 Definition 4 ($\varepsilon$-bracketing number).
@@
 (2).
+Bracketing number $\mathcal{N}_{\mathcal{R}}(\varepsilon)$ measures the complexity of function class $\mathcal{R}$ and takes $\widetilde{\mathcal{O}}(d)$ in linear reward model [18].
+Then, before giving the theorem of policy improvement, we first assume the reward class $\mathcal{R}$ is bounded, that is
+Assumption 3 (Boundedness).
+For any $R\in\mathcal{R}$ and any state-action pairs $(s,a)$, the equation $|R(s,a)|\leq R_{max}$ holds.
 Theorem 2.
-Under Assumption 1, for any $\delta\in(0,1]$, the policy $\widehat{\pi}$ learned by LEASE, with high probability $1-\delta$, satisfies that $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$ is upper bound by
+Under Assumptions 1 and 3, for any $\delta\in(0,1]$, the policy $\widehat{\pi}$ learned by LEASE, with high probability $1-\delta$, satisfies that $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$ is upper bound by
 $\xi+\frac{1+\mathscr{C}_{\mathcal{R}}(\widehat{\pi})}{1-\gamma}\Bigg(\sqrt{\frac{4C}{NL^{2}}\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}$ (14)
 $+{\color[rgb]{0,0,0}\sqrt{\frac{4R_{max}^{2}\log\big(1/\delta\big)}{NL}}}\Bigg),$
-where $C>0$ is the absolute constant defined in Eq.
-(2), $N$ is the size of preference dataset, and $L$ is the length of trajectory.
-The term $\xi$ is performance gap depending on offline algorithm itself.
-The proof of Theorem 2 can be found in Appendix A-C.
-The $\xi$ is constant when offline algorithm is determined.
-Therefore, the tighter bound can be achieved through reducing performance gap caused by reward model (see Proposition 3).
-It can be reduced to small value $\varpi$ with sample complexity of
+where $C>0$ is the absolute constant defined in Lemma 2, $N$ is the size of preference dataset, and $L$ is the length of trajectory.
+The term $\xi$ is the performance gap depending on the offline algorithm itself.
+The proof of Theorem 2 can be found in Appendix C.
+The bounds of the PbRL policy learned by LEASE comprise two distinct components: the first term $\xi$ stems from the inherent gap of the offline RL itself, and the second arises from the discrepancy induced by the reward model.
+The first term $\xi$ is constant when the offline algorithm is determined.
+The second term is determined by the performance of the reward model.
+Therefore, the tighter bound can be achieved through reducing the performance gap caused by the reward model (see Proposition 3).
+It can be reduced to a small value $\varpi$ with a sample complexity of
 $N=\widetilde{\mathcal{O}}\Bigg(\frac{4\big(1+\mathscr{C}_{\mathcal{R}}(\widehat{\pi})\big)^{2}}{\varpi^{2}(1-\gamma)^{2}}\Bigg($ $\sqrt{\frac{C\log(\mathcal{N}_{\mathcal{R}}(1/N)/\delta)}{L^{2}}}$ (15)
-$+\sqrt{\frac{R_{max}^{2}\log\big{(}1/\delta\big{)}}{L}}\Bigg{)}^{2}\Bigg{)},$
-where concentrability coefficient $\mathscr{C}_{\mathcal{R}}(\widehat{\pi})$ can become smaller when the learned policy $\widehat{\pi}$ is close to behavior policy $\mu$.
-Bracketing number $\log(\mathcal{N}_{\mathcal{R}}(1/N))$ measures the complexity of function class $\mathcal{R}$ and takes $\widetilde{\mathcal{O}}(d)$ in linear reward model [18].
-Notably, LEASE can learn accurate reward model through data augmentation under fewer $N_{l}$ preference data ($N_{l}<N$).
-Thus, the performance gap can become tighter under fewer data, that is LEASE can have higher sample efficiency compared with [18].
-More discussion can be found in Section VIII-A.
-VII Experiments
-This section gives the details of experiments and focuses on answering the following questions:
-  • $\mathbf{Q_{1}}$: How well the accuracy of reward model through data augmentation under fewer preference data?
-  • $\mathbf{Q_{2}}$: How well does LEASE perform compared with offline PbRL baseline in standard benchmark tasks?
-  • $\mathbf{Q_{3}}$: How the proposed selecting mechanism and parameter settings affect the performance of reward and policy?
-We answer these questions on two distinct domains (mujoco and adroit) within the D4RL benchmark [22], covering several locomotion and manipulation tasks.
-The code for LEASE is available at anonymous.link.
-VII-A Experiment Setup
+$+\sqrt{\frac{R_{max}^{2}\log\big(1/\delta\big)}{L}}\Bigg)^{2}\Bigg).$
+[…]
+VI Experiments
+This section gives the details of experiments and presents comprehensive experimental evaluations of the proposed algorithm, including the accuracy assessment of reward model learning, performance analysis of policy learning, and detailed parameter sensitivity studies.
+We focus on answering the following questions:
+  • $\mathbf{Q_{1}}$: How well does the accuracy of the reward model through data augmentation under fewer preference data?
+  • $\mathbf{Q_{2}}$: How well does LEASE perform compared with the offline PbRL baseline in standard benchmark tasks?
+  • $\mathbf{Q_{3}}$: How does the proposed selection mechanism and parameter settings affect the performance of reward and policy?
+We answer these questions on two distinct domains (mujoco and adroit), covering several locomotion and manipulation tasks, within the D4RL benchmark [22]. The code for LEASE is available at Code.
+VI-A Experiment Setup
 Figure 2: The description of Mujoco tasks (locomotion tasks) and Adroit tasks (manipulation tasks).
-Experimental Datasets.
+VI-A1 Experimental Datasets
 The offline dataset and preference dataset of experiments originate from [22] and [21], respectively.
 The offline datasets are based on the D4RL dataset [22].
-We select medium and medium-expert two types dataset for Mujoco tasks, and human and expert for Adroit tasks.
-The difference between different dataset in certain task lies in the collected policy.
+We select medium and medium-expert two types of datasets for Mujoco tasks and human and expert for Adroit tasks.
+The difference between datasets in a certain task lies in the policies they collect.
+The Mujoco tasks include halfcheetah-v2, hopper-v2, and walker2d-v2.
+The Adroit tasks include pen-v1, door-v1, and hammer-v1.
 Fig.
-2 depicts the six tasks for D4RL benchmark.
+2 depicts the six tasks for the D4RL benchmark.
 For three mujoco tasks, the goal is to control the robot’s various joints to achieve faster and more stable locomotion.
 For adroit tasks, they involve controlling a 24-DoF simulated Shadow Hand robot on a sparse reward, high-dimensional robotic manipulation task.
-The preference dataset are from the previous work [21], where they provided baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF).
+The preference dataset is from the previous work [21], where they provided the baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF).
 One is crowd-sourced labels obtained by crowd-sourcing, and the other is synthetic labels based on ground truth reward.
-To test LEASE performance for two types preference data, the labels of preference is from human-realistic feedback for mujoco tasks, and ground truth reward for adroit tasks.
+To test LEASE performance for two types of preference data, the labels of preference are from human-realistic feedback for MuJoco tasks, and ground truth reward for Adroit tasks.
 This paper aims to test the performance under fewer preference data without online interaction.
-We take the first $100$ data of [21] as $2000$ preference dataset, and denote method without data augmentation as FEWER.
+We take the first $100$ data of [21] as $2000$ preference dataset, and denote the method without data augmentation as FEWER.
+VI-A2 Experimental Details
+The Offline PbRL setting involves two steps: reward learning and policy learning.
+Here, we directly use offline RL algorithms CQL [4] and IQL [25] to perform policy learning.
+The performance of these algorithms is evaluated based on the cumulative return.
+For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [22].
+The normalized score $\tilde{J}$ is computed by:
+$\tilde{J}=\frac{J-J_{r}}{J_{e}-J_{r}}\times 100,$ (16)
+where $J_{r}$, $J_{e}$, and $J$ are the expected returns of a random policy, an expert policy, and the trained policy by offline RL algorithms, respectively.
+Table I gives the hyperparameter configuration for reward and transition training, and policy learning.
+The details are given below:
 TABLE I: Base parameter configuration of method LEASE.
 Model Parameter Value
-Transition Model Model learning rate $1\times e^{-3}$
+Transition Model Model learning rate $1\times 10^{-3}$
 Number of hidden layers 4
@@
 Number of elites 5
-Reward Model Model learning rate $3\times e^{-4}$
+Reward Model Model learning rate $3\times 10^{-4}$
 Number of hidden layers 3
@@
 Number of labeled dataset $N_{l}$ 100
-Policy Model Actor learning rate (IQL) $1(3)\times e^{-4}$
-Critic learning rate $3\times e^{-4}$
+Policy Model Actor learning rate (IQL) $1(3)\times 10^{-4}$
+Critic learning rate $3\times 10^{-4}$
 Number of hidden layers(IQL) 3(2)
@@
 Batch size 256
-Experimental Details.
-Offline PbRL setting involves two steps: reward learning and policy learning.
-Here, we directly use offline RL algorithms CQL [4] and IQL [25] to perform policy learning.
-The performance of these algorithms is evaluated based on the cumulative return.
-For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [22].
-The normalized score $\tilde{J}$ is computed by:
-$\tilde{J}=\frac{J-J_{r}}{J_{e}-J_{r}}\times 100,$ (16)
-where $J_{r}$, $J_{e}$ and $J$ are the expected return of a random policy, an expert policy, and the trained policy by offline RL algorithms, respectively.
-Table I gives the hyper-parameter configuration for reward and transition training, and policy learning.
-The details are given the below:
-Figure 3: The relationship between preference model accuracy and the number of preference dataset $N_{l}$.
-TABLE II: The hyperparameters for LEASE under CQL and IQL algorithms.
-Task Name CQL IQL
-$H$ $\kappa_{p}$ $\kappa_{\tau}$ $H$ $\kappa_{p}$ $\kappa_{\tau}$
-walker2d-m $10$ $0.85$ $0.05$ $10$ $0.75$ $0.08$
-walker2d-m-e $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
-hopper-m $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
-hopper-m-e $100$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
-halfcheetah-m $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
-halfcheetah-m-e $200$ $0.85$ $0.05$ $10$ $0.75$ $0.08$
-pen-human $10$ $0.95$ $0.05$ $10$ $0.85$ $0.08$
-pen-expert $10$ $0.90$ $0.05$ $10$ $0.95$ $0.08$
-door-human $10$ $0.95$ $0.05$ $10$ $0.90$ $0.08$
-door-expert $20$ $0.95$ $0.05$ $10$ $0.95$ $0.08$
-hammer-human $10$ $0.95$ $0.05$ $10$ $0.90$ $0.08$
-hammer-expert $20$ $0.99$ $0.05$ $10$ $0.95$ $0.08$
-1) Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models.
+Figure 3: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) hopper-medium-expert and (b) halfcheetah-medium-expert datasets, where the value predicted by the trained reward model and ground-truth reward value are both normalized to $[0,1]$.
+From left to right are methods LEASE, FEWER and FRESH, respectively.
+Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models.
 Each model consists of a $4$-layer feed-forward neural network with $200$ hidden units.
-The model training employs maximum likelihood estimation with a learning rate of $1\times e^{-3}$ and Adam optimizer.
+The model training employs maximum likelihood estimation with a learning rate of $1\times 10^{-3}$ and Adam optimizer.
 Following [21], we train an ensemble of $3$ reward models.
 Each reward model includes a $3$-layer feed-forward neural network with $256$ hidden units.
-The reward model training uses cross-entropy loss with a learning rate of $3\times e^{-4}$ and Adam optimizer.
-2) Policy optimization: The offline algorithms include CQL and IQL.
-In CQL algorithm, the actor network $\pi_{\theta}$ and critic network $Q_{\omega}$ adopt a $3$-layer feed-forward neural network with learning rate of $1\times e^{-4}$ and $3\times e^{-4}$, respectively.
+The reward model training uses cross-entropy loss with a learning rate of $3\times 10^{-4}$ and Adam optimizer.
+Policy optimization: The offline algorithms include CQL [4] and IQL [25].
+In CQL algorithm, the actor network $\pi_{\theta}$ and critic network $Q_{\omega}$ adopt a $3$-layer feed-forward neural network with learning rate of $1\times 10^{-4}$ and $3\times 10^{-4}$, respectively.
 The cql weight and temperature are set as $5.0$ and $1.0$ for all tasks, respectively.
-In IQL algorithm, the actor network and critic network adopt a $2$-layer neural network with learning rate of $3\times e^{-4}$ and $3\times e^{-4}$, respectively.
-The expectile and temperature are set as $0.7$ and $3.0$ for mujoco tasks and $0.8$ and $3.0$ for adroit tasks.
-3) Parameter tuning: Theorem 1 indicates that the performance of reward model is influenced by the selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$ and the quality of the generated data.
-Therefore, the key hyperparameters should be tuned for LEASE including: rollout horizon $H\in\{10,20,100,200\}$, probability confidence $\kappa_{p}(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})\in\{0.75,0.85,0.90,0.95,0.99\}$ and uncertainty variance $\kappa_{\tau}(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})\in\{0.05,0.08\}$.
-Table II shows the above hyperparameters for LEASE under CQL and IQL algorithms.
-Figure 4: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) hopper-medium-expert and (b) halfcheetah-medium-expert datasets, where the value predicted by the trained reward model and ground-truth reward value are both normalized to $[0,1]$.
-From left to right are methods LEASE, FEWER and FRESH, respectively.
-VII-B Results for Reward Model
-The effect of the number of preference data for preference model.
-Learning an accurate reward is difficult in the offline PbRL setting when preference labels are scarce.
+In the IQL algorithm, the actor network and critic network adopt a $2$-layer neural network with learning rate of $3\times 10^{-4}$ and $3\times 10^{-4}$, respectively.
+Similar to [25], the expectile and temperature are set as $0.7$ and $3.0$ for mujoco tasks and $0.8$ and $3.0$ for adroit tasks.
+[…]
+TABLE II: Comparison results of pseudo-labeling accuracy between using the selection mechanism (Ours), only using confidence (No-U), only using uncertainty (No-C), and not using the selection mechanism (No-C-U).
+Task Name Ours No-U No-C No-C-U
+pen-expert $\mathbf{87.3\pm 0.5}$ $85.9\pm 0.8$ $86.7\pm 0.6$ $86.8\pm 0.5$
+door-expert $\mathbf{89.3\pm 0.6}$ $87.8\pm 0.2$ $86.0\pm 0.0$ $85.9\pm 0.8$
+hammer-expert $\mathbf{85.5\pm 0.2}$ $84.4\pm 0.4$ $82.5\pm 1.2$ $83.6\pm 0.5$
+Average $\mathbf{87.4\pm 0.4}$ $86.0\pm 0.5$} $85.1\pm 0.6$} $85.4\pm 0.6$}
+TABLE III: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds.
+$\pm$ captures the standard deviation over seeds.
+Bold indicates the performance within 2% of the best-performing algorithm for each offline algorithm.
+Task Name CQL [4] IQL [25]
+URLHF FEWER LEASE URLHF FEWER LEASE
+walker2d-m $6.0\pm 0.9 77$$7.4\pm 0.6$ $\mathbf{78.4\pm 0.9}$ $\mathbf{78.4\pm 0.5}$ $71.8\pm 0.8$ $73.0\pm 2.1$
+walker2d-m-e $92.8\pm 22.4$ $77.7\pm 0.3$ $\mathbf{98.6\pm 18.1}$ $\mathbf{109.4\pm 0.1}$ $105.2\pm 3.1$ $\mathbf{108.1\pm 0.5}$
+hopper-m $4.7\pm 3.4 55$$5.8\pm 2.8$ $\mathbf{56.5\pm 0.6}$ $50.8\pm 4.2$ $\mathbf{56.7\pm 2.6}$ $\mathbf{56.0\pm 0.5}$
+hopper-m-e $\mathbf{57.4\pm 4.9}$ $53.6\pm 0.9$ $56.4\pm 0.8$ $\mathbf{94.3\pm 7.4}$ $6.0\pm 1.4 55$$5.9\pm 1.9$
+halfcheetah-m $\mathbf{43.4\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.3\pm 0.2}$ $42.2\pm 0.1$ $\mathbf{43.0\pm 0.3}$
+halfcheetah-m-e $\mathbf{62.7\pm 7.1}$ $48.3\pm 0.7$ $53.2\pm 3.1$ $\mathbf{91.0\pm 2.3}$ $59.1\pm 4.9$ $60.2\pm 1.2$
+Mujoco Average $\mathbf{64.5\pm 6.5}$ $59.4\pm 0.9$ $\mathbf{64.4\pm 4.0}$ $\mathbf{77.9\pm 2.5}$ $5.2\pm 2.2 66$$6.4\pm 1.1$
+pen-human $\mathbf{9.8\pm 14.1}$ $0.5\pm 3.0$ $3.8\pm 4.6$ $50.2\pm 15.8$ $67.3\pm 10.0$ $\mathbf{75.6\pm 3.3}$
+pen-expert $\mathbf{138.3\pm 5.2}$ $128.1\pm 0.7$ $130.6\pm 1.8$ $\mathbf{132.9\pm 4.6}$ $04.1\pm 12.9 11$$13.8\pm 6.3$
+door-human $\mathbf{4.7\pm 5.9}$ $0.2\pm 1.0$ $\mathbf{4.7\pm 8.8}$ $3.5\pm 3.2$ $4.0\pm 2.5$ $\mathbf{5.7\pm 0.8}$
+door-expert $\mathbf{103.9\pm 0.8}$ $103.0\pm 0.9$ $\mathbf{103.2\pm 0.7}$ $\mathbf{105.4\pm 0.4}$ $104.5\pm 0.6$ $\mathbf{105.2\pm 0.2}$
+hammer-human $\mathbf{0.9\pm 0.3}$ $0.3\pm 0.0$ $0.3\pm 0.0$ $\mathbf{1.4\pm 1.0}$ $1.2\pm 0.2$ $\mathbf{1.4\pm 0.9}$
+hammer-expert $120.2\pm 6.8$ $124.1\pm 2.1$ $\mathbf{126.3\pm 1.2}$ $\mathbf{127.4\pm 0.2}$ $125.2\pm 2.3$ $\mathbf{126.3\pm 0.1}$
+Adroit Average $\mathbf{63.0\pm 5.5}$ $59.4\pm 1.3$ $61.5\pm 2.9$ $\mathbf{70.1\pm 4.2}$ $67.1\pm 4.8$ $\mathbf{71.3\pm 1.9}$
+VI-B Results for Reward Model
+VI-B1 The effect of the selection mechanism for pseudo-label accuracy
+We employ a pre-trained reward model to annotate unlabeled preference data generated during the process.
+This paper proposes a novel screening mechanism to select data samples with both high confidence and low uncertainty, thereby enhancing the accuracy of pseudo-labels.
+The pseudo-label is generated by a preference model, which is constructed based on the reward model as formulated in Eq.
+(2).
+Table II compares the pseudo-label accuracy under four different conditions: 1) using the proposed screening mechanism (combining both confidence and uncertainty thresholds), 2) employing only the confidence threshold (No-U), 3) applying only the uncertainty threshold (No-C), and 4) without any screening mechanism (No-C-U).
+The accuracy of the pseudo-label generated by the reward model is evaluated on all preference datasets.
+Figure 4: The relationship between pseudo-labeling accuracy and the number of preference dataset $N_{l}$.
+Table II shows that the proposed screening mechanism significantly improves pseudo-label accuracy, i.e., reduces error $\eta$ defined in Assumption 2.
+The pseudo-label accuracy is improved by $4.5\%$, $2.8\%$, and $6.9\%$ for pen-expert, door-expert, and hammer-expert, respectively.
+Moreover, we also analyze the effect of the number of preference data on the preference model.
 Fig.
-3 illustrates the relationship between the prediction accuracy of preference model without data augmentation and the number of preference data $N_{l}$, where preference model is based on reward model (Eq.
-(2)).
-Here, we use the last $200$ of $2000$ preference data that are not seen in training stage as evaluation dataset.
+4 illustrates the relationship between the prediction accuracy of the preference model without data augmentation and the number of preference data $N_{l}$.
+Here, we use the last $200$ of $2000$ preference data that are not seen in the training stage as the evaluation dataset.
 It indicates that as the number of preference data $N_{l}$ increases, the prediction accuracy of the preference model shows an upward trend.
 This also validates Theorem 1 that increasing the number of samples $N_{l}$ can reduce the generalization error.
-TABLE III: Comparison results of pseudo-labeling accuracy and agent performance between using confidence and uncertainty and only using confidence.
-The accuracy of pseudo-label generated by reward model is evaluated on all preference dataset.
-Selecting Mechanism pen-expert door-expert hammer-expert
-label accuracy agent performance label accuracy agent performance label accuracy agent performance
-Confidence and uncertainty $87.3\pm 0.5$ $\mathbf{132.5\pm 2.3}$ $89.3\pm 0.6$ $\mathbf{103.2\pm 0.7}$ $85.5\pm 0.2$ $\mathbf{126.3\pm 1.2}$
-Only confidence $85.9\pm 0.8$ $126.8\pm 0.1$ $87.8\pm 0.2$ $101.1\pm 2.2$ $84.4\pm 0.4$ $118.2\pm 9.6$
-The effect of data augmentation for the performance of reward model.
-We compare the accuracy of reward model before and after data augmentation under fewer preference dataset.
+VI-B2 The effect of data augmentation on the performance of the reward model
+To answer question $\mathbf{Q_{1}}$, we compare the accuracy of the reward model before and after data augmentation under a smaller preference dataset.
 Fig.
 3 shows the comparison results between prediction and ground truth of reward, where the predicted and true rewards are both normalized to $[0,1]$.
-We randomly sample $500$ data from unlabeled datasets that are not seen in training stage for evaluation.
-This figure indicates the linear relationship between reward predicted by LEASE and ground truth is better than that of other two methods, where method not using selecting mechanism is denoted as FRESH.
-The prediction of reward model learned by method FRESH is very narrow and the accuracy is greatly reduced compared with FEWER.
-This is because generated data of FRESH are not selected by $f(\sigma_{0},\sigma_{1})$, which causes substantial errors for the labels of generated data and leads to the collapse of training.
-This validates that the performance of reward model can be improved through data augmentation and selecting mechanism, where the designed selecting mechanism has a greater impact on reward model performance.
-TABLE IV: The performance of the learned transition model, where error is calculated by the sum of the mean square values of predicted and true value in each dimension.
-Task Name walker2d-m-e hopper-m-e halfcheetah-m-e
-Error $6.57\pm 0.10$ $0.38\pm 0.01$ $16.19\pm 0.48$
-Task Name pen-expert door-expert hammer-expert
-Error $1.05\pm 0.01$ $0.04\pm 0.01$ $0.58\pm 0.03$
-The effect of uncertainty for reward model and agent performance.
-Table III compares the accuracy of pseudo-labels and performance between the proposed method LEASE and the confidence-only method.
-The latter method is the Surf [3] that only uses confidence for screening data.
-This table validates that LEASE outperforms the confidence-only approach in pseudo-labels accuracy and agent performance.
-The performance improved by $4.5\%$, $2.8\%$ and $6.85\%$ for pen-expert, door-expert and hammer-expert, respectively.
-Additionally, there is a difference between the accuracy of the preference model $P(\sigma_{0}\succ\sigma_{1})$ (pseudo-label accuracy) and the accuracy of the reward model $R(s,a)$ (See Eq.
-(2)).
-A decline in the accuracy of the reward model may not necessarily lead to a significant drop in the accuracy of the preference model, as the preference model is related to the summation of reward model over trajectories.
-VII-C Results for Transition model
-The accuracy of transition model.
-Table IV shows the accuracy result for the learned transition model, where the accuracy of transition model is measured by the sum of the mean square values of the predicted value and the true value in each dimension.
-Note that, for walker2d and halfcheetah tasks, the error of the learned transition model is greatly larger than hopper tasks in mujoco environment.
-This is mainly because the state space of walker2d and hopper is larger than hopper, and the physical model of them are more complex.
-TABLE V: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds.
-$\pm$ captures the standard deviation over seeds.
-Bold indicates the performance within 2% of the best performing algorithm for each offline algorithm.
-Task Name CQL [4] IQL [25]
-URLHF FEWER LEASE URLHF FEWER LEASE
-walker2d-m $76.0\pm 0.9$ $77.4\pm 0.6$ $\mathbf{78.4\pm 0.9}$ $\mathbf{78.4\pm 0.5}$ $71.8\pm 0.8$ $74.6\pm 1.8$
-walker2d-m-e $92.8\pm 22.4$ $77.7\pm 0.3$ $\mathbf{98.6\pm 18.1}$ $\mathbf{109.4\pm 0.1}$ $105.2\pm 3.1$ $\mathbf{108.1\pm 0.5}$
-hopper-m $54.7\pm 3.4$ $55.8\pm 2.8$ $\mathbf{56.5\pm 0.6}$ $50.8\pm 4.2$ $\mathbf{56.7\pm 2.6}$ $\mathbf{56.0\pm 0.5}$
-hopper-m-e $\mathbf{57.4\pm 4.9}$ $53.6\pm 0.9$ $56.4\pm 0.8$ $\mathbf{94.3\pm 7.4}$ $56.0\pm 1.4$ $55.9\pm 1.9$
-halfcheetah-m $\mathbf{43.4\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.3\pm 0.2}$ $42.2\pm 0.1$ $\mathbf{43.0\pm 0.3}$
-halfcheetah-m-e $\mathbf{62.7\pm 7.1}$ $48.3\pm 0.7$ $53.2\pm 3.1$ $\mathbf{91.0\pm 2.3}$ $59.1\pm 4.9$ $62.4\pm 1.4$
-Mujoco Average $\mathbf{64.5\pm 6.5}$ $59.4\pm 0.9$ $\mathbf{64.4\pm 4.0}$ $\mathbf{77.9\pm 2.5}$ $65.2\pm 2.2$ $66.7\pm 1.0$
-pen-human $\mathbf{9.8\pm 14.1}$ $0.5\pm 3.0$ $3.8\pm 4.6$ $50.2\pm 15.8$ $67.3\pm 10.0$ $\mathbf{75.6\pm 3.3}$
-pen-expert $\mathbf{138.3\pm 5.2}$ $128.1\pm 0.7$ $132.5\pm 2.3$ $\mathbf{132.9\pm 4.6}$ $104.1\pm 12.9$ $113.8\pm 6.3$
-door-human $\mathbf{4.7\pm 5.9}$ $0.2\pm 1.0$ $\mathbf{4.7\pm 8.8}$ $3.5\pm 3.2$ $4.0\pm 2.5$ $\mathbf{5.9\pm 0.5}$
-door-expert $\mathbf{103.9\pm 0.8}$ $103.0\pm 0.9$ $\mathbf{103.2\pm 0.7}$ $\mathbf{105.4\pm 0.4}$ $104.5\pm 0.6$ $\mathbf{105.2\pm 0.2}$
-hammer-human $\mathbf{0.9\pm 0.3}$ $0.3\pm 0.0$ $0.3\pm 0.0$ $1.4\pm 1.0$ $1.2\pm 0.2$ $\mathbf{1.7\pm 0.4}$
-hammer-expert $120.2\pm 6.8$ $124.1\pm 2.1$ $\mathbf{126.3\pm 1.2}$ $\mathbf{127.4\pm 0.2}$ $125.2\pm 2.3$ $\mathbf{126.3\pm 0.1}$
-Adroit Average $\mathbf{63.0\pm 5.5}$ $59.4\pm 1.3$ $61.8\pm 3.0$ $\mathbf{70.1\pm 4.2}$ $67.1\pm 4.8$ $\mathbf{71.4\pm 1.8}$
-TABLE VI: The comparison results between the performance using selecting mechanism and that not using. The latter method is denoted as FRESH. $\uparrow$ denotes the improvement of performance.
+We randomly sample $500$ data points from unlabeled datasets that are not seen in the training stage for evaluation.
+This figure indicates that the linear relationship between reward predicted by LEASE and ground truth is better than that of the other two methods, where the method not using a selecting mechanism is denoted as FRESH.
+The prediction of the reward model learned by method FRESH is very narrow, and the accuracy is greatly reduced compared with FEWER.
+This is because the generated data of FRESH are not selected by $f(\sigma_{0},\sigma_{1})$, which causes substantial errors for the labels of generated data and leads to the collapse of training.
+This validates that the performance of the reward model can be improved through data augmentation and a selection mechanism, where the designed selection mechanism has a greater impact on the reward model’s performance.
+VI-C Results for Policy Performance
+TABLE IV: The comparison results between the performance using the selection mechanism and that not using it. The latter method is denoted as FRESH. $\uparrow$ denotes the improvement of performance.
 Task Name walker2d-m hopper-m halfcheetah-m walker2d-m-e hopper-m-e halfcheetah-m-e
@@
 IQL FRESH $72.7\pm 4.1$ $53.5\pm 0.8$ $42.5\pm 0.1$ $105.3\pm 1.0$ $53.9\pm 0.1$ $54.5\pm 2.2$
-LEASE $\mathbf{74.6\uparrow(2.6\%)}$ $\mathbf{56.0\uparrow(4.5\%)}$ $\mathbf{43.0\uparrow(1.3\%)}$ $\mathbf{108.1\uparrow(2.7\%)}$ $\mathbf{55.9\uparrow(3.7\%)}$ $\mathbf{62.4\uparrow(14.6\%)}$
-TABLE VII: The prediction result of preference model for different rollout length $H$.
-Pretrained represents the accuracy of preference model trained with fewer preference data, and updated represents that updated with generated data.
-Task name walker2d-m-e hopper-m-e halfcheetah-m-e
-Pretrained $0.85$ $0.77$ $0.72$
-Updated $H=10$ $\mathbf{0.85}$ $0.74$ $\mathbf{0.78}$
-$H=100$ $0.84$ $\mathbf{0.79}$ $0.74$
-$H=200$ $0.83$ $0.76$ $0.77$
-The effect of rollout length $H$ of transition model for preference model.
-Table VII shows the prediction accuracy of the trained preference model for different rollout length $H$.
-Here, we take CQL as example and select the last $200$ of $2000$ preference data as evaluation dataset that are not seen in in reward training stage.
-The accuracy is calculated through the gap between predicted preference and true preference based on two trajectories.
-It shows that long horizon easily brings low performance of reward model since the cumulative error of the trained transition model, but long horizon is beneficial to enhance the performance of reward model under some conditions.
-The accuracy of preference model can be improved through data augmentation under most tasks.
-Note that the rollout length $H$ is equal to the length of generated unlabeled data $L$.
-The $H$ is influenced by the accuracy of trained transition model.
-As $H$ increases, the accuracy of prediction decreases and large error will bring the unstable agent training.
-However, the long horizon of trajectory can better describe the preference of human feedback and can reduce sample complexity (See Eq.
-(15)).
-Additionally, the results in table VII can not completely reflect the generalization performance of reward model since the evaluation data are limited and there is difference between preference model and reward model (See Eq.
-(2)).
-TABLE VIII: Results of the effect of the learned transition model for agent performance.
-No. hopper-medium pen-expert
-Error Performance Error Performance
-$1$ $0.35\pm 0.01$ $56.5\pm 0.60$ $1.05\pm 0.01$ $132.5\pm 2.3$
-$2$ $0.49\pm 0.04$ $54.8\pm 0.85$ $1.42\pm 0.05$ $126.4\pm 8.6$
-$3$ $1.19\pm 0.05$ $52.9\pm 0.92$ $2.36\pm 0.24$ $87.7\pm 41.8$
-The effect of the accuracy of transition model for agent performance.
-Table VIII shows the detailed results for the effect of transition model accuracy for agent performance, where the smaller the model error means the higher the accuracy.
-The change in transition model error is achieved by changing the number of training steps.
-The base offline RL algorithm is CQL.
-It shows that the lower the accuracy of the transition model, the poorer performance of the agent.
-VII-D Results for Policy Performance
-The comparison results on benchmark tasks.
-Here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms.
-Table V gives the results of the average normalized score with standard deviation, where m and m-e denote medium and medium-expert respectively.
-The offline PbRL algorithm under sufficient feedback (URLHF) are obtained from paper [21].
-This table shows that LEASE can greatly improve performance under fewer preference data compared with FEWER, and achieve comparable performance to URLHF that demand large human feedback.
-This validates the content of Theorem 2 that LEASE can reduce the performance gap caused by reward model under fewer preference data and has high sample efficiency.
-TABLE IX: The comparison results for the D4RL tasks under different number of preference data $N_{l}$. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score.
+LEASE $\mathbf{73.0\uparrow(0.4\%)}$ $\mathbf{56.0\uparrow(4.5\%)}$ $\mathbf{43.0\uparrow(1.3\%)}$ $\mathbf{108.1\uparrow(2.7\%)}$ $\mathbf{55.9\uparrow(3.7\%)}$ $\mathbf{60.2\uparrow(10.5\%)}$
+VI-C1 The comparison results on benchmark tasks
+To answer question $\mathbf{Q_{2}}$, here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms on six types of tasks and 12 datasets in total.
+Table III presents a comparative analysis of he results of the average normalized score with standard deviation under three distinct scenarios: 1) performance with extensive preference data (URLHF), 2) performance with limited data without employing data augmentation techniques (FEWER), and 3) performance with limited data while incorporating data augmentation (our proposed method LEASE).
+In this table, m and m-e represent medium and medium-expert, respectively.
+The results of the offline PbRL algorithm under sufficient feedback (URLHF) are obtained from the paper [21].
+As demonstrated in Table III, LEASE exhibits significantly superior performance compared to FEWER.
+This empirical evidence substantiates that data augmentation can effectively enhance the agent’s capability.
+Furthermore, compared to the performance achieved with extensive preference data URLHF, LEASE attains comparable results in most cases.
+However, the proposed method exhibits slightly inferior performance on specific tasks, such as hopper-medium-expert and halfcheetah-medium-expert datasets.
+This limitation primarily stems from the excessively narrow distribution of the expert dataset, which consequently induces a constrained distribution in the generated preference data.
+As a result, the reward model’s performance cannot be effectively enhanced through data augmentation, thereby restricting the potential improvement in agent performance.
+However, overall, the results in Table III can validate the content of Theorem 2, which shows that LEASE can reduce the performance gap caused by the reward model and improve the agent’s performance with fewer preference data.
+TABLE V: The comparison results for the D4RL tasks under different number of preference data $N_{l}$. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score.
 CQL [4] IQL [25]
@@
 walker2d-medium $77.7\pm 1.3$ $\mathbf{78.4\pm 0.9}$ pen-human $71.9\pm 7.4$ $\mathbf{75.6\pm 3.3}$
-walker2d-medium-expert $98.0\pm 18.6$ $\mathbf{98.6\pm 18.1}$ pen-expert $102.5\pm 12.8$ $\mathbf{113.8\pm 6.3}$
-hopper-medium $\mathbf{56.8\pm 1.5}$ ${56.5\pm 0.6}$ door-human $4.4\pm 1.2$ $\mathbf{5.9\pm 0.5}$
+[…]
+hopper-medium $\mathbf{56.8\pm 1.5}$ $56.5\pm 0.6$} door-human $4.4\pm 1.2$ $\mathbf{5.7\pm 0.8}$
 hopper-medium-expert $54.5\pm 1.2$ $\mathbf{56.4\pm 0.8}$ door-expert $\mathbf{105.2\pm 0.1}$ $\mathbf{105.2\pm 0.2}$
-halfcheetah-medium $43.4\pm 0.4$ $\mathbf{43.5\pm 0.1}$ hammer-human $1.2\pm 0.4$ $\mathbf{1.7\pm 0.4}$
+halfcheetah-medium $43.4\pm 0.4$ $\mathbf{43.5\pm 0.1}$ hammer-human $1.2\pm 0.4$ $\mathbf{1.4\pm 0.9}$
 halfcheetah-medium-expert $51.0\pm 0.8$ $\mathbf{53.2\pm 3.1}$ hammer-expert $\mathbf{126.4\pm 0.1}$ $126.3\pm 0.1$
-Mujoco Average $63.6\pm 4.0$ $\mathbf{64.4\pm 4.0}$ Adroit Average $68.6\pm 3.7$ $\mathbf{71.4\pm 1.8}$
-The effect of selecting mechanism for agent performance.
-Table VI compares the agent performance between the proposed method LEASE and the method not using selecting mechanism (We denote this method as FRESH).
-This table indicates that the application of selecting mechanism $f(\sigma_{0},\sigma_{1})$ can efficiently improve agent performance.
-This results also further validates the effectiveness and advantages of the designed selecting mechanism.
-The effect of the number of preference data $N_{l}$ for agent performance.
+Mujoco Average $63.6\pm 4.0$ $\mathbf{64.4\pm 4.0}$ Adroit Average $68.6\pm 3.7$ $\mathbf{71.3\pm 1.9}$
+VI-C2 The effect of selection mechanism on agent performance
+To answer question $\mathbf{Q_{3}}$, we compare the agent performance between the proposed method LEASE and the method not using the selection mechanism (We denote this method as FRESH).
 Fig.
-4 indicates that the number of preference data $N_{l}$ influences the accuracy of reward model directly.
+3 shows that the performance of the reward model can be significantly improved by using the selection mechanism.
+Table IV compares the results between LEASE and FRESH under three locomotion tasks based on CQL and IQL offline RL algorithms.
+This table indicates that the application of the selection mechanism $f(\sigma_{0},\sigma_{1})$ can efficiently improve agent performance, and the average performance of the policy learned by LEASE can be improved by $5.6\%$ compared with not using the selection mechanism.
+These results also further validate the effectiveness and advantages of the designed selection mechanism.
+VI-C3 The effect of the number of preference data $N_{l}$ for agent performance
+Fig.
+4 indicates that the number of preference data $N_{l}$ influences the accuracy of the reward model directly.
 URLHF [21] trains reward model with $2000$ preference dataset.
-Our method LEASE aims to achieve comparable performance with URLHF under fewer preference dataset.
-Table IX shows the agent performance when the number of preference dataset is $20$ and $100$, where we use CQL algorithm for mujoco tasks and IQL algorithm for adroit tasks.
-This table indicates LEASE still can improve agent performance under very fewer preference dataset.
+This paper aims to achieve competitive performance with limited preference data, that is, achieve comparable performance with URLHF under a smaller preference dataset.
+We further investigate the impact of the number of preference data on the agent’s performance.
+Table V compares the agent’s performance when the number of preference datasets is $20$ and $100$, using the CQL algorithm for MuJoCo tasks and the IQL algorithm for Adroit tasks.
+This table indicates LEASE can still improve agent performance under a very small preference dataset.
 Notably, the performance gap between $100$ preference data and $20$ preference data is not much.
-This indicates that LEASE have potential to perform well under more fewer preference dataset.
-VIII Discussion
-VIII-A Discussion for Offline PbRL Theory
-In offline RL field, high sample efficiency refers that the agent can achieve comparable performance under fewer data compared with the performance under large data.
+This indicates that LEASE has the potential to perform well under a smaller preference dataset.
+VI-D Parameter Analysis
+TABLE VI: The results of parameter analysis for the confidence threshold $\kappa_{p}$ and the uncertainty threshold $\kappa_{\tau}$.
+CQL $\kappa_{\tau}=0.05$ $\kappa_{\tau}=0.08$
+$\kappa_{p}=0.85$ $\kappa_{p}=0.95$ $\kappa_{p}=0.85$ $\kappa_{p}=0.95$
+walker2d-m $\mathbf{78.4\pm 0.9}$ $70.6\pm 1.8$ $72.4\pm 1.1$ $70.3\pm 2.1$
+pen-expert $127.0\pm 3.4$ $\mathbf{130.6\pm 1.8}$ $122.9\pm 1.8$ $125.8\pm 2.4$
+IQL $\kappa_{\tau}=0.05$ $\kappa_{\tau}=0.08$
+$\kappa_{p}=0.85$ $\kappa_{p}=0.95$ $\kappa_{p}=0.85$ $\kappa_{p}=0.95$
+walker2d-m $72.6\pm 0.1$ $69.1\pm 1.5$ $\mathbf{73.0\pm 2.1}$ $68.9\pm 3.3$
+[…]
+The selection mechanism $f(\sigma_{0},\sigma_{1})$ for the reward model has two critical parameters: the confidence threshold $\kappa_{p}$ and the uncertainty threshold $\kappa_{\tau}$.
+The selection of two parameters depends on multiple factors, including the complexity of the task’s reward function, the performance of the pre-trained reward model, and the specific offline RL algorithm employed (since the collected preference data is inherently linked to the policy learned by the algorithm).
+Given the substantial computational cost inherent in RL training, exhaustive parameter tuning is often impractical.
+To mitigate this, our work systematically analyzes the impact of key parameters on task performance across multiple algorithms, offering practical guidance for their configuration.
+Table VI presents a comparative analysis of the performance of different offline RL algorithms (CQL and IQL) under varying parameter configurations for both locomotion and manipulation tasks.
+For the uncertainty threshold $\kappa_{\tau}$, the results demonstrate that CQL exhibits a smaller uncertainty threshold compared to IQL.
+This phenomenon can be attributed to CQL’s inherent conservatism, which leads to a narrower data distribution when the learned policy interacts with the environmental model (i.e., the generated data is more concentrated within a specific region).
+Consequently, the reward model exhibits lower prediction uncertainty, allowing for the selection of a reduced uncertainty threshold.
+For the confidence threshold $\kappa_{p}$, manipulation tasks require a higher confidence threshold than locomotion tasks.
+This discrepancy arises from the simpler reward structure of manipulation tasks, which typically rely on discrete variables (See Fig.
+A.1 in Appendix D).
+As a result, the pre-trained reward model achieves higher accuracy in manipulation tasks, allowing for the use of a more stringent confidence threshold to enhance the quality of generated data.
+For additional experiments of the LEASE—such as evaluations of the reward model on alternative tasks, its application within model-based RL algorithms, and an extended comparative analysis—please refer to Appendix D.
+VII Discussion
+VII-A Discussion for Offline PbRL Theory
+In the offline RL field, high sample efficiency refers to the agent’s ability to achieve comparable performance under fewer data compared with the performance under large data.
 In this paper, the data refers to the preference dataset.
-The labeled preference dataset, each trajectory of length $L$, is collected through real-time human feedback under policy $\mu$, which demands tremendous human effort, thus the collected cost of preference data is higher than fixed offline data.
-The unlabeled dateset is generated through trained transition without real-time interaction under learned policy $\pi^{t}$ at time $t$.
+The labeled preference dataset, each trajectory of length $L$, is collected through real-time human feedback under policy $\mu$, which demands tremendous human effort; thus, the collected cost of preference data is higher than fixed offline data.
+The unlabeled dataset is generated through a trained transition without real-time interaction under the learned policy $\pi^{t}$ at time $t$.
 Through data augmentation, the sample efficiency can be significantly reduced.
-The performance gap caused by reward can be reduced to $\varpi$ under fewer labeled dataset.
-Zhan et al.[18] also developed systematical theory for offline PbRL and introduced the concentrability coefficient $\mathscr{C}_{r}(\pi)$ for PbRL, defined as
+The performance gap caused by reward can be reduced to $\varpi$ under a smaller labeled dataset.
+Zhan et al.[18] also developed a systematic theory for offline PbRL, but did not focus on the generalization performance of the reward model.
+Moreover, they also introduced the concentrability coefficient $\mathscr{C}_{r}(\pi)$ for PbRL, defined as
 $\max\Bigg\{0,\sup_{r}\frac{\mathbb{E}_{\sigma^{0}\sim\pi,\sigma^{1}\sim\pi_{\text{ref}}}\big[r^{*}(\sigma^{0},\sigma^{1})-{r}(\sigma^{0},\sigma^{1})\big]}{\sqrt{\mathbb{E}_{\sigma^{0}\sim\mu_{0},\sigma^{1}\sim\mu_{1}}\big|r^{*}(\sigma^{0},\sigma^{1})-{r}(\sigma^{0},\sigma^{1})\big|^{2}}}\Bigg\},$
@@
 However, it is based on trajectory and is difficult to combine with other offline RL theories.
-The proposed concentrability coefficient in our paper defined in Eq.
-(13) is based on state-action pairs.
-Moreover, the performance gap of offline PbRL between behavior policy and learned policy is influenced by offline algorithm itself and the performance of the learned reward (preference) model.
-However, [18] fails to consider the gap caused by offline algorithm itself and the sampling error.
-The theory developed in our paper can be easily combined with other offline algorithm.
-The method in [18] can learn $\varpi$-optimal policy with a sample complexity of
+The proposed concentrability coefficient in our paper, defined in Eq.
+(13), is based on state-action pairs.
+Moreover, the performance gap of offline PbRL between the behavior policy and learned policy is influenced by the offline algorithm itself and the performance of the learned reward (preference) model.
+However, [18] fails to consider the gap caused by the offline algorithm itself and the sampling error.
+The theory developed in our paper can be easily combined with another offline algorithm.
+Then, the method in [18] can learn $\varpi$-optimal policy with a sample complexity of
 $N=\widetilde{\mathcal{O}}\Bigg(\frac{c^{2}k^{2}\mathscr{C}_{r}^{2}(\widehat{\pi})}{\varpi^{2}}\log\left(\frac{\mathcal{N}_{r}(1/N)}{\delta}\right)\Bigg),$ (17)
@@
 (15) contains more useful information.
-For example, the sample complexity is related to the length of preference data and the range of reward function, and can be reduced when the learned policy is closed to behavior policy.
-VIII-B Discussion for Broader Impacts
+For example, the sample complexity is related to the length $L$ of preference data and the range $R_{max}$ of the reward function, and can be reduced when the learned policy is close to the behavior policy $\mu$.
+VII-B Discussion for Broader Impacts
 Actual application.
-In actual scenarios, especially for human-in-loop-control, such as exoskeleton robot assistance or rehabilitation, designing a high sample efficient RL algorithm is greatly significant.
-Firstly, many rewards are difficult to describe in mathematical terms in some situation, such as the comfort of interaction.
-Human feedback or preference is the better way to reflect the above indexes.
-In addition, in human in-the-loop control, obtaining preference data through interaction can easily cause fatigue, and inappropriate interaction may cause damage to human.
-It is a necessity to learn reward model from limited preference dataset.
-Therefore, PbRL has a great potential to improve control performance in some scenarios where the reward function is difficult to describe and human is in the control loop.
+In actual scenarios, especially for human-in-loop control, such as exoskeleton robot assistance or rehabilitation, designing a highly sample-efficient RL algorithm is greatly significant.
+Firstly, many rewards are difficult to describe in mathematical terms in some situations, such as the comfort of interaction.
+Human feedback or preference is a better way to reflect the above indices.
+In addition, in human-in-the-loop control, obtaining preference data through interaction can easily cause fatigue, and inappropriate interaction may cause damage to humans.
+It is a necessity to learn a reward model from a limited preference dataset.
+Then, for robot assistance or rehabilitation, the goal of the robot is to reduce human metabolic cost and guarantee the comfort of interaction.
+However, the reward model for this goal is difficult to design.
+Therefore, PbRL has a great potential to improve control performance in some scenarios where the reward function is difficult to describe and humans are in the control loop.
 Theoretical study.
@@
 Moreover, the theory of LEASE is based on state-action pairs, which is consistent with most offline algorithms theory.
-Therefore, the theory developed in this paper can be easily combined with other offline algorithms theory and easily used to build theory of policy improvement guarantee, which can provide the theoretical basis for offline PbRL and facilitate the development of offline PbRL theory.
-IX Conclusion
+Therefore, the theory developed in this paper can be easily combined with other offline algorithm theories and easily used to build the theory of policy improvement guarantee, which can provide the theoretical basis for offline PbRL and facilitate the development of offline PbRL theory.
+VIII Conclusion
 This paper proposes a novel offline PbRL algorithm (LEASE) with high sample efficiency.
-LEASE can achieve comparable performance under fewer preference dataset.
-By selecting high confidence and low variance data, the stability and accuracy of reward model are guaranteed.
-Moreover, this paper provides the theoretical analysis for LEASE, including generalization of reward model and policy improvement guarantee.
-This theory can be easily connected with other offline algorithms.
-The theoretical and experimental results demonstrate that the data selecting mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve performance of reward model and the performance learned by LEASE can be guaranteed under fewer preference dataset.
-However, there is still the gap between the true and the learned reward model.
-Future works can focus on how to further reduce this gap under the limited preference data or how to achieve conservative estimation for state-action pairs where the learned reward model predicts inaccurately.
+LEASE can achieve comparable performance under fewer preference datasets.
+By selecting high-confidence and low-variance data, the stability and accuracy of the reward model are guaranteed.
+Moreover, this paper provides the theoretical analysis for LEASE, including the generalization of the reward model and policy improvement guarantee.
+The theoretical and experimental results demonstrate that the data selection mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve the performance of the reward model, and the performance learned by LEASE can be guaranteed under fewer preference datasets.
+Future work can focus on reducing this gap further under limited preference data or on achieving conservative estimation for state-action pairs where the learned reward model predicts inaccurately.
 Appendix A Related Proofs
@@
 The Proof for Proposition 1: This proof follows the previous work [29]. Obviously, the largest pseudo-labeling error $\eta$ in Assumption 2 includes exactly two types of pseudo-labeling error:
-$\eta_{1}=\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]},$ (A.1)
-$\eta_{2}=\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]},$
+[…]
+[…]
 where $\eta_{1}$ represents the error ratio of classifying the first category as the second category, and $\eta_{2}$ is the error ratio of classifying the second category as the first category.
-Then, we prove the gap between the unlabeled loss with pseudo-label ^$\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ from two sides: ^$\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$ and ^$\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$.
-Notably, we ignore the selecting mechanism in the below proof since it doesn’t influence the proof result.
-Step 1: Proving upper bound : ^$\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$.
-^$\widehat{\mathcal{L}}_{u}(\psi)$
+Then, we prove the gap between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ from two sides: $\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$ and $\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$.
+Notably, we ignore the selection mechanism in the proof below since it doesn’t influence the proof result.
+Step 1: Proving upper bound : $\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$.
+$\widehat{\mathcal{L}}_{u}(\psi)$
 $=$ $\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$
-$=$ $-\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5\big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
-$+\mathbbm{1}\big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5\big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
+[…]
+[…]
 $\leq$ $-\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\big[y^{j}=0\big]\log P_{\psi}\big(\sigma_{0}^{j}\succ\sigma_{1}^{j}\big)$
 $+\mathbbm{1}\big[y^{j}=1\big]\log P_{\psi}\big(\sigma_{1}^{j}\succ\sigma_{0}^{j}\big)$
-$+\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
-$+\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
-$\leq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta_{1}\max_{j}\Big{\{}-\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}\Big{\}}$
+[…]
+[…]
+$\leq$ $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta_{1}\max_{j}\Big\{-\log P_{\psi}\big(\sigma_{0}^{j}\succ\sigma_{1}^{j}\big)\Big\}$
 $+\eta_{2}\max_{j}\Big\{-\log P_{\psi}\big(\sigma_{1}^{j}\succ\sigma_{0}^{j}\big)\Big\}$
-$\leq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega.$
-Step 2: Proving low bound : ^$\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$.
-^$\widehat{\mathcal{L}}_{u}(\psi)$
+$\leq$ $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega.$
+Step 2: Proving low bound : $\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$.
+$\widehat{\mathcal{L}}_{u}(\psi)$
 $=$ $\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$
@@
 $+\mathbbm{1}\big[y^{j}=1\big]\log P_{\psi}\big(\sigma_{1}^{j}\succ\sigma_{0}^{j}\big)$
-$-\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
-$-\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
-$\geq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega.$
-Combing Step 1 and Step 2, we can obtain the following result:
-|^$\Big{|}\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\Big{|}\leq\eta\Omega.$ (A.2)
+[…]
+[…]
+$\geq$ $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega.$
+Combining Step 1 and Step 2, we can obtain the following result:
+$\Big|\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\Big|\leq\eta\Omega.$ (A.2)
 This completes the proof of Theorem 1.
@@
 where $\mathbb{E}_{\widehat{\mathcal{D}}}\left[f\right]=\sum_{i=1}^{m}\left[g(x_{i})\right]/m$ is the empirical form of $\mathbb{E}_{\mathcal{D}}\left[f\right]$.
-This proof can be found in Theorem 3.3 of work [24].
+This proof can be found in Theorem 3.3 of the work [24].
 Proposition 2.
@@
 Then, for any sample $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the following relation holds between the empirical Rademacher complexities of $\Pi(\mathcal{R})$ and $\mathcal{F}$:
-^$\widehat{\Re}_{\widehat{\mathcal{D}}}(\mathcal{F})\leq 2\widehat{\Re}_{\widehat{\mathcal{D}}}\big{(}\Pi(\mathcal{R})\big{)}.$ (A.4)
+$\widehat{\Re}_{\widehat{\mathcal{D}}}(\mathcal{F})\leq 2\widehat{\Re}_{\widehat{\mathcal{D}}}\big(\Pi(\mathcal{R})\big).$ (A.4)
 Proof. According to Definition 1, the empirical Rademacher complexity of $\mathcal{F}$ can be written as:
-^$\widehat{\Re}_{\widehat{\mathcal{D}}}(\mathcal{F})$ (A.5)
+[…]
 $=$ $\mathbb{E}_{\bm{\sigma}}\bigg[\sup_{R\in\mathcal{R}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\big[-(1-y)\log P\left(\sigma_{0}\succ\sigma_{1}\right)$
@@
 Combing Eqs.
-(A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error ^$\widehat{\mathcal{L}}(\psi)$ holds:
+(A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error $\widehat{\mathcal{L}}(\psi)$ holds:
 $\mathcal{L}(\psi)\leq\widehat{\mathcal{L}}(\psi)+4\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}\big(\Pi(\mathcal{R})\big)+3\sqrt{\frac{\log(2/\delta)}{2N}}.$ (A.6)
-However, the dataset for training reward model includes two parts: the labeled dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$ and unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$, and empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)=\widehat{\mathcal{L}}_{l}(\psi)+\widehat{\mathcal{L}}_{u}(\psi)$.
-Let ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ be the empirical error under unlabeled dataset with true label, then
+However, the dataset for training reward model includes two parts: the labeled dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$ and unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$, and empirical error $\widehat{\mathcal{L}}_{R}(\psi)=\widehat{\mathcal{L}}_{l}(\psi)+\widehat{\mathcal{L}}_{u}(\psi)$.
+Let $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ be the empirical error under an unlabeled dataset with true label, then
 $\mathcal{L}_{R}(\psi)\leq\widehat{\mathcal{L}}_{l}(\psi)+\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+4\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}\big(\Pi(\mathcal{R})\big)+3\sqrt{\frac{\log(2/\delta)}{2(N_{l}+N_{u})}},$
 where $\widehat{\mathcal{D}}$ is the input combination of labeled and unlabeled dataset, denoted as $\{(\sigma_{0},\sigma_{1})^{(i)}\}_{i=1}^{N_{l}+N_{u}}$.
-According to Proposition 1, that is ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\leq\widehat{\mathcal{L}}_{u}(\psi)+\eta\Omega.$, we can derive
+According to Proposition 1, that is $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\leq\widehat{\mathcal{L}}_{u}(\psi)+\eta\Omega.$, we can derive
 $\mathcal{L}_{R}(\psi)\leq\widehat{\mathcal{L}}_{R}(\psi)+\eta\Omega+4\widehat{\mathfrak{R}}_{{\widehat{\mathcal{D}}}}\big(\Pi(\mathcal{R})\big)+3\sqrt{\frac{\log(2/\delta)}{2(N_{l}+N_{u})}}.$
@@
 A-C Proof of Theorem 2
-Before proof, we give the related Assumption and Lemmas for Theorem 2. Firstly, we assume the the reward class $\mathcal{R}$ is realizable (Assumption 1) and bounded (Assumption 3).
-Assumption 3 (Boundedness).
-For any $R\in\mathcal{R}$ and any state-action pairs $(s,a)$, the equation $|R(s,a)|\leq R_{max}$ holds.
 Proposition 3.
@@
 The proof of Proposition 3 can be found in Appendix A-D.
-Note that, the above reward model is only trained from offline preference dataset.
+Note that the above reward model is only trained from the offline preference dataset.
 Through generating more preference data, the upper bound can be tighter.
-However, the generated data may be inaccuracy, which may improve the reward gap.
-The selecting mechanism can effectively solve this problem.
-The Proof for Theorem 2: This Theorem aims to give the lower bound of term $J(\widehat{\pi},R^{*})-J(\mu,R^{*})$. We prove this from ideal case and actual case.
-1) Ideal case: There is no need to consider distribution shift and empirical error problem.
-The policy $\widehat{\pi}$ is directly learned by $\widehat{\pi}=\max_{\pi}J(\pi,\widehat{R})$.
-Thus, $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})$ holds.
-Then, we can derive
+However, the generated data may be inaccurate, which may improve the reward gap.
+The selection mechanism can effectively solve this problem.
+The Proof for Theorem 2: This Theorem aims to give the lower bound of term $J(\widehat{\pi},R^{*})-J(\mu,R^{*})$. We prove this from the ideal case and the actual case.
+[…]
 $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$ (A.8)
@@
 $\leq$ $J(\mu,R^{*})-J(\mu,\widehat{R})+J(\widehat{\pi},\widehat{R})-J(\widehat{\pi},R^{*})$
-$\leq$ $\big{|}J(\mu,R^{*})-J(\mu,\widehat{R})\big{|}+\big{|}J(\widehat{\pi},R^{*})-J(\widehat{\pi},\widehat{R})\big{|}.$
-Since $J(\pi,R):=\mathbb{E}_{(s,a)\sim d_{T}^{\pi}(s,a)}[R(s,a)]/(1-\gamma)$, $|J(\pi,R^{*})-J(\pi,\widehat{R})|$ can be written as
+[…]
+[…]
 $\big|J(\pi,R^{*})-J(\pi,\widehat{R})\big|=\frac{1}{1-\gamma}\Big|\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]\Big|.$
 Then, according to Definition 3, we have
-$\Big{|}J(\pi,R^{*})-J(\pi,\widehat{R})\Big{|}\leq\frac{\mathscr{C}_{\mathcal{R}}(\pi)}{1-\gamma}\Big{|}\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}\Big{|}.$
+$\Big|J(\pi,R^{*})-J(\pi,\widehat{R})\Big|=\frac{\mathscr{C}_{\mathcal{R}}(\pi)}{1-\gamma}\Big|\mathbb{E}_{(s,a)\sim d_{T}^{\mu 0}}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]\Big|.$
 By Proposition 3, the following inequality holds
@@
 $\bigg(\sqrt{\frac{4C}{NL^{2}}\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}+{\color[rgb]{0,0,0}\sqrt{\frac{4R_{max}^{2}\log\big(1/\delta\big)}{NL}}}\bigg).$
-2) Actual case: Offline RL suffers from distribution shift problem and it is a necessity to consider empirical error.
-Therefore, many offline RL algorithms incorporate conservatism into policy to overcome distribution shift, that is learning policy through $\widehat{\pi}=\max_{\pi}J(\pi,\widehat{R})-P(\pi)$.
-Therefore, $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})-\xi$ instead of $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})$ in actual implantation.
-The term $\xi$ depends on the algorithm itself.
-Here, we take CQL as a example.
-The policy gap is proven in Theorem 3.6 of [4].
-The detailed is given as follow.
+[…]
 $\xi=$ $\underbrace{\frac{\gamma C_{T,R}R_{max}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d_{T}^{\mu}(s)}\sqrt{\frac{|\mathcal{A}|(1+D(s))}{|\mathcal{D}(s)|}}}_{:=\xi_{1}}$ (A.10)
 $-\underbrace{\frac{\alpha}{1-\gamma}\mathbb{E}_{s\sim d_{T}^{\mu}(s)}\bigg[D(s)\bigg]}_{:=\xi_{2}},$
-where ||$\cdot$|| ⋅ | denotes cardinality of a specific set, $D(s)=\sum_{a}\widehat{\pi}(a|s)(\frac{\widehat{\pi}(a|s)}{\mu(a|s)}-1)$ and $C_{T,R}$ is the empirical coefficient.
+where $|\cdot|$ denotes cardinality of a specific set, $D(s)=\sum_{a}\widehat{\pi}(a|s)(\frac{\widehat{\pi}(a|s)}{\mu(a|s)}-1)$ and $C_{T,R}$ is the empirical coefficient.
 It consists of two terms: the first term $\xi_{1}$ captures the decrease in policy performance due to sampling error.
-The second term $\xi_{2}$ captures the increase in policy performance in empirical setting [4].
+The second term $\xi_{2}$ captures the increase in policy performance in the empirical setting [4].
 Therefore, when considering distribution shift and empirical error, we can derive
@@
 $\leq$ $J(\mu,R^{*})-J(\mu,\widehat{R})+J(\widehat{\pi},\widehat{R})+\xi-J(\widehat{\pi},R^{*})$
-$\leq$ $\xi+\big{|}J(\mu,R^{*})-J(\mu,\widehat{R})\big{|}+\big{|}J(\widehat{\pi},R^{*})-J(\widehat{\pi},\widehat{R})\big{|}.$
+[…]
 Furthermore, according to Eq. (A.9), we have
@@
 $\mathcal{N}_{\mathcal{R}}(1/N)$ is the bracketing number defined in Definition 3.
-This proof can be found in Lemma 2 of previous work [18] and Proposition 14 of [23].
+This proof can be found in Lemma 2 of the previous work [18] and Proposition 14 of [23].
 The Proof for Proposition 3: We prove this Proposition from two steps.
@@
 $\geq$ $\frac{1}{4}\bigg|\sum_{l=1}^{L}\sum_{j=0}^{1}\Big(R^{*}(s_{j}^{l},a_{j}^{l})-\widehat{R}(s_{j}^{l},a_{j}^{l})\Big)\bigg|.$
-Then, combing Eqs. (A.12) and (A.14), we can further derive
+Then, combining Eqs. (A.12) and (A.14), we can further derive
 $\bigg|\sum_{i=1}^{N}\sum_{l=1}^{L}\sum_{j=0}^{1}\Big(R^{*}(s_{j}^{i,l},a_{j}^{i,l})-\widehat{R}(s_{j}^{i,l},a_{j}^{i,l})\Big)\bigg|$ (A.15)
@@
 (A.15), we set $X_{i}$ as $R^{*}(s_{j},a_{j})-R(s_{j},a_{j})$ and $X_{i}\in[-2R_{max},2R_{max}])$.
-Then, the below equation holds
+Then, the following equation holds.
 $\bigg|\mathbb{E}_{(s,a)\sim d_{T}^{\mu}(s,a)}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]\bigg|$ (A.17)
@@
 This completes proof of Proposition 3.
+Figure A.1: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) pen-expert and (b) hammer-expert datasets.
+TABLE A.1: Comparison results between URLHF using fewer preference data and LEASE, where the offline RL algorithm is CQL and URLHF using fewer data is denoted as $\texttt{URLHF}^{*}$.
+Task Name $\textbf{URLHF}^{*}$ URLHF FEWER LEASE
+walker2d-m $76.1\pm 0.8$ $6.0\pm 0.9 77$$7.4\pm 0.6$ $\mathbf{78.4\pm 0.9}$
+walker2d-m-e $86.8\pm 18.6$ $92.8\pm 22.4$ $77.7\pm 0.3$ $\mathbf{98.6\pm 18.1}$
+hopper-m $\mathbf{56.6\pm 2.4}$ $4.7\pm 3.4 55$$5.8\pm 2.8$ $\mathbf{56.5\pm 0.6}$
+hopper-m-e $55.3\pm 0.9$ $\mathbf{57.4\pm 4.9}$ $53.6\pm 0.9$ $56.4\pm 0.8$
+halfcheetah-m $43.3\pm 0.2$ $\mathbf{43.4\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.5\pm 0.1}$
+halfcheetah-m-e $58.9\pm 2.3$ $\mathbf{62.7\pm 7.1}$ $48.3\pm 0.7$ $53.2\pm 3.1$
+Mujoco Average $62.8\pm 4.2$ $\mathbf{64.5\pm 6.5}$ $59.4\pm 0.9$ $\mathbf{64.4\pm 4.0}$
+pen-human $\mathbf{17.7\pm 13.0}$ $9.8\pm 14.1$ $0.5\pm 3.0$ $3.8\pm 4.6$
+pen-expert $114.6\pm 53.7$ $\mathbf{138.3\pm 5.2}$ $128.1\pm 0.7$ $130.6\pm 1.8$
+door-human $1.7\pm 1.1$ $\mathbf{4.7\pm 5.9}$ $0.2\pm 1.0$ $\mathbf{4.7\pm 8.8}$
+door-expert $\mathbf{103.3\pm 0.5}$ $\mathbf{103.9\pm 0.8}$ $103.0\pm 0.9$ $\mathbf{103.2\pm 0.7}$
+hammer-human $0.7\pm 0.1$ $\mathbf{0.9\pm 0.3}$ $0.3\pm 0.0$ $0.3\pm 0.0$
+hammer-expert $117.4\pm 2.7$ $120.2\pm 6.8$ $124.1\pm 2.1$ $\mathbf{126.3\pm 1.2}$
+Adroit Average $59.2\pm 11.8$ $\mathbf{63.0\pm 5.5}$ $59.4\pm 1.3$ $61.5\pm 2.9$
+A-E Other Experimental Results
+These part presents supplementary experimental results, including: 1) the performance evaluation of the reward model on other tasks, 2) a comparative analysis between the proposed method LEASE and URLHF when trained on identical datasets, 3) an investigation of our framework’s effectiveness in model-based offline RL settings, and 4) further analysis on the performance on benchmark tasks.
+The comprehensive results provide additional empirical validation of the performance of the proposed method.
+A-E1 The other comparison results for reward model performance
+Fig.
+A.1 shows the comparison between the prediction value by the learned rewards and their ground truths for LEASE, FEWER, and FRESH under Adroit tasks, where the offline algorithm is IQL.
+The predicted and true rewards are both normalized to $[0,1]$.
+We randomly sample $500$ data from unlabeled datasets that are not seen in the training stage for evaluation, like Fig.
+4.
+Since the reward has a value in certain positions in Adroit tasks, the figure is in the form of multiple scattered lines in Fig.
+A.1.
+As evidenced by the experimental results shown in this figure, the reward model trained under the LEASE framework achieves the smallest variance in predicted reward values, which shows that the data augmentation and selection mechanism can improve the reward model’s performance.
+A-E2 The further comparison results between LEASE and URLHF
+Considering that it is crucial to include a comparison with URLHF using the same amount of data as LEASE.
+To further show the superior performance of the proposed method, we compare LEASE to the baseline algorithm URLHF with the same amount of data as LEASE, where the latter method is denoted as $\text{URLHF}^{*}$.
+Table A.1 shows that the average performance of LEASE is superior to that of the baseline algorithm URLHF using the same amount of data with LEASE.
+The experimental results demonstrate that the proposed algorithm LEASE achieves a performance improvement of approximately 3.9% compared to $\text{URLHF}^{*}$.
+This further validates the effectiveness and advantages of the proposed framework and method.
+TABLE A.2: Comparison results of offline RL algorithms under the designed framework.
+Task Name $\textbf{CQL}^{*}$ $\textbf{IQL}^{*}$ $\textbf{COMBO}^{*}$
+walker2d-m $78.4\pm 0.9$ $73.0\pm 2.1$ $71.6\pm 2.4$
+walker2d-m-e $98.6\pm 18.1$ $108.1\pm 0.5$ $79.1\pm 1.1$
+hopper-m $56.5\pm 0.6$ $56.0\pm 0.5$ $54.8\pm 0.9$
+hopper-m-e $6.4\pm 0.8 55$$5.9\pm 1.9$ $54.9\pm 1.1$
+halfcheetah-m $43.5\pm 0.1$ $43.0\pm 0.3$ $42.9\pm 0.1$
+halfcheetah-m-e $53.2\pm 3.1$ $60.2\pm 1.2$ $73.8\pm 7.0$
+Mujoco Average $4.4\pm 4.0 66$$6.4\pm 1.1$ $62.9\pm 2.1$
+A-E3 The results for LEASE under model-based RL offline algorithm
+To enhance the generalizability of the proposed framework, we extended our experimental validation beyond model-free offline RL approaches (such as CQL and IQL) to include model-based offline RL methodologies.
+Table A.2 shows the results of COMBO [49] under the designed framework, where * denotes the results of offline RL algorithms under our designed framework.
+For COMBO hyperparameters, the probability confidence $\kappa_{p}$, and uncertainty variance $\kappa_{\tau}$ are set as $0.85$ and $0.05$ for all tasks, respectively.
+Note that in our framework, model-based methods do not necessarily perform better than model-free methods.
+Model-based RL methods focus on how to learn a conservative policy by regularizing $Q$ values or penalizing rewards to alleviate the effects of inaccurate model data.
+These methods critically depend on the accuracy of the reward model.
+Consequently, model-based approaches are more sensitive to reward modeling errors than model-free methods, which can hinder the improvement of agent performance.
+Therefore, model-based offline RL requires a higher accuracy of the reward model than model-free RL.
+A-E4 Further analysis for the results on benchmark tasks
+Table A.1 shows that LEASE doesn’t achieve better results on all tasks.
+Compared to URLHF*, LEASE shows significant improvements on several tasks, such as walker2d-m-e, pen-expert, and hammer-expert.
+For tasks like walker2d-m, hopper-m-e, halfcheetah-m, and door-human, the performance shows slight improvements.
+It is widely accepted that the current improved RL algorithms struggle to perform well across all tasks.
+The network architectures and parameters of URLHF and LEASE differ, which could also lead to the performance discrepancy.
+It can also be observed that in the hopper, halfcheetah, and door environments, the performance improvement of LEASE over URLHF* is minimal, and in some cases, even lower.
+This is because the accuracy of the reward function in these environments does not have a significant impact on performance.
+Table A.3 has validated this conclusion.
+This table shows the performance comparison of the agent in the above three environments using official URLHF code, with different numbers of preference data $100$, $500$, and $1000$.
+TABLE A.3: Comparison results between different numbers of preference data.
+Number hopper-medium halfcheetah-medium door-expert
+$100$ $56.6\pm 2.4$ $43.3\pm 0.2$ $103.3\pm 0.5$
+$500$ $55.8\pm 0.1$ $43.4\pm 0.2$ $103.8\pm 1.0$
+$1000$ $55.6\pm 2.8$ $43.3\pm 0.2$ $103.6\pm 0.4$

### v3

@@
 Reinforcement learning (RL) has been widely applied in robot control [1, 10] and healthcare [12, 11] fields.
-However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [39]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9].
+However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [39, 50]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9].
 Much research has been conducted to solve the above two challenges.
@@
 II Related Works
-This section presents a brief overview of the algorithms and related theories of offline PbRL, and discusses the differences and advantages between our work and the previous related works.
+This section provides a brief overview of the algorithms and related theories in offline PbRL, and discusses the differences and advantages between our work and previous associated works.
 The algorithm for offline PbRL.
-Offline PbRL eliminates the demand for interaction with the environment and handcrafted-designed reward.
+Offline PbRL eliminates the demand for interaction with the environment and handcrafted-designed rewards.
 OPAL [34] is the first algorithm that combines offline RL and PbRL.
@@
 Instead, our method LEASE trains a simple transition model to generate unlabeled data and designs a selection mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer datasets and less time.
+The aforementioned approach first learns an explicit reward function from preference data and then utilizes it to guide policy learning.
+However, in complex tasks, the reward function may contain substantial errors, which can degrade the performance of policy learning.
+Recently, preference-based direct policy learning has gained significant research attention [51, 52, 53, 54], as it directly optimizes the policy based on user preferences.
+The most representative work in this area is the Direct Preference Optimization (DPO) algorithm [51], which formulates an initial preference optimization objective.
+Subsequent research has focused on designing different forms of preference optimization functions to further improve policy learning performance, including simpler preference optimization formulations [52, 53] and optimization objectives based on prospect theory [54].
+This direction constitutes a key focus of preference-based policy optimization.
 The theory for offline PbRL.
@@
 Given a pair of segments $(\sigma_{0},\sigma_{1})$, human choose which segment is preferred, i.e., $y\in\{0,1\}$ [14].
-The preference label $y=1$ indicates $\sigma_{1}\succ\sigma_{0}$ and $y=0$ indicates $\sigma_{0}\succ\sigma_{1}$, where $\sigma_{i}\succ\sigma_{j}$ denotes that the segment $i$ is preferable to the segment $j$.
+The preference label $y=1$ indicates $\sigma_{1}\succ\sigma_{0}$, while $y=0$ indicates $\sigma_{0}\succ\sigma_{1}$, where $\sigma_{i}\succ\sigma_{j}$ denotes that segment $i$ is preferred over segment $j$.
 The preference data is stored in the dataset $\mathcal{D}_{l}$ as a tuple $(\sigma_{0},\sigma_{1},y)$.
@@
 $+$ $(1-y)\log P_{\psi}\left(\sigma_{0}\succ\sigma_{1}\right)\big].$
+Conversely, given a learned reward model, the preference label between a pair of trajectories ($\sigma_{0},\sigma_{1}$) is determined as follows: if the probability computed by Eq.
+(2) exceeds 0.5, trajectory 0 ($\sigma_{0}$) is inferred to be preferred over trajectory 1 ($\sigma_{1}$); otherwise, preference is assigned to $\sigma_{1}$.
 After obtaining the reward function $\widehat{R}$, offline PbRL optimizes the policy by maximizing the expected discounted return $J(\pi,R)$ like standard offline RL.
@@
 $\mathcal{L}_{T}(\varphi)=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\text{offline }}}\big[\log\widehat{T}_{\varphi}(s^{\prime}~|~s,a)\big].$ (4)
-In practical implantation, the transition model is approximated by Gaussian distribution and the MLE method is employed to train $N_{T}$ ensemble transition models $\{T^{i}_{\varphi}=\mathcal{N}(\mu^{i}_{\varphi},\sigma^{i}_{\varphi})\}_{i=1}^{N_{T}}$ [37].
+In practical implementation, to account for the inherent uncertainty in the next-state prediction, a Gaussian distribution is employed to approximate the transition model.
+To further mitigate the uncertainty associated with model predictions, an ensemble of $N_{T}$ transition models $\{T^{i}_{\varphi}=\mathcal{N}(\mu^{i}_{\varphi},\sigma^{i}_{\varphi})\}_{i=1}^{N_{T}}$ is trained.
+These models are then optimized using a maximum likelihood function [37].
 The samples are generated through $H$-step rollouts.
@@
 This assumption states that there must exist a reward function $R\in\mathcal{R}$ which can well-approximate the optimal reward function when the preference data can cover all data space.
-However, preference data tend to be limited, even fewer than a hundred in real-world scenarios.
-Therefore, different from previous PbRL [3, 35, 33], this paper aims to study how to learn the better reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$.
+However, preference data tend to be limited, with even fewer than a hundred in real-world scenarios.
+Therefore, different from previous PbRL [3, 35, 33], this paper aims to study how to learn a better reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$.
 To broaden the coverage of preference data, motivated by model-based offline RL methods, we train the transition model through Eq.
@@
 Then, the pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
-$\widehat{y}\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)=\mathbbm{1}\big[P\left(\sigma_{1}^{u}\succ\sigma_{0}^{u};{\psi}\right)>0.5\big],$ (5)
+$\widehat{y}\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)=\mathbbm{1}\big[P\left(\sigma_{0}^{u}\succ\sigma_{1}^{u};{\psi}\right)<0.5\big],$ (5)
 where $\mathbbm{1}(\cdot)$ is indicator function.
+When the probability is less than 0.5, it indicates that $\sigma_{1}\succ\sigma_{0}$ is highly likely to hold, and the preference label is accordingly set to 1; otherwise, it is set to 0.
 The reward function can be updated through the collected labeled dataset $\mathcal{D}_{l}$ and the generated unlabeled dataset $\mathcal{D}_{u}$.
@@
 The reward learning involves two stages: pretraining based on collected labeled data $\mathcal{D}_{l}$ and updating based on $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$ during policy learning, which is essentially semi-supervised learning [26].
-This section focuses on designing a data selection function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of generated preference data, and explaining the factors that influence the generalization ability of the reward model.
+However, the generated preference labels may contain errors, which can degrade the performance of the reward model.
+Therefore, this section focuses on designing a data selection function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of the generated preference data and explains the factors influencing the generalization capability of the reward model.
 IV-A Uncertainty-aware Pseudo-label Selection
@@
 where $\kappa_{p}$ and $\kappa_{\tau}$ are thresholds of confidence and uncertainty respectively.
-$f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes the generated data $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected, and $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ denotes the data is not selected.
+Here, $f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes that the generated data pair $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected for inclusion, whereas $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ indicates its exclusion.
 Intuitively, more rigorous data selection criteria, characterized by higher confidence and lower variance, lead to a reduction in the number of qualified samples.
 While such a strategy improves the quality of the generated data, it could lead to a narrower data distribution, which might adversely affect the precision of the trained reward model.
-Therefore, appropriate selection of parameters $\kappa_{p}$ and $\kappa_{\tau}$ is required to enhance the performance of the reward model.
+Therefore, an appropriate selection of parameters $\kappa_{p}$ and $\kappa_{\tau}$ is required to enhance the performance of the reward model.
+The unlabeled dataset selected through $f(\sigma_{0}^{u},\sigma_{1}^{u})$ is denoted as $\widetilde{\mathcal{D}}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{\widetilde{N}_{u}}$, where $\widetilde{N}_{u}$ is the number of generated dataset after screening.
 Then, combining Eq.
@@
 (5), then the pseudo-labeling error for the unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(j)}\}_{j=1}^{N_{u}}$ is smaller than $\eta$, i.e., $\sum_{j=1}^{N_{u}}\mathbbm{1}[\widehat{y}^{j}\neq y^{j}]/N_{u}\leq\eta$.
-Note that the bound of pseudo-label error rate $\eta$ of the generated data filtered by the proposed selection mechanism is significantly lower than that of the unfiltered data, as the mechanism effectively eliminates low-confidence and high-uncertainty samples.
-This observation is experimentally validated in Table II and Fig.
+Thanks to its capacity to effectively filter out samples with low confidence and high uncertainty, the screening mechanism mitigates the adverse impact of negative samples on reward model training, thereby contributing to improved reward model performance.
+Consequently, the bound on the pseudo-label error rate $\eta$ for the filtered generated data is significantly lower than that for the unfiltered data.
+This observation is experimentally validated in Table III and Fig.
 3.
@@
 The proof of Theorem 1 can be found in Appendix B.
-Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by empirical error $\widehat{\mathcal{L}}_{R}(\psi)$, pseudo-labeling error $\eta\Omega$, Rademacher complexity and constant terms.
-The constant term can become small when the number of unlabeled data increases, but it may cause unstable training when labels are inaccurate.
-According to Proposition 1, empirical error and pseudo-labeling error can be reduced through the selection mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, thus, a better generalization ability, that is tighter upper bound, can be achieved.
+Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by the empirical error $\widehat{\mathcal{L}}_{R}(\psi)$, the pseudo-label error $\eta\Omega$, the Rademacher complexity, and a constant term.
+While the constant term diminishes as the amount of unlabeled data increases, inaccurate labels can lead to training instability.
+According to Proposition 1, a selection mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$ can be employed to reduce both the empirical error and the pseudo-label error, thereby achieving better generalization performance, that is, a tighter upper bound.
+Conversely, without such a screening mechanism, the pseudo-label error rate would rise, resulting in a larger upper bound for the expected error of the reward model and ultimately degrading its performance.
 Moreover, Theorem 1 has generality and can apply to methods that train a reward model using pseudo-labels.
@@
 This section aims to describe the details of LEASE and establish the theory for a safe policy improvement guarantee.
-In this paper, we directly use the proposed offline RL algorithms, such as CQL [4] and IQL [25], to perform policy learning, and do not improve the offline RL algorithm itself.
+In this paper, we directly use the proposed offline RL algorithms, such as Conservative Q-Learning (CQL) [4] and Implicit Q-Learning (IQL) [25], to perform policy learning, and do not improve the offline RL algorithm itself.
 V-A The Implementation Details of LEASE
@@
 The proposed algorithmic framework comprises three key components: reward model training, transition model training, and policy training.
-Initially, the transition model is trained using limited offline datasets, while the reward model undergoes pretraining with a small-scale preference dataset.
-The reward update condition is triggered when the number of unlabeled data in the buffer reaches its maximum capacity.
-Subsequently, the policy training phase commences.
-During this phase, the transition model is periodically employed to generate unlabeled synthetic data, and the reward model undergoes iterative updates to enhance its performance.
+Initially, the transition model is trained on a limited offline dataset, whereas the reward model is pretrained using a small-scale preference dataset.
+The reward update mechanism is activated once the volume of unlabeled data in the buffer reaches its upper capacity limit.
+Subsequently, the policy training phase is initiated.
+In this phase, the transition model is periodically leveraged to generate unlabeled synthetic data, and the reward model undergoes iterative updates to further boost its performance.
 V-B Safe Policy Improvement Guarantee
@@
 For offline PbRL, we further investigate whether similar policy improvement guarantees can be established.
+Unlike offline RL, the reward function in PbRL is unknown, and the learned reward function may contain errors, which introduces an additional performance gap.
 Therefore, this part aims to develop a theoretical guarantee of policy improvement for LEASE, that is, giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$.
@@
 The proof of Theorem 2 can be found in Appendix C.
-The bounds of the PbRL policy learned by LEASE comprise two distinct components: the first term $\xi$ stems from the inherent gap of the offline RL itself, and the second arises from the discrepancy induced by the reward model.
-The first term $\xi$ is constant when the offline algorithm is determined.
-The second term is determined by the performance of the reward model.
-Therefore, the tighter bound can be achieved through reducing the performance gap caused by the reward model (see Proposition 3).
+The performance bound for the PbRL policy learned by LEASE consists of two distinct components: the first term, denoted as $\xi$, stems from the inherent gap of the offline RL process itself; the second term arises from the discrepancy induced by the reward model.
+When the offline algorithm is fixed, the first term $\xi$ is constant.
+The second term depends on the performance of the learned reward model.
+Consequently, the proposed framework is readily compatible with other offline RL algorithms, and a tighter overall bound can be achieved by reducing the performance gap attributable to the reward model (see Proposition 3).
 It can be reduced to a small value $\varpi$ with a sample complexity of
@@
   • $\mathbf{Q_{2}}$: How well does LEASE perform compared with the offline PbRL baseline in standard benchmark tasks?
-  • $\mathbf{Q_{3}}$: How does the proposed selection mechanism and parameter settings affect the performance of reward and policy?
+  • $\mathbf{Q_{3}}$: How do the proposed selection mechanism and parameter settings affect the performance of reward and policy?
 We answer these questions on two distinct domains (mujoco and adroit), covering several locomotion and manipulation tasks, within the D4RL benchmark [22]. The code for LEASE is available at Code.
@@
 The offline datasets are based on the D4RL dataset [22].
-We select medium and medium-expert two types of datasets for Mujoco tasks and human and expert for Adroit tasks.
+We select two types of datasets for Mujoco tasks: medium and medium-expert, and human and expert for Adroit tasks.
 The difference between datasets in a certain task lies in the policies they collect.
@@
 2 depicts the six tasks for the D4RL benchmark.
-For three mujoco tasks, the goal is to control the robot’s various joints to achieve faster and more stable locomotion.
-For adroit tasks, they involve controlling a 24-DoF simulated Shadow Hand robot on a sparse reward, high-dimensional robotic manipulation task.
+The objective of the three Mujoco tasks is to control the robot’s joints to achieve faster and more stable motion, with action dimensions ranging from 3 to 8, and the reward function is defined in the form of continuous values.
+In contrast, the Adroit task involves controlling a simulated Shadow Hand robot with an action dimension of 24, which is a high-dimensional robotic manipulation task characterized by sparse rewards—only a few specific reward values are provided.
 The preference dataset is from the previous work [21], where they provided the baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF).
-One is crowd-sourced labels obtained by crowd-sourcing, and the other is synthetic labels based on ground truth reward.
+One is crowd-sourced labels obtained through crowd-sourcing, and the other is synthetic labels based on ground-truth rewards.
 To test LEASE performance for two types of preference data, the labels of preference are from human-realistic feedback for MuJoco tasks, and ground truth reward for Adroit tasks.
@@
 The details are given below:
+Figure 3: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) hopper-medium-expert and (b) halfcheetah-medium-expert datasets, where the value predicted by the trained reward model and ground-truth reward value are both normalized to $[0,1]$.
+From left to right are methods LEASE, FEWER and FRESH, respectively.
 TABLE I: Base parameter configuration of method LEASE.
@@
 Batch size 256
-Figure 3: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) hopper-medium-expert and (b) halfcheetah-medium-expert datasets, where the value predicted by the trained reward model and ground-truth reward value are both normalized to $[0,1]$.
-From left to right are methods LEASE, FEWER and FRESH, respectively.
 Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models.
@@
 Each reward model includes a $3$-layer feed-forward neural network with $256$ hidden units.
-The reward model training uses cross-entropy loss with a learning rate of $3\times 10^{-4}$ and Adam optimizer.
+The training of the reward model uses cross-entropy loss with a learning rate of $3\times 10^{-4}$ and the Adam optimizer.
 Policy optimization: The offline algorithms include CQL [4] and IQL [25].
@@
 […]
-TABLE II: Comparison results of pseudo-labeling accuracy between using the selection mechanism (Ours), only using confidence (No-U), only using uncertainty (No-C), and not using the selection mechanism (No-C-U).
-Task Name Ours No-U No-C No-C-U
-pen-expert $\mathbf{87.3\pm 0.5}$ $85.9\pm 0.8$ $86.7\pm 0.6$ $86.8\pm 0.5$
-door-expert $\mathbf{89.3\pm 0.6}$ $87.8\pm 0.2$ $86.0\pm 0.0$ $85.9\pm 0.8$
-hammer-expert $\mathbf{85.5\pm 0.2}$ $84.4\pm 0.4$ $82.5\pm 1.2$ $83.6\pm 0.5$
-Average $\mathbf{87.4\pm 0.4}$ $86.0\pm 0.5$} $85.1\pm 0.6$} $85.4\pm 0.6$}
-TABLE III: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds.
+Figure 4: The relationship between pseudo-labeling accuracy and the number of preference dataset $N_{l}$.
+TABLE II: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds.
 $\pm$ captures the standard deviation over seeds.
@@
 VI-B Results for Reward Model
-VI-B1 The effect of the selection mechanism for pseudo-label accuracy
-We employ a pre-trained reward model to annotate unlabeled preference data generated during the process.
-This paper proposes a novel screening mechanism to select data samples with both high confidence and low uncertainty, thereby enhancing the accuracy of pseudo-labels.
-The pseudo-label is generated by a preference model, which is constructed based on the reward model as formulated in Eq.
-(2).
-Table II compares the pseudo-label accuracy under four different conditions: 1) using the proposed screening mechanism (combining both confidence and uncertainty thresholds), 2) employing only the confidence threshold (No-U), 3) applying only the uncertainty threshold (No-C), and 4) without any screening mechanism (No-C-U).
-The accuracy of the pseudo-label generated by the reward model is evaluated on all preference datasets.
-Figure 4: The relationship between pseudo-labeling accuracy and the number of preference dataset $N_{l}$.
-Table II shows that the proposed screening mechanism significantly improves pseudo-label accuracy, i.e., reduces error $\eta$ defined in Assumption 2.
-The pseudo-label accuracy is improved by $4.5\%$, $2.8\%$, and $6.9\%$ for pen-expert, door-expert, and hammer-expert, respectively.
-Moreover, we also analyze the effect of the number of preference data on the preference model.
-Fig.
-4 illustrates the relationship between the prediction accuracy of the preference model without data augmentation and the number of preference data $N_{l}$.
-Here, we use the last $200$ of $2000$ preference data that are not seen in the training stage as the evaluation dataset.
-It indicates that as the number of preference data $N_{l}$ increases, the prediction accuracy of the preference model shows an upward trend.
-This also validates Theorem 1 that increasing the number of samples $N_{l}$ can reduce the generalization error.
-VI-B2 The effect of data augmentation on the performance of the reward model
+VI-B1 The effect of data augmentation on the performance of the reward model
 To answer question $\mathbf{Q_{1}}$, we compare the accuracy of the reward model before and after data augmentation under a smaller preference dataset.
@@
 This figure indicates that the linear relationship between reward predicted by LEASE and ground truth is better than that of the other two methods, where the method not using a selecting mechanism is denoted as FRESH.
-The prediction of the reward model learned by method FRESH is very narrow, and the accuracy is greatly reduced compared with FEWER.
-This is because the generated data of FRESH are not selected by $f(\sigma_{0},\sigma_{1})$, which causes substantial errors for the labels of generated data and leads to the collapse of training.
-This validates that the performance of the reward model can be improved through data augmentation and a selection mechanism, where the designed selection mechanism has a greater impact on the reward model’s performance.
+The reward model learned by the method FRESH exhibits a notably narrow prediction range, and its accuracy significantly drops compared with that of FEWER.
+This is because the data generated by FRESH are not filtered by $f(\sigma_{0},\sigma_{1})$, leading to substantial errors in data labeling and consequently causing training collapse.
+These results validate that the performance of a reward model can be enhanced through data augmentation and selection mechanisms, with the design of the selection mechanism having a more pronounced impact on the model’s overall performance.
+TABLE III: Comparison results of pseudo-labeling accuracy between using the selection mechanism (Ours), only using confidence (No-U), only using uncertainty (No-C), and not using the selection mechanism (No-C-U).
+Task Name Ours No-U No-C No-C-U
+pen-expert $\mathbf{87.3\pm 0.5}$ $85.9\pm 0.8$ $86.7\pm 0.6$ $86.8\pm 0.5$
+door-expert $\mathbf{89.3\pm 0.6}$ $87.8\pm 0.2$ $86.0\pm 0.0$ $85.9\pm 0.8$
+hammer-expert $\mathbf{85.5\pm 0.2}$ $84.4\pm 0.4$ $82.5\pm 1.2$ $83.6\pm 0.5$
+Average $\mathbf{87.4\pm 0.4}$ $86.0\pm 0.5$} $85.1\pm 0.6$} $85.4\pm 0.6$}
+VI-B2 The effect of the selection mechanism for pseudo-label accuracy
+This paper introduces a novel screening mechanism designed to select unlabeled preference data samples characterized by both high confidence and low uncertainty, thereby improving pseudo-label accuracy.
+The pseudo-labels are generated by a preference model derived from a pre-trained reward model, as defined in Eq.
+(2).
+The performance of this approach is evaluated in terms of pseudo-label accuracy across all preference datasets.
+As shown in Table III, the proposed screening mechanism—which simultaneously applies confidence and uncertainty thresholds—is compared against three baseline conditions: employing only the confidence threshold (No-U), employing only the uncertainty threshold (No-C), and using no screening mechanism (No-C-U).
+Table III demonstrates that the proposed screening mechanism substantially enhances the pseudo-label accuracy, i.e., it reduces the error $\eta$ defined in Assumption 2.
+For the Pen, Door, and Hammer experts, the pseudo-label accuracy is improved by 4.5%, 2.8%, and 6.9%, respectively.
+Furthermore, we analyze the influence of the amount of preference data on the preference model.
+Fig.
+4 illustrates the relationship between the prediction accuracy of the preference model—without data augmentation—and the quantity of preference data $N_{l}$.
+Here, we use the final 200 preference samples (out of a total of 2000) that were not observed during training as the evaluation dataset.
+The results indicate that as the amount of preference data $N_{l}$ increases, the prediction accuracy of the preference model exhibits an upward trend.
+This observation also corroborates Theorem 1, which states that increasing the number of samples $N_{l}$ reduces the generalization error.
 VI-C Results for Policy Performance
@@
 To answer question $\mathbf{Q_{2}}$, here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms on six types of tasks and 12 datasets in total.
-Table III presents a comparative analysis of he results of the average normalized score with standard deviation under three distinct scenarios: 1) performance with extensive preference data (URLHF), 2) performance with limited data without employing data augmentation techniques (FEWER), and 3) performance with limited data while incorporating data augmentation (our proposed method LEASE).
+Table II presents a comparative analysis of he results of the average normalized score with standard deviation under three distinct scenarios: 1) performance with extensive preference data (URLHF), 2) performance with limited data without employing data augmentation techniques (FEWER), and 3) performance with limited data while incorporating data augmentation (our proposed method LEASE).
 In this table, m and m-e represent medium and medium-expert, respectively.
 The results of the offline PbRL algorithm under sufficient feedback (URLHF) are obtained from the paper [21].
-As demonstrated in Table III, LEASE exhibits significantly superior performance compared to FEWER.
+As demonstrated in Table II, LEASE exhibits significantly superior performance compared to FEWER.
 This empirical evidence substantiates that data augmentation can effectively enhance the agent’s capability.
 Furthermore, compared to the performance achieved with extensive preference data URLHF, LEASE attains comparable results in most cases.
-However, the proposed method exhibits slightly inferior performance on specific tasks, such as hopper-medium-expert and halfcheetah-medium-expert datasets.
-This limitation primarily stems from the excessively narrow distribution of the expert dataset, which consequently induces a constrained distribution in the generated preference data.
-As a result, the reward model’s performance cannot be effectively enhanced through data augmentation, thereby restricting the potential improvement in agent performance.
-However, overall, the results in Table III can validate the content of Theorem 2, which shows that LEASE can reduce the performance gap caused by the reward model and improve the agent’s performance with fewer preference data.
+However, the proposed method exhibits slightly inferior performance on certain specific tasks, such as the hopper-medium-expert and halfcheetah-medium-expert datasets.
+This limitation primarily stems from the overly narrow distribution of the expert datasets, which consequently constrains the distribution of the generated preference data.
+As a result, data augmentation fails to effectively enhance the performance of the reward model, thereby limiting the potential improvement of the agent’s performance.
+However, overall, the results in Table II can validate the content of Theorem 2, which shows that LEASE can reduce the performance gap caused by the reward model and improve the agent’s performance with fewer preference data.
 TABLE V: The comparison results for the D4RL tasks under different number of preference data $N_{l}$. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score.
@@
 This table indicates LEASE can still improve agent performance under a very small preference dataset.
-Notably, the performance gap between $100$ preference data and $20$ preference data is not much.
-This indicates that LEASE has the potential to perform well under a smaller preference dataset.
+It is noteworthy that the performance gap between using 100 preference datasets and 20 preference datasets is not substantial.
+This is primarily because the rewards in the D4RL dataset are relatively simple; a small amount of preference data is sufficient to learn a reasonably accurate reward model.
+Coupled with data augmentation, the performance of the reward function can be further improved, thereby enhancing the agent’s overall capability.
+This indicates that LEASE has the potential to perform well even with a limited number of preference datasets.
 VI-D Parameter Analysis
@@
 Actual application.
-In actual scenarios, especially for human-in-loop control, such as exoskeleton robot assistance or rehabilitation, designing a highly sample-efficient RL algorithm is greatly significant.
-Firstly, many rewards are difficult to describe in mathematical terms in some situations, such as the comfort of interaction.
-Human feedback or preference is a better way to reflect the above indices.
-In addition, in human-in-the-loop control, obtaining preference data through interaction can easily cause fatigue, and inappropriate interaction may cause damage to humans.
-It is a necessity to learn a reward model from a limited preference dataset.
-Then, for robot assistance or rehabilitation, the goal of the robot is to reduce human metabolic cost and guarantee the comfort of interaction.
-However, the reward model for this goal is difficult to design.
-Therefore, PbRL has a great potential to improve control performance in some scenarios where the reward function is difficult to describe and humans are in the control loop.
+The high-sample-efficiency preference-based offline RL algorithm proposed in this work aims to learn robot control policies from offline data with minimal human feedback, under the condition of an unknown reward function.
+This approach is particularly suitable for scenarios where the reward function is difficult to specify mathematically and where human-robot interaction entails high cost.
+For instance, in the field of exoskeleton robot assistance, objectives such as maximizing assistive comfort and minimizing metabolic cost are challenging to define precisely, and prolonged human-in-the-loop interaction can lead to user fatigue.
+Therefore, it is imperative to learn a reward model from a limited preference dataset and utilize it to efficiently guide policy optimization.
+The deployment of the algorithm for practical control involves three main stages: 1) collecting an offline dataset (containing only $(s,a,s^{\prime})$ tuples) through methods such as teleoperation, to train a transition model; 2) acquiring a limited set of preference data by having users select their preference between pairs of trajectories, which is then used to learn a reward function; and 3) training the robot control policy offline using the collected dataset and the learned reward model, followed by its deployment to the actual control system.
+The ultimate goal is to optimize a control policy offline with minimal human-robot interaction efficiently.
 Theoretical study.
@@
 Moreover, the theory of LEASE is based on state-action pairs, which is consistent with most offline algorithms theory.
-Therefore, the theory developed in this paper can be easily combined with other offline algorithm theories and easily used to build the theory of policy improvement guarantee, which can provide the theoretical basis for offline PbRL and facilitate the development of offline PbRL theory.
+The performance gap in offline PbRL is determined by two independent components, with the first term corresponding to the inherent gap of offline RL itself.
+Given this gap, the preference-based gap can be directly derived.
+Therefore, the theory developed in this paper can be easily combined with other offline algorithm theories and used to build a theory of policy improvement guarantee, providing a theoretical basis for offline PbRL and facilitating the development of offline PbRL theory.
 VIII Conclusion
@@
 The theoretical and experimental results demonstrate that the data selection mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve the performance of the reward model, and the performance learned by LEASE can be guaranteed under fewer preference datasets.
-Future work can focus on reducing this gap further under limited preference data or on achieving conservative estimation for state-action pairs where the learned reward model predicts inaccurately.
+Future works can focus on the following two aspects: 1) Since the trained transition model may have errors, which will affect the performance of the reward model to some extent, the quality of generated data can be considered in the future to further improve the accuracy of the reward function; 2) There is still a gap between the trained reward model and the real model.
+Other efficient methods of data augmentation can be used in the future to further narrow this gap.
 Appendix A Related Proofs
@@
 The Proof for Theorem 1: This proof is based on Proposition 1 and 2.
-Combing Eqs.
+Combining Eqs.
 (A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error $\widehat{\mathcal{L}}(\psi)$ holds:
@@
 $\leq$ $\sqrt{CN\log\left(\frac{\mathcal{N}_{\mathcal{R}}(1/N)}{\delta}\right)}.$
+Figure A.1: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) pen-expert and (b) hammer-expert datasets.
 Step 2: Bound the reward difference $|R^{*}(s,a)-\widehat{R}(s,a)|$.
@@
 This completes proof of Proposition 3.
-Figure A.1: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) pen-expert and (b) hammer-expert datasets.
 TABLE A.1: Comparison results between URLHF using fewer preference data and LEASE, where the offline RL algorithm is CQL and URLHF using fewer data is denoted as $\texttt{URLHF}^{*}$.
@@
 A.1 shows the comparison between the prediction value by the learned rewards and their ground truths for LEASE, FEWER, and FRESH under Adroit tasks, where the offline algorithm is IQL.
-The predicted and true rewards are both normalized to $[0,1]$.
-We randomly sample $500$ data from unlabeled datasets that are not seen in the training stage for evaluation, like Fig.
+For evaluation, we randomly sample $500$ data points from the unlabeled dataset that are not seen during the training phase, as illustrated in Fig.
 4.
-Since the reward has a value in certain positions in Adroit tasks, the figure is in the form of multiple scattered lines in Fig.
-A.1.
+Because rewards in Adroit tasks are limited to a few specific values (i.e., discrete states), while rewards in Mujoco tasks are continuous, Fig.
+A.1 is presented as multiple scatter lines, in contrast to Fig.
+4.
 As evidenced by the experimental results shown in this figure, the reward model trained under the LEASE framework achieves the smallest variance in predicted reward values, which shows that the data augmentation and selection mechanism can improve the reward model’s performance.
@@
 For COMBO hyperparameters, the probability confidence $\kappa_{p}$, and uncertainty variance $\kappa_{\tau}$ are set as $0.85$ and $0.05$ for all tasks, respectively.
-Note that in our framework, model-based methods do not necessarily perform better than model-free methods.
-Model-based RL methods focus on how to learn a conservative policy by regularizing $Q$ values or penalizing rewards to alleviate the effects of inaccurate model data.
-These methods critically depend on the accuracy of the reward model.
-Consequently, model-based approaches are more sensitive to reward modeling errors than model-free methods, which can hinder the improvement of agent performance.
-Therefore, model-based offline RL requires a higher accuracy of the reward model than model-free RL.
+It is important to note that, within our framework, model-based methods do not necessarily outperform model-free approaches.
+Model-based offline reinforcement learning typically emphasizes learning conservative policies through techniques such as regularizing Q-values or penalizing rewards, which mitigates the impact of inaccuracies in the model-generated data.
+However, these methods rely heavily on the accuracy of the reward model.
+Consequently, model-based approaches are more sensitive to errors in the reward model than their model-free counterparts, which may constrain potential improvements in agent performance.
+Therefore, model-based offline reinforcement learning generally requires a higher degree of reward model accuracy compared to model-free methods.
 A-E4 Further analysis for the results on benchmark tasks